### v2

# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

Abstract
In this article, we propose and study a stochastic preconditioned Douglas–Rachford splitting method to solve saddle-point problems that have separable dual variables. We prove the almost sure convergence of the iteration sequences in Hilbert spaces for a class of convex-concave and nonsmooth saddle-point problems. We also provide the sublinear convergence rate for the ergodic sequence concerning the expectation of the restricted primal-dual gap functions. Numerical experiments show the high efficiency of the proposed stochastic preconditioned Douglas–Rachford splitting methods.
Key words.
Douglas-Rachford splitting, stochastic optimization, saddle-point problems, almost sure convergence, preconditioning techniques.
MSC codes.
65K10, 65F08, 49K35, 90C25.
1 Introduction
The Douglas–Rachford (henceforth abbreviated as DR) splitting method is an iterative splitting method for finding a zero point of the sum of maximal set-valued monotone operators, which is a powerful optimization tool for convex optimization and has plenty of applications [3, 8, 14]. The method originated from splitting linear operators for solving heat conduction equations [13] and was extended to a more general nonlinear maximal monotone operator splitting framework [18, 23]. The convergence of the DR splitting method in Hilbert spaces can be found in [14] through a proximal point framework. The DR splitting method is essentially an implicit and unconditionally stable splitting algorithm for arbitrary step sizes. Solving implicit equations including linear equations can be computationally demanding for DR methods. In [8], within the generalized proximal-point framework, an efficient preconditioned DR framework with a convergence guarantee aiming at a specific class of convex-concave saddle-point problems is developed. It can incorporate the classical preconditioning techniques for linear systems into the nonlinear DR iterations. We refer to [7] for applications of preconditioned DR and [6] for its accelerated variants.
Due to a wide range of applications in various fields including big data science and machine learning, there are lots of studies on stochastic first-order algorithms [4, 11] and especially for saddle-point problems [1, 10, 11]. In this paper, we study a stochastic preconditioned DR splitting framework which includes the stochastic DR without preconditioners [11] as a special case for solving the following saddle-point problems with separable dual variables [1, 11].
$\min_{x\in\operatorname{dom}{\mathcal{F}}}\max_{y\in\operatorname{dom}{\mathcal{G}}}\mathcal{L}(x,y),\quad\mathcal{L}(x,y):=\mathcal{F}(x)+\sum_{i=1}^{n}(\langle\mathcal{K}_{i}x,y_{i}\rangle-\mathcal{G}_{i}(y_{i})).$ (1.1)
where $\mathcal{F}:\mathbb{X}\rightarrow\mathbb{R}_{\infty}:=\mathbb{R}\cup\{+\infty\}$ and each $\mathcal{G}_{i}:\mathbb{Y}_{i}\rightarrow\mathbb{R}_{\infty}$ are proper, convex and lower semi-continuous. Here, $\mathbb{X}$ and $\mathbb{Y}_{i}$ are real Hilbert spaces and each $\mathcal{K}_{i}:\mathbb{X}\rightarrow\mathbb{Y}_{i}$ is linear and continuous. The space $\mathbb{Y}$ is a product space, i.e., $\mathbb{Y}=\prod^{n}_{i=1}\mathbb{Y}_{i}$ with $y=(y_{1},y_{2},\ldots,y_{n})^{T}$. The saddle-point problem (1.1) can be reformulated as the following common primal-dual formulation [3]:
$\min_{x\in\operatorname{dom}{\mathcal{F}}}\max_{y\in\operatorname{dom}{\mathcal{G}}}\mathcal{L}(x,y),\quad\mathcal{L}(x,y):=\mathcal{F}(x)+\langle\mathcal{K}x,y\rangle-\mathcal{G}(y),$ (1.2)
where the dual function $\mathcal{G}$ is separable, i.e., $\mathcal{G}(y)=\sum^{n}_{i=1}\mathcal{G}_{i}(y_{i})$ with $\mathcal{G}:\mathbb{Y}\rightarrow\mathbb{R}_{\infty}$ and $\mathcal{K}=[\mathcal{K}_{1},\mathcal{K}_{2},\ldots,\mathcal{K}_{n}]^{T}:\mathbb{X}\rightarrow\mathbb{Y}$. When strong duality holds [3], the solutions of (1.1) are exactly the solutions of the following primal and dual problems:
$\min_{x\in\mathbb{X}}\mathcal{P}(x),\ \mathcal{P}(x):=\mathcal{F}(x)+\sum_{i=1}^{n}\mathcal{G}_{i}^{*}(\mathcal{K}_{i}x),\quad\max_{y\in\mathbb{Y}}\mathcal{D}(y),\ \mathcal{D}(y):=-\mathcal{F}^{*}(-\mathcal{K}^{*}y)-\sum_{i=1}^{n}\mathcal{G}_{i}(y_{i}),$ (1.3)
with $\mathcal{F}^{*}$ and $\mathcal{G}_{i}^{*}$ denoting the Fenchel conjugate functionals of $\mathcal{F}$ and $\mathcal{G}_{i}$, respectively.
Compared to related work, our contributions are as follows. First, different from the stochastic first-order primal-dual method [10] or stochastic classical DR splitting method without preconditioners [11], preconditioners are introduced and the challenges in the convergence analysis are addressed. Actually, preconditioners already introduced challenges for the deterministic preconditioned DR and extra effort is necessary for the corresponding convergence analysis [8, 6]. For the stochastic preconditioned DR (henceforth abbreviated as SPDR), the almost sure convergence framework of the stochastic fixed point iteration based on firmly nonexpansive operators developed in [11] cannot be applied directly here. We circumvent these difficulties by analyzing the relation between the stochastic updates and the one-step deterministic updates based on the current stochastic iteration. We mainly employed the celebrated stochastic quasi-Fejér monotonicity [11, 15] and Opial’s lemma [2, 20]. Second, we show the SPDR splitting method has a $\mathcal{O}(1/k)$ rate for the ergodic sequence respecting restricted primal-dual gap functions where $k$ is the iteration number. Since taking the expectation of the supremum of the primal-dual gap is a very subtle issue [1], a detailed analysis is devoted to this part. Finally, we show the high efficiency of the proposed SPDR compared to state-of-art methods including the stochastic primal-dual method [10] with some numerical experiments.
The rest of the paper is organized as follows. In section 2, we give a brief introduction of the preconditioned deterministic DR and the proposed stochastic preconditioned DR splitting method along with some necessary analysis tools. In section 3, we discuss the almost sure convergence of the proposed SPDR splitting method to a fixed point using stochastic quasi-Fejér monotonicity [11] and Opial’s lemma [2]. In section 4, we show the sublinear convergence rate of ergodic sequences respecting the expectation of the restricted primal-dual gap. In section 5, we extend these results to the stochastic and preconditioned DR method for quadratic-linear problems (henceforth abbreviated as SPDRQ). In section 6, we present detailed numerical experiments for total generalized variation (TGV) regularized problems to show the competitive performance of the proposed stochastic and preconditioned DR methods. Finally, we give some conclusions in section 7.
2 Preconditioned DR framework
Now, let us turn to the deterministic preconditioned Douglas–Rachford splitting method (PDR), which aims at solving the saddle-point problem (1.2). It reads as follows [8, 7, 6]:
$\left\{\begin{array}[]{ll}x^{k+1}=N_{1}^{-1}[(N_{1}-I)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{\ast}y^{k+1}],\\ y^{k+1}=N_{2}^{-1}[(N_{2}-I)y^{k}+\bar{y}^{k}+\tau\mathcal{K}x^{k+1}],\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=(I+\tau\partial\mathcal{G})^{-1}[2y^{k+1}-\bar{y}^{k}],\\ \bar{x}^{k+1}=\bar{x}^{k}+x_{\operatorname{test}}^{k+1}-x^{k+1},\\ \bar{y}^{k+1}=\bar{y}^{k}+y_{\operatorname{test}}^{k+1}-y^{k+1},\end{array}\right.$ (PDR)
where $\sigma$ and $\tau$ are arbitrary positive primal and dual step sizes, $N_{1}\geq I$, $N_{2}\geq I$ with $N_{1}:\mathbb{X}\rightarrow\mathbb{X}$ and $N_{2}:\mathbb{Y}\rightarrow\mathbb{Y}$ being bounded, linear and self-adjoint operators, and $x_{\operatorname{test}}^{k+1}$ and $y_{\operatorname{test}}^{k+1}$ are transitional variables. Denoting by $z=(x,y)\in\mathbb{X}\times\mathbb{Y}$ and $\bar{z}=(\bar{x},\bar{y})\in\mathbb{X}\times\mathbb{Y}$, the iteration (PDR) with $N_{1}=I$, $N_{2}=I$ and $\sigma=\tau$ becomes the original DR splitting method [14, 18] based on the following maximal monotone operator splitting by the first-order optimality conditions of (1.2),
$0\in Az+Bz,\ \ \text{with}\quad A=\begin{pmatrix}0&\mathcal{K}^{*}\\ -\mathcal{K}&0\end{pmatrix},\quad B=\begin{pmatrix}\partial\mathcal{F}&0\\ 0&\partial\mathcal{G}\end{pmatrix}.$
Note that the first two equations of (PDR) are implicit. However, if we choose $N_{2}=I$ and an operator $M$ such that
$N_{1}=M-\sigma\tau\mathcal{K}^{*}\mathcal{K}$ (2.1)
where $M:X\rightarrow X$ is linear, continuous, self-adjoint and satisfies $M\geq T:=I+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, we can update $x^{k+1}$ first with the following classical preconditioned iteration
$x^{k+1}=x^{k}+M^{-1}(b^{k}-Tx^{k}),\quad b^{k}=\bar{x}^{k}-\sigma\mathcal{K}^{*}\bar{y}^{k},$ (2.2)
and then obtain the linear update $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ according to the second equation of (PDR) followed by $x^{k+1}$ [8, 7, 6]. Classical preconditioning techniques can be employed with (2.2) which can be seen as an iterative method for solving $Tx^{k+1}=b^{k}$. This linear equation arises from the linear update step in the original Douglas–Rachford splitting method where $N_{1}=I$ and $N_{2}=I$. Preconditioned DR splitting can bring high efficiency with great flexibility together with convergence guarantees using finitely many iterations of any feasible preconditioner [8, 7, 6]. A preconditioner is called feasible if $M\geq T$ as in (2.2) includes a lot of classical efficient preconditioners [8]. Let us denote by $u:=(x,y,\bar{x},\bar{y})\in\mathbb{U}:=(\mathbb{X}\times\mathbb{Y})^{2}$ and introduce an associated positive semidefinite and possibly degenerate inner product as
$\langle u,u^{\prime}\rangle_{\mathbf{M}}:=\langle\mathbf{M}u,u^{\prime}\rangle=\frac{1}{\sigma}[\langle(N_{1}-I)x,x^{\prime}\rangle+\langle\bar{x},\bar{x}^{\prime}\rangle]+\frac{1}{\tau}[\langle(N_{2}-I)y,y^{\prime}\rangle+\langle\bar{y},\bar{y}^{\prime}\rangle],$ (2.3)
with $\mathbf{M}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\sigma}I,\tfrac{1}{\tau}I]$. Denoting by $\mathcal{T}:\mathbb{U}\rightarrow\mathbb{U}$ as the iteration operator that maps $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ to $u^{k+1}=(x^{k+1},y^{k+1},\bar{x}^{k+1},\bar{y}^{k+1})$ according to (PDR), we can rewrite the iteration (PDR) as
$u^{k+1}:=\mathcal{T}u^{k}.$ (2.4)
Now, let us give several lemmas which are convenient for analyzing the stochastic PDR method in the sequel, whose proofs can be found in [8, 7, 6] and are omitted here.
Lemma 1.
For ${u^{k}={(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})}}$, we have
$\frac{1}{\sigma}\langle x^{k+1}-\bar{x}^{k+1},x_{\operatorname{test}}^{k+1}\rangle+\frac{1}{\tau}\langle y^{k+1}-\bar{y}^{k+1},y_{\operatorname{test}}^{k+1}\rangle=\langle u^{k}-u^{k+1},u^{k+1}\rangle_{\mathbf{M}}.$ (2.5)
Hence, employing Lemma 1 yields the following proposition.
Proposition 1.
If ${u={(x,y,\bar{x},\bar{y})}}$ with $(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$, and $\bar{y}=y-\tau\mathcal{K}x$, then we have
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})$ $\leq\langle u^{k}-u^{k+1},u^{k+1}-u\rangle_{\mathbf{M}}$
$=\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}}^{2}).$ (2.6)
[…]
Lemma 2.
The iteration (PDR) for (1.2) possesses the following properties:
  • (i)
    A point $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ is a fixed point of (2.4) if and only if $(x^{*},y^{*})$ is a saddle point for (1.2) and $x^{*}+\sigma\mathcal{K}^{*}y^{*}=\bar{x}^{*},y^{*}-\tau\mathcal{K}x^{*}=\bar{y}^{*}$.
  • (ii)
    If w-${\lim_{j\rightarrow\infty}}(\bar{x}^{k_{j}},\bar{y}^{k_{j}})=(\bar{x}^{*},\bar{y}^{*})$ and $\lim_{j\rightarrow\infty}((N_{1}-I)(x^{k_{j}}-x^{k_{j}+1}),(N_{2}-I)(y^{k_{j}}-y^{k_{j}+1}),\bar{x}^{k_{j}}-\bar{x}^{k_{j}+1},\bar{y}^{k_{j}}-\bar{y}^{k_{j}+1})=(0,0,0,0)$, then $(x^{k_{j}+1},y^{k_{j}+1},\bar{x}^{k_{j}+1},\bar{y}^{k_{j}+1})$ converges weakly to a fixed point.
In the special case where $\mathcal{F}$ is a quadratic-linear form as follows:
$\mathcal{F}(x)=\langle\frac{1}{2}Qx-f,x\rangle,$
where $Q$ is a self-adjoint, linear, and bounded positive semi-definite operator and $f\in X$, we can get a more compact preconditioned Douglas–Rachford splitting method for quadratic problems, which is called PDRQ [8, 7, 6]. Denoting by $u=(x,y,\bar{y})\in\mathbb{U}_{Q}:=\mathbb{X}\times\mathbb{Y}\times\mathbb{Y}$, $N_{1}\geq 0$, $N_{2}\geq I$, the PDRQ iteration reads as,
$\left\{\begin{array}[]{l}x^{k+1}=(N_{1}+\sigma Q)^{-1}[N_{1}x^{k}-\sigma(\mathcal{K}^{*}y^{k+1}-f)],\\ y^{k+1}=N_{2}^{-1}[(N_{2}-I)y^{k}+\bar{y}^{k}+\tau\mathcal{K}x^{k+1}],\\ y_{\operatorname{test}}^{k+1}=(I+\tau\partial G)^{-1}[2y^{k+1}-\bar{y}^{k}],\\ \bar{y}^{k+1}=\bar{y}^{k}+y_{\operatorname{test}}^{k+1}-y^{k+1}.\end{array}\right.$ (PDRQ)
Moreover, if we choose $N_{2}=I$ and $N_{1}$ such that
$N_{1}=M_{Q}-\sigma Q-\sigma\tau\mathcal{K}^{*}\mathcal{K},$ (2.7)
where $M_{Q}:X\rightarrow X$ is linear, continuous, self-adjoint and feasible meaning $M_{Q}\geq T_{Q}:=\sigma Q+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, then we can update $x^{k+1}$ first with the following classical preconditioned iteration
$x^{k+1}=x^{k}+M_{Q}^{-1}(b^{k}-Tx^{k}),\quad b^{k}=\sigma f-\sigma\mathcal{K}^{*}\bar{y}^{k}.$ (2.8)
The linear update of $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ is then according to the second line in (PDRQ) [8, 7, 6]. Denoting by
$\mathbf{M}_{Q}=\operatorname{Diag}[\tfrac{1}{\sigma}N_{1},\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\tau}I],$
we have the associated inner product again,
$\langle u,u^{\prime}\rangle_{\mathbf{M}_{Q}}=\frac{1}{\sigma}\langle N_{1}x,x^{\prime}\rangle+\frac{1}{\tau}[\langle(N_{2}-I)y,y^{\prime}\rangle+\langle\bar{y},\bar{y}^{\prime}\rangle].$ (2.9)
Denoting by $\mathcal{T}_{Q}$ as the iteration operator associated with (PDRQ), we can rewrite the iteration (PDRQ) as
$u^{k+1}:=\mathcal{T}_{Q}u^{k}.$ (2.10)
Similarly, we have the following proposition and lemma for PDRQ.
Proposition 2.
If ${u={(x,y,\bar{y})}}$, with $\bar{y}=y-\tau\mathcal{K}x$, then we have the following estimate:
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})$ $\leq\langle u^{k}-u^{k+1},u^{k+1}-u\rangle_{\mathbf{M}_{Q}}$
$=\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}_{Q}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}_{Q}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}_{Q}}^{2}).$ (2.11)
Similar to Lemma 2, we give the following lemma with omitted proof, which will also turn out useful for the almost sure convergence of stochastic (2) method introduced in the paper.
Lemma 3.
The iteration (PDRQ) for (1.2) possesses the following properties:
  • (i)
    A point $u^{*}=(x^{*},y^{*},\bar{y}^{*})$ is a fixed point of (2.10) if and only if $(x^{*},y^{*})$ is a saddle point for (1.2) and $Qx^{*}+\mathcal{K}^{*}y^{*}=f$, $y^{*}-\tau\mathcal{K}x^{*}=\bar{y}^{*}$.
  • (ii)
    If w-${\lim_{j\rightarrow\infty}}\bar{y}^{k_{j}}=\bar{y}^{*}$ and ${\lim_{j\rightarrow\infty}}(N_{1}(x^{k_{j}}-x^{k_{j}+1}),(N_{2}-I)(y^{k_{j}}-y^{k_{j}+1}),\bar{y}^{k_{j}}-\bar{y}^{k_{j}+1})=(0,0,0)$, then $(x^{k_{j}+1},y^{k_{j}+1},\bar{y}^{k_{j}+1})$ converges weakly to a fixed point.
With these preparations for deterministic PDR (or PDRQ), let us now turn to our focus, to the stochastic PDR method.
2.1 The stochastic preconditioned DR (SPDR) method
Let us begin with the stochastic setting. Denote by $(\Omega,\mathfrak{F},\mathcal{P})$ the underlying probability space. Let $\{i_{0},i_{1},i_{2},\ldots\}$ be index selection variables, i.e., independent random variables on the latter space with values in $\{1,\ldots,n\}$ which are identically distributed. We denote the probability of selecting an index $i\in\{1,\ldots,n\}$ by $p_{i}$ where we assume $10<p_{i}<$ 1 for each $i$ and recall that $\sum_{i=1}^{n}p_{i}=1$. Introduce $P=\operatorname{Diag}[p_{1},p_{2},\ldots,p_{n}]$ and observe that $P^{-1}=\operatorname{Diag}[\frac{1}{p_{1}},\frac{1}{p_{2}},\ldots,\frac{1}{p_{n}}]$. A Borel random vector $y$ in a separable Hilbert space $\mathbb{H}$ is a measurable map from $\Omega$ into $\mathbb{H}$ where $\Omega$ is equipped with $\mathfrak{F}$ and $\mathbb{H}$ is equipped with the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the open sets of $\mathbb{H}$ (see [11] or Chapter 2 of [17]). Further, $(\mathfrak{F}_{k})_{k\in\mathbb{N}}$ denotes the filtration generated by the indices $\{i_{0},i_{1},\ldots,i_{k}\}$, i.e., the $\sigma$-algebras generated by $i_{j}^{-1}\{i\}$ for $j=0,\ldots,k$ and $i=1,\ldots,n$ [22] (Chapter 7). Then, $\mathfrak{F}_{k}\subset\mathfrak{F}_{k+1}$ for each $k\in\mathbb{N}$. Henceforth, we denote by $\mathbb{E}^{k}:=\mathbb{E}[\cdot\mid\mathfrak{F}_{k}]$ the conditional expectation with respect to $\mathfrak{F}_{k}$, and by $\mathbb{E}$ the full expectation. Recall that if a Borel random vector $y$ is $\mathfrak{F}_{k}$-measurable, i.e., the preimages of Borel sets are in $\mathfrak{F}_{k}$, then $\mathbb{E}[y\mid\mathfrak{F}_{k}]=y$.
Let us first extend the preconditioned Douglas–Rachford splitting method to a stochastic setting. For arbitrary positive step sizes $\sigma$, $\tau$ and $(x^{0},y^{0},\bar{x}^{0},\bar{y}^{0})\in(\mathbb{B}_{1}\times\mathbb{B}_{2})^{2}$, $y^{0}_{\operatorname{test}}\in\mathbb{B}_{2}$, the stochastic PDR (SPDR) reads as follows.
$\left\{\begin{array}[]{ll}x^{k+1}=\left(N_{1}+\sigma\tau\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(\left(N_{1}-I\right)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right),\\ y^{k+1}=\left\{\begin{array}[]{ll}N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}x^{k+1}]&i=i_{k},\\ y^{k}_{i}&\forall i\neq i_{k},\\ \end{array}\right.\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2y^{k+1}_{i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{x}^{k+1}=\bar{x}^{k}+(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}]-x^{k+1},\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+y_{\operatorname{test},i}^{k+1}-y_{i}^{k+1}&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k},\end{array}\right.\end{array}\right.$ (SPDR)
where $N_{1}\geq I$ and $N_{2}=\operatorname{Diag}[N_{2,1},\ldots,N_{2,n}]$ with each $N_{2,i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ being self-adjoint, bounded and linear such that $N_{2,i}\geq I$. Although the update of $x^{k+1}$ in the above (SPDR) seems complicated, however, it is convenient for analysis and the preconditioned iteration (2.2) can be recovered from (SPDR) with $N_{2}=I$ and $N_{1}$ as in (2.1).
Next, suppose that the bounded set $\mathbb{B}:=\mathbb{B}_{1}\times\mathbb{B}_{2}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ contains a saddle point. We give some distance functions including the following restricted primal-dual gap $\mathfrak{G}_{\mathbb{B}_{1}\times\mathbb{B}_{2}}$ associated with $\mathbb{B}$:
$\mathfrak{G}_{\mathbb{B}_{1}\times\mathbb{B}_{2}}(x,y):=\sup_{\tilde{y}\in\mathbb{B}_{2}}\mathcal{L}(x,\tilde{y})-\inf_{\tilde{x}\in\mathbb{B}_{1}}\mathcal{L}(\tilde{x},y):=\sup_{\tilde{z}\in\mathbb{B}}\mathcal{H}(z,\tilde{z}),\quad\mathcal{H}(z,\tilde{z}):=\mathcal{L}(x,\tilde{y})-\mathcal{L}(\tilde{x},y),$ (2.12)
where $z=(x,y)$ as before and $\tilde{z}:=(\tilde{x},\tilde{y})$. For any proper, convex, and lower semi-continuous function $\zeta:\mathbb{X}\rightarrow\mathbb{R}_{\infty}$, the Bregman distance at $s$ with $q\in\partial\zeta(s)$ is defined as
$D_{\zeta}^{q}(t,s)=\zeta(t)-\zeta(s)-\langle q,t-s\rangle.$
If $\tilde{z}$ is a saddle point, we can interpret $\mathcal{H}(z,\tilde{z})$ as the following Bregman distance for $h(z):=\mathcal{F}(x)+\mathcal{G}(y)$ with respect to the point $q:=(-\mathcal{K}^{*}\tilde{y},\mathcal{K}\tilde{x})$ which is included in the subgradient $\partial h(\tilde{z})$ by optimality, i.e.,
$\mathcal{H}(z,\tilde{z}):=D_{h}^{q}(z,\tilde{z})=D_{\mathcal{F}}^{-\mathcal{K}^{*}\tilde{y}}(x,\tilde{x})+D_{\mathcal{G}}^{\mathcal{K}\tilde{x}}(y,\tilde{y}).$
With these preparations, now let us turn to the almost sure convergence analysis of (SPDR).
3 Almost sure convergence of stochastic PDR
In this section, we present the almost sure convergence of iteration sequences generated by (SPDR) to a solution of (1.2). To avoid any ambiguity, henceforth, the notation $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ along with $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ always represents the stochastic sequences produced by (SPDR). In order to write the iteration in a more compact form, we introduce the notation $\tilde{u}^{k+1}$ for one step of the deterministic update of $u^{k}$, i.e.,
$\tilde{u}^{k+1}:=\mathcal{T}u^{k},\quad\text{with}\ \ \tilde{u}^{k+1}=(\tilde{x}^{k+1},\tilde{y}^{k+1},\tilde{\bar{x}}^{k+1},\tilde{\bar{y}}^{k+1}),$ (3.1)
where $\mathcal{T}$ is the same as (2.4) defined by (PDR). The stochastic update $u^{k+1}$ in (SPDR) can then be written as
$x^{k+1}=\tilde{x}^{k+1},\quad x_{\operatorname{test}}^{k+1}=\tilde{x}_{\operatorname{test}}^{k+1},\quad\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1},$ (3.2)
and
$y^{k+1}_{i}=\begin{cases}\tilde{y}^{k+1}_{i}&\text{if}\ i=i_{k},\\ y^{k}_{i}&\text{if}\ i\neq i_{k},\end{cases}\quad y^{k+1}_{\operatorname{test},i}=\begin{cases}\tilde{y}^{k+1}_{\operatorname{test},i}&\text{if}\ i=i_{k},\\ y^{k}_{\operatorname{test},i}&\text{if}\ i\neq i_{k},\end{cases}\quad\bar{y}^{k+1}_{i}=\begin{cases}\tilde{\bar{y}}^{k+1}_{i}&\text{if}\ i=i_{k},\\ \bar{y}^{k}_{i}&\text{if}\ i\neq i_{k}.\end{cases}$ (3.3)
Note that $x^{k},x^{k}_{\operatorname{test}},\bar{x}^{k},y^{k},y^{k}_{\operatorname{test}},\bar{y}^{k}$ according to (3.2) and (3.3) are $\mathfrak{F}_{k}$-measurable Borel random vectors by continuity of $\mathcal{T}$ and construction.
The stochastic iteration sequence $\{u^{k}\}$ and the corresponding one-step deterministic update $\tilde{u}^{k+1}$ can be shown in the following diagram
[…]
$\downarrow\ \;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\qquad\qquad\downarrow{\mathcal{T}}$
$\tilde{u}^{1}\,\ \quad\tilde{u}^{2}\,\ \quad\tilde{u}^{3}\,\ \quad\tilde{u}^{4}\,\ \quad\cdots\,\ \quad\tilde{u}^{k+1}=\mathcal{T}u^{k}\cdots.$
Let us begin the convergence analysis with the following estimate.
Lemma 4.
For any $(x,y)\in\operatorname{dom}{\mathcal{F}}\times\operatorname{dom}{\mathcal{G}}$, $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$ and $\bar{y}=y-\tau\mathcal{K}x$, it holds that
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\left(\|u^{k}-u\|^{2}_{\mathbf{M}_{p}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}\right),$ (3.5)
where $\mathbf{M}_{p}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\tau}(N_{2}-I)\mathbf{P}^{-1},\frac{1}{\sigma}I,\tfrac{1}{\tau}\mathbf{P}^{-1}]$ with $\mathbf{P}=\operatorname{Diag}[p_{1}I,\ldots,p_{n}I]$ and consequently, $\mathbf{P}^{-1}=\operatorname{Diag}[p_{1}^{-1}I,\ldots,p_{n}^{-1}I]$.
Proof.
With Proposition 1, $u=(x,y,\bar{x},\bar{y})$ and $\tilde{u}^{k+1}$ as in (3.1), (3.2) and (3.3), we arrive at
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\big{(}\|u^{k}-u\|^{2}_{\mathbf{M}}-\|\tilde{u}^{k+1}-u\|^{2}_{\mathbf{M}}-\|\tilde{u}^{k+1}-u^{k}\|^{2}_{\mathbf{M}}\big{)}$ (3.6)
Now, let us reformulate $\|\tilde{y}^{k+1}-\cdot\|_{\frac{(N_{2}-I)}{\tau}}^{2}$ and $\|\tilde{\bar{y}}^{k+1}-\cdot\|_{\tau^{-1}}^{2}$. First, noting that for $\mathfrak{F}_{k}$-measurable Borel random vectors $s^{k}=(s^{k}_{i})_{i=1,\ldots,n}$, $\tilde{s}^{k+1}=(\tilde{s}^{k+1}_{i})_{i=1,\ldots,n}$ on $\mathbb{Y}$, the update rule
$s_{i}^{k+1}=\left\{\begin{array}[]{ll}\tilde{s}_{i}^{k+1}&i=i_{k},\\ s_{i}^{k}&\forall i\neq i_{k},\end{array}\right.$
and any continuous $\varphi:Y\to\mathbb{R}$, it holds that
$\mathbb{E}^{k}\varphi(s_{i}^{k+1})=p_{i}\varphi(\tilde{s}_{i}^{k+1})+(1-p_{i})\varphi(s_{i}^{k}),$ (3.7)
which can be rewritten to
$\varphi(\tilde{s}_{i}^{k+1})=\frac{1}{p_{i}}\mathbb{E}^{k}\varphi(s_{i}^{k+1})-(\frac{1}{p_{i}}-1)\varphi(s_{i}^{k}).$ (3.8)
In particular, for any block diagonal operator $\mathbf{B}=\operatorname{Diag}[\mathbf{B}_{1},\ldots,\mathbf{B}_{n}]$ of symmetric and positive semidefinite operators $\mathbf{B}_{1},\ldots,\mathbf{B}_{n}$, this implies that
$\|\tilde{s}^{k+1}-\cdot\|_{\mathbf{B}}^{2}-\|s^{k}-\cdot\|_{\mathbf{B}}^{2}=\mathbb{E}^{k}\|s^{k+1}-\cdot\|_{\mathbf{P}^{-1}\mathbf{B}}^{2}-\|s^{k}-\cdot\|_{\mathbf{P}^{-1}\mathbf{B}}^{2}.$ (3.9)
Hence, using (3.9) with $\tilde{y}^{k+1}$ and $\tilde{\bar{y}}^{k+1}$ as $\tilde{s}^{k+1}$, $\tfrac{1}{\tau}(N_{2}-I)$ and $\tfrac{1}{\tau}I$ as $\mathbf{B}$, together with the fact that $x^{k+1}$ and $\bar{x}^{k+1}$ are $\mathfrak{F}_{k}$ measurable, we get
$\|\tilde{u}^{k+1}-u\|_{\mathbf{M}}^{2}-\|u^{k}-u\|_{\mathbf{M}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u\|_{\mathbf{M}_{p}}^{2}-\|u^{k}-u\|_{\mathbf{M}_{p}}^{2},$
$\|\tilde{u}^{k+1}-u^{k}\|_{\mathbf{M}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u^{k}\|_{\mathbf{M}_{p}}^{2},$
and consequently, (3.5). ∎
Moreover, we see, for instance,
$(N_{2}-I)\mathbf{P}^{-1}=(p_{1}^{-1}(N_{2,1}-I),\ldots,p_{n}^{-1}(N_{2,n}-I))=\mathbf{P}^{-1}(N_{2}-I).$
For this reason, we will tacitly use that $\mathbf{P}$ (and $\mathbf{P}^{-1}$) always commutes with the linear operators considered in this paper unless otherwise specified.
Let us give some notations first. Denote $\mathcal{Z}^{*}$ as the set for the saddle points of (1.1) and $\mathbb{F}\subset\mathbb{U}$ as the set for the fixed points for $\mathcal{T}$. Introduce $\mathcal{M}_{0}:=\operatorname{Diag}[{N_{1}-I},N_{2}-I,I,I]$ and
$\mathbb{F}_{m}=\mathbb{F}\mathcal{M}_{0}:=\{((N_{1}-I)x^{*},(N_{2}-I)y^{*},\bar{x}^{*},\bar{y}^{*})\ |\ u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}\}.$
Now, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$ to a fixed point of $\mathcal{T}$ as $k\to\infty$. Inspired by [11], we mainly employ the stochastic quasi-Fejér monotonicity [11], Opial’s lemma [2] (Lemma 21), and Lemma 2 to show the almost sure convergence of sequence $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ first. We then “glue” $\{u^{k+1},k\in\mathbb{N}\}$ to $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ and finish the proof.
Theorem 1.
The iteration sequence $\{u^{k}\}_{k}=\{(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})\}_{k}$ produced by (SPDR) weakly converges to a fixed point of $\mathcal{T}$, i.e., $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})\in\mathbb{F}$ almost surely, and $(x^{**},y^{**})\in\mathcal{Z}^{*}$ is a saddle point for (1.1) with $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**}$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover, the transitional sequence $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ converges weakly to the saddle point $(x^{**},y^{**})$ almost surely.
Proof.
Inserting $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ a fixed point of $\mathcal{T}$ in (3.5). By Lemma 2, $(x^{*},y^{*})$ is a saddle point of (1.1). Then it follows for every $(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}$, $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),(x^{*},y^{*}))\geq 0$. Denoting
$\Delta^{k}=\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p}},$
and employing Lemma 4 as well as $\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}=\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}$, we have
$\mathbb{E}(\Delta^{k+1}|\mathfrak{F}_{k})=\mathbb{E}^{k}\Delta^{k+1}\leq\Delta^{k}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}=\Delta^{k}-\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}.$ (3.10)
Using the Robbins–Siegmund theorem [21] (see also [11, 1]) on (3.10), we conclude that
$\sum_{k=0}^{\infty}\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}<\infty,\quad\|\mathcal{T}u^{k}-u^{k}\|^{2}_{\mathbf{M}}\rightarrow 0,$ (3.11)
almost surely in $\Omega$. Recalling the definition of $\mathbf{M}_{p}$, we immediately see that also $\|\mathcal{T}u^{k}-u^{k}\|_{\mathbf{M}_{p}}\to 0$ as $k\to\infty$ almost surely. Using further that $\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}=\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|$ for each $k$ and applying Proposition 2.3 in [11], we conclude that $\{\mathbf{M}_{p}^{1/2}u^{k}\}_{k}$ is bounded almost surely and that the limit
$l(u^{*}):=\lim_{k\rightarrow+\infty}\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}$ (3.12)
exists almost surely. In the following, we will establish the almost sure weak convergence of $\{u^{k}\}_{k}$. Our first goal is to show that $\{\tilde{u}^{k}\}_{k}$ converges weakly to a fixed point of $\mathcal{T}$. We will begin with proving the almost sure weak convergence of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ with Opial’s lemma [2] (Lemma 21). This can be done by showing that almost surely, $\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|$ converges and every weak sequential limit point of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k+1}\}_{k}$ belongs to $\mathbf{M}_{p}^{1/2}\mathbb{F}$. According to the triangle inequality, we have
$\big{|}\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|-\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|\big{|}\leq\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{k})\|=\|\mathcal{T}u^{k}-u^{k}\|_{\mathbf{M}_{p}},$
so
$\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|\rightarrow l(u^{*})$ (3.13)
almost surely as $k\to\infty$. In particular, $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ is bounded almost surely. Next, we show that for each event such that (3.11), (3.12) and (3.13) hold, each weak accumulation point of $\{\tilde{u}^{k}\}_{k}$ is a fixed point. For that purpose, note that due to (3.12), we have for the sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ and each $\mathbf{M}_{p}^{1/2}u^{*}\in\mathbf{M}_{p}^{1/2}\mathbb{F}$ that $\|\mathbf{M}_{p}^{1/2}\tilde{u}^{k}-\mathbf{M}_{p}^{1/2}u^{*}\|$ converges. Now, let $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ be a weakly converging subsequence. Then, as $\|\mathbf{M}_{p}^{1/2}(\mathcal{T}u^{k}-u^{k})\|\to 0$ as $k\to\infty$, $\{\mathbf{M}_{p}^{1/2}u^{k_{j}-1}\}_{j}$ converges weakly to the same limit and looking at the definition of $\mathbf{M}_{p}$, we conclude that $\bigl{\{}\bigl{(}(N_{1}-I)x^{k_{j}-1},(N_{2}-I)y^{k_{j}-1},\bar{x}^{k_{j}-1},\bar{y}^{k_{j}-1}\bigr{)}\bigr{\}}_{j}$ converges weakly. We further have that $\bigl{(}(N_{1}-I)(\tilde{x}^{k_{j}}-x^{k_{j}-1}),(N_{2}-I)(\tilde{y}^{k_{j}}-y^{k_{j}-1}),\tilde{\bar{x}}^{k_{j}}-\bar{x}^{k_{j}-1},\tilde{\bar{y}}^{k_{j}}-\bar{y}^{k_{j}-1}\bigr{)}\to 0$ as $j\to\infty$ such that with Lemma 2 (ii), we obtain that $\{(\tilde{x}^{k_{j}},\tilde{y}^{k_{j}},\tilde{\bar{x}}^{k_{j}},\tilde{\bar{y}}^{k_{j}})\}$ converges weakly to a fixed point $(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}$. In particular, by Lemma 2 (i), $(x^{**},y^{**})$ is a saddle point for (1.1) and $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**},y^{**}-\tau\mathcal{K}x^{**}=\bar{y}^{**}$. It is then immediate that the weak limit of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ belongs to $\mathbf{M}_{p}^{1/2}\mathbb{F}$. By Opial’s lemma, we have that the whole sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ converges weakly to an element in $\mathbf{M}_{p}^{1/2}\mathbb{F}$. Repeating the above arguments with $k_{j}=j+1$, it follows that $\{\tilde{u}^{k}\}_{k}$ converges weakly to the fixed point $u^{**}$ of $\mathcal{T}$. Finally, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$. First, note that since for fixed $i=1,\ldots,n$, the probability $p_{i}$ of $i_{k}=i$ is positive and all $i_{k}$ are independent random variables, the probability of $i_{k}=i$ only occurring finitely many times has to be zero. For that reason, almost surely, for each $i=1,\ldots,n$, the sequences $\{k_{i,j}\}_{j}$ such that $k_{i,j}=\max\ \{k:k\leq j,\ i_{k-1}=i\}$, i.e., the sequences that yield the index of the last activation of $i$, are increasing and satisfy $\lim_{j\to\infty}k_{i,j}=\infty$. Fix an event for which $\{\tilde{u}^{k}\}_{k}$ converges weakly to a fixed point $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}$ and for which $k_{i,j}$ tends to infinity as $j\to\infty$ for all $i=1,\ldots,n$. We will show that $\{u^{k}\}_{k}$ converges weakly, as $k\to\infty$, to $u^{*}$. For that purpose, observe that by (3.2), we have $x^{k}=\tilde{x}^{k}\to x^{**}$, $x^{k}_{\operatorname{test}}=\tilde{x}^{k}+\tilde{\bar{x}}^{k}-\bar{x}^{k-1}\to x^{**}$ and $\bar{x}^{k}=\tilde{\bar{x}}^{k}\to\bar{x}^{**}$ in the weak sense as $k\to\infty$. By (3.3) and the definition of $k_{i,j}$, we see that $y^{j}_{i}=\tilde{y}^{k_{i,j}}_{i}$, $y^{j}_{\operatorname{test},i}=\tilde{y}^{k_{i,j}}_{i}+\tilde{\bar{y}}^{k_{i,j}}_{i}-\bar{y}^{(k_{i,j})-1}_{i}$ and $\bar{y}^{j}_{i}=\tilde{\bar{y}}^{k_{i,j}}_{i}$ for each $j$. Hence, letting $j\to\infty$, we see that $y^{j}_{i}\to y^{**}_{i}$, $y^{j}_{\operatorname{test},i}\to y^{**}_{i}$ and $\bar{y}^{j}_{i}\to\bar{y}^{**}_{i}$ in the weak sense. As the events for which this follows occur almost surely, the proof is complete. ∎
Table 1: Existing research and our contributions
Preconditioner setting Almost sure convergence Restricted primal-dual gap
$N_{1}=I;N_{2}=I$ Already have in [11] Presented in this paper
$N_{1}>I;N_{2}>I$ Easy to prove with [1] Presented in this paper
$N_{1}\geq I,N_{1}\neq I;N_{2}=I$ Presented in this paper Presented in this paper
As shown in Table 1, our framework covers some results in existing research. The commonly used preconditioners for stochastic PDR are the case $N_{1}\geq I,N_{1}\neq I$ and $N_{2}=I$, leading to difficulties in the proof. Next, let us turn to the convergence rate of the restricted primal-dual gap, which is also a challenge due to the preconditioners and the transitional variables.
4 Sublinear convergence of stochastic PDR for restricted primal-dual gap
Taking full expectation in (3.10), induction with $k$ yields
$\mathbb{E}[\Delta^{k}]\leq\Delta^{0}=\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}},\quad\sum_{i=0}^{k}\mathbb{E}[\|u^{i+1}-u^{i}\|^{2}_{\mathbf{M}_{p}}]\leq\Delta^{0}.$ (4.1)
Let us begin with the following lemma which gives the control of $\mathcal{T}u^{k}$ by the weighted $\mathbf{M}_{p}$ norm of $u^{k}$. Applying Lemma 2.5 in [8] with $\mathbf{M}_{p}$, we have the following estimate.
Lemma 5.
There exists a constant $C>0$ such that for all $u^{1},u^{2}\in\mathbb{U}$ and $\tilde{u}^{1}=\mathcal{T}u^{1}$, $\tilde{u}^{2}=\mathcal{T}u^{2}$ satisfies
$\|\tilde{u}^{1}-\tilde{u}^{2}\|^{2}\leq C\|u^{1}-u^{2}\|_{\mathbf{M}_{p}}^{2},$ (4.2)
where the constant $C$ is only related to $N_{1}$, $N_{2}$, $\sigma$, $\tau$ and $\mathcal{K}$.
The following simple lemma can deal with the main challenge in $y^{k+1}_{\operatorname{test}}$ for the gap estimate.
Lemma 6.
Supposing $\delta\in(0,1)$ is a constant, given a non-decreasing sequence $S_{k}$ and a bounded sequence $h_{k}$ satisfying $0<S_{k}\leq h_{k}+(1-\delta)S_{k-1}$ for $k\geq 1$, we have
$S_{k}\leq\sum_{i=0}^{k-1}(1-\delta)^{i}h_{k-i}+(1-\delta)^{k}S_{0}.$ (4.3)
Proof.
An induction of $S_{k}$ gives (4.3). Furthermore, if $h_{k}$ is bounded by $M$, we have $S_{k}\leq[1-(1-\delta)^{k}]M/{\delta}+(1-\delta)^{k}S_{0}\leq{M}/{\delta}+(1-\delta)^{k}S_{0}$. ∎
Lemma 7.
The estimates of $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ and $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|$ are as follows. First, with notations $s^{k}$ and $\varphi$ defined in (3.7), we deduce
  • (i)
    $\mathbb{E}\varphi(s_{i}^{k+1})=\sum_{n=0}^{k}\mathbb{E}[p_{i}(1-p_{i})^{n}\varphi(\tilde{s}_{i}^{k+1-n})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$
  • (ii)
    $\mathbb{E}\|s^{k+1}-s^{k}\|^{2}_{P^{-1}-I}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k}\|^{2}_{I-P}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k+1}\|^{2}.$
Then, it follows
  • (iii)
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}\leq n(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C$ is as in Lemma 5.
  • (iv)
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}\leq D\Delta^{0}$, where $D>0$ is a positive constant.
  • (v)
    $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}\leq M_{K}$ with $M_{K}\leq M_{\infty}$ and $M_{\infty}$ being a positive constant.
Proof.
For (i), since $\mathbb{E}\mathbb{E}^{k}[\cdot]=\mathbb{E}[\cdot]$, a direct induction gives,
$\mathbb{E}\varphi(s_{i}^{k+1})$ $=\mathbb{E}\mathbb{E}^{k}\varphi(s_{i}^{k+1})=\mathbb{E}[p_{i}\varphi(\tilde{s}_{i}^{k+1})+(1-p_{i})\varphi(s_{i}^{k})]$
$=\mathbb{E}p_{i}\varphi(\tilde{s}_{i}^{k+1})+\mathbb{E}\mathbb{E}^{k-1}(1-p_{i})\varphi(s_{i}^{k})$
$=\mathbb{E}[p_{i}\varphi(\tilde{s}_{i}^{k+1})+p_{i}(1-p_{i})\varphi(\tilde{s}_{i}^{k})+(1-p_{i})^{2}\varphi(s_{i}^{k-1})]=\cdots$
$=\mathbb{E}[\sum_{n=0}^{k}p_{i}(1-p_{i})^{n}\varphi(\tilde{s}_{i}^{k+1-n})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$ (4.4)
which leads to (i). For (ii), the trick is to find the relations among those three items. That is, for every $k$, $\|\tilde{s}^{k+1}-s^{k}\|^{2}=\|s^{k+1}-s^{k}\|^{2}+\|\tilde{s}^{k+1}-s^{k+1}\|^{2}$ since $\|s_{i}^{k+1}-s_{i}^{k}\|^{2}$ will be zero except $i=i_{k}$ and $\|\tilde{s}_{i}^{k+1}-s_{i}^{k+1}\|^{2}$ will be zero except $i\neq i_{k}$. Using $\mathbb{E}\|s^{k+1}-s^{k}\|^{2}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k}\|_{P}^{2}$, the statement (ii) follows. Now, let us prove $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test},i}-y_{i}^{*}\|^{2}$ is uniformly bounded for all $k$.
$\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test},i}-y_{i}^{*}\|^{2}$ $=\mathbb{E}\|\tilde{\bar{y}}_{i}^{k+1}-\bar{y}_{i}^{k}+\tilde{y}_{i}^{k+1}-y_{i}^{*}\|^{2}\leq\mathbb{E}\|\tilde{\bar{y}}^{k+1}-\bar{y}^{*}+\bar{y}^{*}-\bar{y}^{k}+\tilde{y}^{k+1}-y^{*}\|^{2}$
$\leq 3\mathbb{E}\|\tilde{\bar{y}}^{k+1}-\bar{y}^{*}\|^{2}+3\mathbb{E}\|\tilde{y}^{k+1}-y^{*}\|^{2}+3\tau\mathbb{E}\|\bar{y}^{k}-\bar{y}^{*}\|_{\tau^{-1}P^{-1}}^{2}$
$\leq 3\mathbb{E}\|\mathcal{T}u^{k}-u^{*}\|^{2}+3\tau\mathbb{E}\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p}}$
$\leq(3C+3\tau)\mathbb{E}\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p}}\leq(3C+3\tau)\Delta^{0}$ (4.5)
where the third inequality is deduced from Lemma 5 while the last one is from (4.1). Applying (i) with $\varphi(\cdot)=(\cdot-y^{*}_{i})^{2}$ on $y^{k+1}_{\operatorname{test},i}$ and noting $10<\sum_{n=0}^{k}p_{i}(1-p_{i})^{n}<$ 1, $(1-p_{i})^{k+1}<1$, we arrive at
$\mathbb{E}\|y^{k+1}_{\operatorname{test},i}-y_{i}^{*}\|^{2}$ $=\mathbb{E}\sum_{n=0}^{k}{p_{i}(1-p_{i})^{n}}\|\tilde{y}^{k+1-n}_{\operatorname{test},i}-y_{i}^{*}\|^{2}+(1-p_{i})^{k+1}\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2},$
$\leq(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}.$ (4.6)
Summing (4.6) over all the components for $i=1,\ldots,n$ leads to (iii). Now, let us turn to (iv). Remembering the updates of $\tilde{y}^{k+1}_{\operatorname{test}}$, $\tilde{y}^{k+1}$, $x^{k+1}$ as in (SPDR) and the non-expansiveness of the resolvent operator $(I+\tau\partial\mathcal{G})^{-1}$, it can be readily checked that
$\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}\leq\mathbb{E}\sum_{k=1}^{K}\|(2\tilde{y}^{k+1}-\bar{y}^{k})-(2\tilde{y}^{k}-\bar{y}^{k-1})\|^{2}$
$=\mathbb{E}\sum_{k=1}^{K}\|A_{1}(N_{1}-I)(x^{k}-x^{k-1})+A_{1}(\bar{x}^{k}-\bar{x}^{k-1})+A_{2}(N_{2}-I)(y^{k}-y^{k-1})$
$\quad+A_{3}(\bar{y}^{k}-\bar{y}^{k-1})\|^{2}\leq D\mathbb{E}\sum_{k=1}^{K}\|u^{k}-u^{k-1}\|_{\mathbf{M}_{p}}^{2},$ (4.7)
where $A_{1}=\tau A_{2}\mathcal{K}N_{1}^{-1}$, $A_{2}=2(N_{2}+\tau\sigma\mathcal{K}N_{1}^{-1}\mathcal{K}^{*})^{-1}$, $A_{3}=A_{2}-I$, and
$D=4\max({\sigma(\|A_{1}\|\|(N_{1}-I)\|)^{2},\ {\sigma}\|A_{1}\|^{2},\ \tau(\|A_{2}\|\|(N_{2}-I)\|)^{2},\ {\tau}\|A_{3}\|^{2})}$ (4.8)
Combining (4.7) and (4.1), we obtain (iv). For (v), since $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|_{P}^{2}$, we henceforth focus on $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$. By the inequality $\|a+b\|^{2}\leq(1+\epsilon)\|a\|^{2}+(1+{1}/{\epsilon})\|b\|^{2}$, for any $\epsilon>0$ and $K\geq 1$, we arrive at
$\sum_{k=1}^{K}\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}\leq\sum_{k=1}^{K}[\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+\mathbb{E}(1+\frac{1}{\epsilon})\|\tilde{y}^{k}_{\operatorname{test}}-y^{k}_{\operatorname{test}}\|^{2}]$
$=\sum_{k=1}^{K}\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+\sum_{k=1}^{K}\mathbb{E}(1+\frac{1}{\epsilon})\|\tilde{y}^{k}_{\operatorname{test}}-y^{k-1}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}$ (4.9)
$=\sum_{k=1}^{K}\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+(\sum_{k=1}^{K-1}\mathbb{E}(1+\frac{1}{\epsilon})\|\tilde{y}^{k+1}_{\operatorname{test}}-y^{k}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}+\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}),$
where the first equality is obtained by (ii). For the convenience of proofs, letting $\epsilon={1}/{\underline{p}}$ with $\underline{p}:=\min_{i}\{p_{i},\ i=1,\ldots,n\}$ henceforth, then $(1+{1}/{\epsilon})(I-\mathbf{P})\leq(1+{1}/{\epsilon})(1-\underline{p})I=(1-\underline{p}^{2})I$. For employing Lemma 6, choosing $\delta=\underline{p}^{2}$ and $S_{0}=0$, we denote for $K\geq 1$,
$S_{K}:=\sum_{k=1}^{K}\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2},\ \ h_{K}:=\sum_{k=1}^{K}\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}.$
The equation (4.9) thus turns to
$0<S_{K}-(1-\delta)S_{K-1}\leq h_{K}.$ (4.10)
For $\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}$, we have
$\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}\leq\mathbb{E}\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}=\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|^{2}_{(1-\delta)I}.$ (4.11)
Then by (iv), $h_{K}$ is bounded, i.e.,
$h_{K}\leq(1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}$ (4.12)
and $\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{1-\delta}^{2}$ is given by (4.11). Thus by (4.10), (4.12), Lemma 6 and the fact $S_{0}=0$, we have
$S_{K}$ $\leq\frac{1-(1-\delta)^{K-1}}{\delta}\left((1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)\leq\frac{1}{\delta}\left((1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right).$
Hence
$\mathbb{E}\sum_{k=0}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$ $=\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}+\mathbb{E}\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}$
$\leq\frac{1-(1-\delta)^{K-1}}{\delta}\left((1+\epsilon)D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
Finally, with (ii) and $\delta=\underline{p}^{2}$, $\epsilon=1/\underline{p}$, we derive
$\mathbb{E}$ $\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\sum_{k=0}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|_{\mathbf{P}}^{2}\leq\mathbb{E}\sum_{k=0}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$
$\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+1/\underline{p})D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{K},$ (4.13)
$\leq{1}/{\underline{p}^{2}}\left((1+1/\underline{p})D\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{\infty}.$
Now, let us extend a lemma in [1] originated from [19] to positive semidefinite weights.
Lemma 8.
With notations $s^{k}$ defined in Lemma 4, picking up a point $s^{0}_{\dagger}\in\mathbb{Y}$ at will, let us define an iteration sequence $v^{k+1}=s^{k}-\tilde{s}^{k+1}-\mathbf{P}^{-1}(s^{k}-s^{k+1})$ with ${s}_{\dagger}^{k+1}={s}_{\dagger}^{k}-\mathbf{P}v^{k+1}$. Then, it follows that for any $s\in\mathbb{Y}$ and a positive semidefinite, self-adjoint, and linear weight $\mathfrak{D}=\emph{Diag}[\mathfrak{D}_{1},\ldots,\mathfrak{D}_{n}]$ with $\mathfrak{D}_{i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ and $\mathfrak{D}\mathbf{P}=\mathbf{P}\mathfrak{D}$ as before,
$\sum_{k=0}^{K}\langle{s}_{\dagger}^{k}-s,v^{k+1}\rangle_{\mathfrak{D}}=\frac{1}{2}\|{s}_{\dagger}^{0}-s\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{s}_{\dagger}^{K+1}-s\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2},$ (4.14)
and
$\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2}]\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|s^{k}-s^{k+1}\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}].$ (4.15)
Moreover, $v^{k}$ and ${s}_{\dagger}^{k}$ are $\mathfrak{F}_{k}$ measurable and $\mathbb{E}^{k}[v^{k+1}]=0$. It follows that
$\mathbb{E}^{k}\langle{s}_{\dagger}^{k},v^{k+1}\rangle=0.$ (4.16)
Proof.
By the definition of $s^{k+1}_{\dagger}$, we have the estimate,
$\frac{1}{2}\|{s}_{\dagger}^{k+1}-s\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}$ $=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}-\langle\mathbf{P}v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathbf{P}^{-1}\mathfrak{D}}+\frac{1}{2}\|\mathbf{P}v^{k+1}\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}$
$=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}}^{2}-\langle v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathfrak{D}}+\frac{1}{2}\|v^{k+1}\|_{\mathbf{P}\mathfrak{D}}^{2}.$ (4.17)
Summing the equality (4), we arrive at (4.14). For the second result, similar to $\mathbb{E}^{k}\|v^{k+1}\|^{2}=\mathbb{E}^{k}\|\mathbb{E}^{k}\xi-\xi\|^{2}\leq\mathbb{E}^{k}\|\xi\|^{2}$, with $\xi^{k}=\mathbf{P}^{-1}(s^{k}-s^{k+1})$ and $\mathbb{E}^{k}\xi=s^{k}-\tilde{s}^{k+1}$, we obtain
$\mathbb{E}[\sum_{k=0}^{K}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2}]=\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}}^{2}]]$
$\leq\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|\mathbf{P}^{-1}(s^{k+1}-s^{k})\|_{\mathfrak{D}\mathbf{P}}^{2}]]=\sum_{k=0}^{K}\mathbb{E}[\|s^{k+1}-s^{k}\|_{\mathfrak{D}\mathbf{P}^{-1}}^{2}].$ (4.18)
By direct calculation, $\mathbb{E}^{k}[v^{k+1}]=0$. Since ${s}_{\dagger}^{k}$ is $\mathfrak{F}_{k}$ measurable, we conclude $\mathbb{E}^{k}\langle{s}_{\dagger}^{k},v^{k+1}\rangle=\langle{s}_{\dagger}^{k},\mathbb{E}^{k}v^{k+1}\rangle=0$. ∎
With the preceding lemmas, we now prove the sublinear convergence of (SPDR). The main ideas are employing the relation between the stochastic Bregman distance $\mathcal{H}(z_{\operatorname{test},K},z)$ and the deterministic Bregman distance $\mathcal{H}(\tilde{z}_{\operatorname{test},K},z)$ together with separating the operation $\mathbb{E}$ from the operation $\sup_{z\in\mathbb{B}}$ when acting on $\mathcal{H}(z_{\operatorname{test},K},z)$.
Theorem 2.
If $u=(x,y,\bar{x},\bar{y}),$ with $\bar{x}=x+\sigma\mathcal{K}^{\ast}y,$ $\bar{y}=y-\tau\mathcal{K}x$, and $z^{k}_{\operatorname{test}}=(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, a bounded domain, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K},$ (4.19)
where the constant is given by
$C_{\mathbb{B}}=\sup_{z\in\mathbb{B}}\frac{1}{2}\Big{[}\|x^{0}-x\|_{\frac{N_{1}-I}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{x}^{0}-x-\sigma\mathcal{K}^{*}y\|_{\frac{1}{\sigma}}+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|_{\frac{\mathbf{P}^{-1}}{\tau}}$
$+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}+\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2}\Big{]}$
$+\frac{1}{2}(\Delta^{0}+\frac{1}{\underline{p}}M_{K})+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(n(3C+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})$
$-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}.$
Here $C$ is produced in Lemma 5, $M_{K}$ is from (4.13), $\Delta^{0}=\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}$, $(x^{*},y^{*})$ is a saddle point, and $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$, $y^{0}_{\dagger,\operatorname{test}}$ are picked from $\mathbb{Y}$ randomly. The same rate also holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
Proof.
Thanks to the identities
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)=\mathcal{F}(x_{\operatorname{test}}^{k+1})+\langle\mathcal{K}x_{\operatorname{test}}^{k+1},y\rangle-\mathcal{G}(y)-\mathcal{F}(x)-\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}\rangle+\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}),$
$\mathcal{H}((x_{\operatorname{test}}^{k+1},{y}_{\operatorname{test}}^{k+1}),z)=\mathcal{F}(x_{\operatorname{test}}^{k+1})+\langle\mathcal{K}x_{\operatorname{test}}^{k+1},y\rangle-\mathcal{G}(y)-\mathcal{F}(x)-\langle\mathcal{K}x,{y}_{\operatorname{test}}^{k+1}\rangle+\mathcal{G}({y}_{\operatorname{test}}^{k+1}),$
and these give rise to
$\mathcal{H}((x_{\operatorname{test}}^{k+1},{y}_{\operatorname{test}}^{k+1}),z)=\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)+\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})+\mathcal{G}(y_{\operatorname{test}}^{k+1}).$ (4.20)
We now separate $\mathcal{H}((x_{\operatorname{test}}^{k+1},{y}_{\operatorname{test}}^{k+1}),z)$ into three parts including $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$, $\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle$ and $\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})$. By the convexity of $\mathcal{H}$ and the property of supremum, we have
$\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}(z^{K}_{\operatorname{test}},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]$
$+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))].$ (4.21)
We will discuss these three parts one by one. First, denote by
$\mathbf{x}^{k}:=(x^{k},\bar{x}^{k}),\quad\mathbf{y}^{k}:=(y^{k},\bar{y}^{k}),\quad\tilde{\mathbf{y}}^{k}:=(\tilde{y}^{k},\tilde{\bar{y}}^{k}),\quad\mathbf{x}:=(x,\bar{x}),\quad\mathbf{y}:=(y,\bar{y}),$ (4.22)
and two diagonal weights $\mathbf{w}_{x}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\sigma}I]$, $\mathbf{w}_{y}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\tau}I]$. Therefore, $\|u^{k}-u\|^{2}_{\mathbf{M_{p}}}=\|\mathbf{x}^{k}-\mathbf{x}\|^{2}_{\mathbf{w}_{x}}+\|\mathbf{y}^{k}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}$, in which $\mathbf{w}_{y}\mathbf{P}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I)\mathbf{P},\tfrac{1}{\tau}\mathbf{P}]$. For $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$, we focus on the iteration $u^{k}\rightarrow\tilde{u}^{k+1}$. Combining Proposition 1 and (3.2), we obtain
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)\leq\Delta_{\mathbf{x}}^{k}+\Delta_{\mathbf{y}}^{k},$ (4.23)
where $\Delta_{\mathbf{x}}$ and $\Delta_{\mathbf{y}}$ are defined as
$\begin{array}[]{c}\Delta_{\mathbf{x}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}}^{2}\Big{)},\\ \Delta_{\mathbf{y}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}\Big{)}.\\ \end{array}$ (4.24)
Let us first reformulate $\Delta_{\mathbf{y}}^{k}$ as follows
$\Delta_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}+\|\mathbf{y}^{k}-\ y\|_{\mathbf{w}_{y}}^{2}$
$+(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\Big{]}$
$=-\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}+\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}+\epsilon_{\mathbf{y}}^{k},$ (4.25)
where $\epsilon_{\mathbf{y}}^{k}$ is defined as follows with $\mathbf{v}^{k+1}:=\mathbf{y}^{k}-\tilde{\mathbf{y}}^{k+1}-\mathbf{P}^{-1}(\mathbf{y}^{k}-\mathbf{y}^{k+1})$ for integer $k\geq 0$,
$\epsilon_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}}^{2}-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\Big{]}$
$=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}}^{2}-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})-2\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}\Big{]}.$ (4.26)
Now, with (4.23), (4.24) and (4), we arrive at
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}}^{2}$
$+\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}\Big{)}+\epsilon_{\mathbf{y}}^{k}$
$=\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{p}}-\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p}}-\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}\Big{)}$
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}+\frac{1}{2}\Big{(}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}}$
$+\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}\textbf{}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}\Big{)},$ (4.27)
where the equality is by adding and subtracting $\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}$ and the definition of $\epsilon_{\mathbf{y}}^{k}$ in (4.26). In (4.27), dropping $-\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}}$ and summing over $k=0,\ldots,K$ yields,
$\sum^{K}_{k=0}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p}}+\frac{1}{2}\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\mathcal{E}^{K}+\sum_{k=0}^{K}[-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}],$ (4.28)
where $\mathcal{E}^{K}$ is defined as follows
$\frac{1}{2}\sum_{k=0}^{K}\left[(\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}})+(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\right].$
Noting there are no items involving $(x,y)$ in $\mathcal{E}^{K}$, we can take expectation directly without concerning taking the supremum over $(x,y)$. With (3.9) and the fact $\mathbb{E}[\mathbb{E}^{k}[\cdot]]=\mathbb{E}[\cdot]$, we get
$\mathbb{E}\mathcal{E}^{K}/K$ $=\frac{1}{2K}\mathbb{E}\sum_{k=0}^{K}\Big{[}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}-\mathbb{E}^{k}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}$
$+\mathbb{E}^{k}(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2})\Big{]}=0.$ (4.29)
For the sum of inner products in (4.28), introducing $\mathbf{y}^{k+1}_{\dagger}=\mathbf{y}^{k}_{\dagger}-\mathbf{P}\mathbf{v}^{k+1}$ and selecting a $\mathbf{y}^{0}_{\dagger}\in\mathbb{B}_{2}\times\mathbb{B}_{2}$ randomly, we have
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}=\langle\mathbf{y}_{\dagger}^{k}-\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}.$
In (4.28), dropping $-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p}}$, dividing both sides by $K$, taking supremum with respect to $z$, and employing (4.14) on the above inner product, we arrive at
$\mathbb{E}[$ $\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\sup_{z\in\mathbb{B}}\frac{1}{2K}[\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\|\mathbf{y}^{0}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}P^{-1}}]$ (4.30)
$-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}+\mathbb{E}\sum_{k=0}^{K}\frac{1}{K}[\frac{1}{2}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}}].$
Dropping $-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}}$, noting $\mathbb{E}\sum_{k=0}^{K}\frac{1}{2}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}]$ by (4.15), together with (4.16) and (4.1), we can obtain
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\frac{1}{2K}[\sup_{z\in\mathbb{B}}(\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\|\mathbf{y}^{0}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}})+\Delta^{0}].$ (4.31)
Second, let us give the upper bounds for $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$. Denote by $v_{\operatorname{test}}^{k+1}=y_{\operatorname{test}}^{k}-\tilde{y}_{\operatorname{test}}^{k+1}-\mathbf{P}^{-1}(y_{\operatorname{test}}^{k}-y_{\operatorname{test}}^{k+1})$, $y^{k+1}_{\dagger,\operatorname{test}}=y^{k}_{\dagger,\operatorname{test}}-\mathbf{P}v_{\operatorname{test}}^{k+1}$ and select a $y^{0}_{\dagger,\operatorname{test}}\in\mathbb{Y}$ randomly. By the definition of $y^{k}_{\dagger,\operatorname{test}}$, $v^{k}_{\operatorname{test}}$ and (4.14), we obtain
$\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle$
$=$ $\sum_{k=0}^{K}-\langle\mathcal{K}x,y_{\operatorname{test}}^{k}-\tilde{y}_{\operatorname{test}}^{k+1}-\mathbf{P}^{-1}(y_{\operatorname{test}}^{k}-{y}_{\operatorname{test}}^{k+1})\rangle+\sum_{k=0}^{K}\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k})\rangle$
$=$ $\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}}-\mathcal{K}x,v^{k+1}_{\operatorname{test}}\rangle-\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}},v^{k+1}_{\operatorname{test}}\rangle+\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-{y}_{\operatorname{test}}^{0})\rangle$
$=$ $\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{K+1}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}-\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}},v^{k+1}_{\operatorname{test}}\rangle$
$+\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-{y}_{\operatorname{test}}^{0})\rangle.$ (4.32)
For the last term in (4.32), we have
$\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0})\rangle\leq\frac{1}{2}\|\mathcal{K}x\|^{2}+\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2},$ (4.33)
In (4.32), dropping $-\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{K+1}-\mathcal{K}x\|^{2}$ and by (4.33), (4.16), we get
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$ $\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K}\|\mathcal{K}x\|^{2}]$
$+\frac{1}{2K}\mathbb{E}\sum_{k=0}^{K}[\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}+\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}.$ (4.34)
For the last item in (4.34), inserting $y^{*}$ and employing Lemma 7, we get,
$\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}$ $\leq\mathbb{E}\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}\|y_{\operatorname{test}}^{K+1}-y^{*}\|^{2}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}\|y^{*}-y_{\operatorname{test}}^{0}\|^{2}$
$\leq\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}(n(3C+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2}).$ (4.35)
Noting $\mathbb{E}\sum_{k=0}^{K}[\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}]\leq\mathbb{E}\sum_{k=0}^{K}\|y_{\operatorname{test}}^{k}-y_{\operatorname{test}}^{k+1}\|_{\mathbf{P}^{-1}}^{2}$ by (4.15) together with (4.13), finally, we arrive at
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K}\|\mathcal{K}x\|^{2}]$
$+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K}(n(3C+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}K}M_{K}.$ (4.36)
Third, we show $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]$ is also bounded. In fact,
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]=\mathbb{E}\frac{1}{K}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})].$ (4.37)
Employing (3.8) on $\mathcal{G}(\tilde{y}^{k+1}_{\operatorname{test}})$, we get $\mathcal{G}(\tilde{y}^{k+1}_{\operatorname{test}})=\mathbb{E}^{k}\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}})-\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})$. Here $\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}}):=\sum^{n}_{i=1}1/p_{i}\mathcal{G}_{i}(y^{k+1}_{\operatorname{test},i})$ and $\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k+1}_{\operatorname{test}}):=\sum^{n}_{i=1}(1/p_{i}-1)\mathcal{G}_{i}(y^{k+1}_{\operatorname{test},i})$. Thus, using the fact $\mathbb{E}\mathbb{E}[\cdot]=\mathbb{E}[\cdot]$ again, we arrive at
$\mathbb{E}\frac{1}{K}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]=\frac{1}{K}\sum_{k=0}^{K}\mathbb{E}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathbb{E}^{k}\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}})+\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]$
$=\frac{1}{K}\sum_{k=0}^{K}[-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k+1}_{\operatorname{test}})+\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]=\frac{1}{K}(\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}k}+1}_{\operatorname{test}})).$ (4.38)
Since $\mathcal{K}_{i}x^{*}\in\partial\mathcal{G}_{i}(y^{*}_{i})$, one has
$\mathcal{G}_{i}(y_{\operatorname{test},i}^{K+1})\geq\mathcal{G}_{i}(y_{i}^{*})+\langle\mathcal{K}_{i}x^{*},y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\rangle\geq\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\|y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\|.$
We thus obtain
$\mathbb{E}[\mathcal{G}_{\mathbf{P}^{-1}-I}(y_{\operatorname{test}}^{K+1})$ $]=\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathbb{E}[\mathcal{G}_{i}(y_{\operatorname{test},i}^{K+1})]\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\mathbb{E}[\|y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\|])$
$\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\sqrt{\mathbb{E}[\|y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\|^{2}]})$
$\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}),$ (4.39)
where the second inequality is by Jensen’s inequality and the third inequality is by Lemma 7. We thus have
$\mathbb{E}\frac{1}{K}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]$ $\leq\frac{1}{K}\Big{(}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})$
$+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}\Big{)}.$ (4.40)
Finally, with (4.21) and combining (4.31), (4.36) and (4.40), we get (4.19) with the constant $C_{\mathbb{B}}$. For the Bregman distance, replacing $z=(x,y)$ by $(x^{*},y^{*})$, one can obtain the same ergodic convergence rate of the ergodic sequences with similar derivation. ∎
5 Stochastic PDR for quadratic problems (SPDRQ)
In this section, we will focus on the following stochastic PDRQ method
$\left\{\begin{array}[]{ll}x^{k+1}=\left(N_{1}+\sigma Q+\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(N_{1}x^{k}+\sigma f-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right)\\ y^{k+1}=\left\{\begin{array}[]{ll}\tilde{y}^{k+1}_{i}=N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}x^{x+1}]&i=i_{k}\\ y^{k}_{i}&\forall i\neq i_{k}\\ \end{array}\right.\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}\tilde{y}_{\operatorname{test},i}^{k+1}=(I+\tau\partial\mathcal{G}_{i})^{-1}[2\tilde{y}_{i}^{k+1}-\bar{y}_{i}^{k}]&i=i_{k}\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k}\\ \end{array}\right.\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\tilde{\bar{y}}^{k+1}_{i}=\bar{y}_{i}^{k}+y_{\operatorname{test},i}^{k+1}-y_{i}^{k+1}&i=i_{k}\\ \bar{y}^{k}_{i}&\forall i\neq i_{k}\\ \end{array}\right.\\ \end{array}\right.$ (SPDRQ)
where $N_{1}\geq 0$, $N_{2}\geq I$ and $N_{2}$ is separable, i.e., $N_{2}y=(N_{2,1}y_{1},\ldots,N_{2,n}y_{n})^{T}$. The preconditioned iteration (2.8) can be recovered from the update $x^{k+1}$ in the above (SPDRQ) with $N_{2}=I$ and $N_{1}$ as in (2.7). We define $x_{\operatorname{test}}^{k+1}=x^{k+1}$. Similarly, with $\mathcal{T}_{Q}$ defined in (2.10), we do one-step deterministic update based on the $u^{k}$ and get $\tilde{u}^{k+1}$, i.e.,
$\tilde{u}^{k+1}:=\mathcal{T}_{Q}u^{k},\quad\text{with}\ \ \tilde{u}^{k+1}=(\tilde{x}^{k+1},\tilde{y}^{k+1},\tilde{\bar{y}}^{k+1}),$ (5.1)
where $x^{k+1}=\tilde{x}^{k+1}$ and $\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1}$. With Proposition 2 and Lemma 3, we can derive the following lemma and theorem similarly.
Lemma 9.
For any $(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}$, a bounded domain, $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$ and $\bar{y}=y-\tau\mathcal{K}x$, denoting $\mathbf{M}_{Qp}:=\operatorname{Diag}[\tfrac{1}{\sigma}N_{1},\tfrac{1}{\tau}(N_{2}-I)\mathbf{P}^{-1},\tfrac{1}{\tau}\mathbf{P}^{-1}]$, it holds that
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y))$ $\leq\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{Qp}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{Qp}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{Qp}}\Big{)}.$ (5.2)
Theorem 3.
The iteration sequence $u^{k}=(x^{k},y^{k},\bar{y}^{k})$ produced by (SPDRQ) converges weakly to a fixed point of $\mathcal{T}_{Q}$, i.e., $u^{**}=(x^{**},y^{**},\bar{y}^{**})$ almost surely, and $(x^{**},y^{**})\in\mathcal{Z}^{*}$ is one of the saddle points of the problem (1.1) with $Qx^{**}+\mathcal{K}^{*}y^{**}=f$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover the sequence $(x^{k+1}_{\operatorname{test}}$, $y^{k+1}_{\operatorname{test}})$ also converges weakly to the saddle point $(x^{**},y^{**})$ almost surely.
Proof.
Inserting $u^{*}=(x^{*},y^{*},\bar{y}^{*})$, a fixed point of $\mathcal{T}_{Q}$, in (5.2), we conclude that $(x^{*},y^{*})$ is a saddle point by Lemma 3. It follows that $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),(x^{*},y^{*}))\geq 0$. Denoting
$\Delta^{k+1}=\|u^{k+1}-u^{*}\|^{2}_{\mathbf{M}_{Qp}},$
we have
$\mathbb{E}^{k}\Delta^{k+1}\leq\Delta^{k}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{Qp}}=\Delta^{k}-\|\mathcal{T}_{Q}u^{k}-u^{k}\|^{2}_{\mathbf{M}_{Q}}.$ (5.3)
The remaining proof is similar to Theorem 1 and we omit it here. ∎
For the gap estimate, we give the following two lemmas which are similar to the SPDR case.
Lemma 10.
There exists a constant $C_{Q}>0$ such that for all $u^{1},u^{2}\in\mathbb{U}_{Q}$ and $\tilde{u}^{1}=\mathcal{T}_{Q}u^{1}$, $\tilde{u}^{2}=\mathcal{T}_{Q}u^{2}$ and it leads to
$\|\tilde{u}^{1}-\tilde{u}^{2}\|^{2}\leq C_{Q}\|u^{1}-u^{2}\|_{\mathbf{M}_{Qp}}^{2}$ (5.4)
where $C_{Q}$ is only related to $N_{1}$, $N_{2}$, $f$, $\sigma$, $\tau$ and $\mathcal{K}$.
Lemma 11.
Similar to Lemma 7, we give the following estimates of $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ and $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|$.
  • (i)
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}<n(3C_{Q}+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C_{Q}$ is given in Lemma 10.
  • (ii)
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}<D_{Q}\Delta^{0}$, where
    $D_{Q}=3\max(\sigma(\|N_{1}\|\|A_{1}\|)^{2},\ \tau(\|(N_{2}-I)\|\|A_{2}\|)^{2},\ \tau\|A_{3}\|^{2})$
    and $A_{2}=2(N_{2}+\tau\sigma\mathcal{K}(N_{1}+\sigma Q)^{-1}\mathcal{K}^{*})^{-1}$, $A_{1}=A_{2}\tau\mathcal{K}(N_{1}+\sigma Q)^{-1}$, $A_{3}=A_{2}-I$.
  • (iii)
    $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}<M_{Q,K}$, where
    $M_{Q,K}=\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}((1+1/\underline{p})D_{Q}\Delta^{0}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{1-\underline{p}^{2}}^{2})+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
For the restricted gap estimate for (SPDRQ), we have the following theorem similarly.
Theorem 4.
If $u=(x,y,\bar{y}),$ with $\bar{y}=y-\tau\mathcal{K}x$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ produced by (SPDRQ) converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}$, a bounded domain, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K},$ (5.5)
where the constant $C_{\mathbb{B}}$ is given by,
$C_{\mathbb{B}}=\sup_{z\in\mathbb{B}}\frac{1}{2}\Big{[}\|x^{0}-x\|_{\frac{N_{1}}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|_{\frac{\mathbf{P}^{-1}}{\tau}}$
$+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}+\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2}\Big{]}$
$+\frac{1}{2}(\Delta^{0}+\frac{1}{\underline{p}}M_{Q,K})+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(n(3C_{Q}+3\tau)\Delta^{0}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})$
$-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{(3C_{Q}+3\tau)\Delta^{0}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}].$ (5.6)
Here $C_{Q}$ is produced in Lemma 10, $M_{Q,K}$ is from Lemma 11, $\Delta^{0}=\|u^{0}-u^{*}\|_{\mathbf{M}_{Qp}}^{2}$, $(x^{*},y^{*})$ is a saddle point, and $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$ and $y^{0}_{\dagger,\operatorname{test}}$ are picked from $\mathbb{B}_{2}$ randomly. Then the same rate holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
6 Numerical Experiments
In our experiments, we focus on the following image deblurring problem (6.1) based on the Kullback-Leibler (KL) divergence and the anisotropic second-order Total Generalized Variation (TGV) [5] regularization
$\min_{u\in U}\mathcal{F}(u)+\mathcal{G}_{1}^{*}(\mathcal{K}_{1}u)+\operatorname{TGV}_{\alpha}(u)=\min_{u\in U,w\in V}\mathcal{F}(u)+\mathcal{G}_{1}^{*}(\mathcal{K}_{1}u)+\alpha_{1}\|\nabla u-w\|_{1}+\alpha_{0}\|\mathcal{E}w\|_{1}.$ (6.1)
Here $U$ and $V$ are finite-dimensional Hilbert spaces and $w=(w^{1},w^{2})$. The two-dimensional problem (6.1) is standard and all definitions along with discretizations will be made clear afterward. Let us first reformulate the problem (6.1) as the saddle-point problem. $\mathcal{G}_{1}^{*}$ is an appropriate data fidelity measure with noisy data. We employ the following KL divergence which is widely used in the presence of Poisson noise
$\mathcal{G}^{*}_{1}(s)=\sum_{i=1}^{N}\begin{cases}s_{i}-b_{i}\log s_{i}&\text{if }s_{i}\geq 0,\\ \infty&\text{ else.}\end{cases}$ (6.2)
We also use the convention $00\log 0=$ 0 and $-a\log(0)=\infty$ for $a>0$. It can be readily checked that its Fenchel conjugate is
$\mathcal{G}_{1}(z)=\sum_{i=1}^{N}\begin{cases}\frac{-b_{j}z_{j}}{1-z_{j}}-\frac{b_{j}}{1-z_{j}}+b_{j}\log(\frac{b_{j}}{1-z_{j}}),&\text{ if }z_{j}\leq 1\text{ and }(b_{j}=0\text{ or }z_{j}<1),\\ \infty&\text{ else, }\end{cases}$ (6.3)
and its resolvent is $(I+\tau\partial\mathcal{G}_{1})^{-1}(z)=\frac{1}{2}(z+1-\sqrt{(z-1)^{2}+4\tau b})$. Here $\mathcal{F}(u)=\mathcal{I}_{\{0\leq u\leq 1\}}(u)$ is a box constraint. It forces $u$ to be non-negative by the requirement of KL divergence and to have an upper bound $1$. The equivalent saddle-point problem of (6.1) is as follows
$\min_{u\in U,w\in V}\max_{s\in U,p\in V,q\in W}\langle\mathcal{K}_{1}u,s\rangle_{U}+\langle\nabla u-w,p\rangle_{V}+\langle\mathcal{E}w,q\rangle_{W}+\mathcal{F}(u)$ (6.4)
$-\mathcal{G}_{1}(s)-\sum_{i=1}^{2}\mathcal{I}_{\{\|p^{i}\|_{\infty}\leq\alpha_{1}\}}(p^{i})-\sum_{j=1}^{4}\mathcal{I}_{\{\|q^{j}\|_{\infty}\leq\alpha_{0}\}}(q^{j}).$
Let us reformulate (6.4) into (1.1). We have the data $\mathcal{F}(x)=\mathcal{F}(u)=\mathcal{I}_{\{0\leq u\leq 1\}}(u)$ with $x=(u,w)$ and $\mathcal{G}^{*}_{i}(y_{i})=\alpha_{1}\|y_{i}\|_{1},$ $i=2,3$, $\mathcal{G}^{*}_{i}(y_{i})=\alpha_{0}\|y_{i}\|_{1},$ $i=4,5,6,7$. $\mathcal{G}_{i}$ is the Fenchel conjugate of $\mathcal{G}_{i}^{*}$. The dual variables $y=(y_{1},\ldots,y_{7})=(s,p,q)$ with $p=(p^{1},p^{2})\in V:=U^{2}$ and $q=(q^{1},q^{2},q^{3},q^{4})\in W:=U^{4}$ is thus obtained. The linear operator $\mathcal{K}:X\rightarrow Y$ as in (1.1) is
$\mathcal{K}=\left(\begin{array}[]{cc}\mathcal{K}_{1}&0\\ \nabla&-I\\ 0&\mathcal{E}\end{array}\right),\quad X=U\times V,\quad Y=U\times V\times W,$ (6.5)
where $\mathcal{K}_{1}=\kappa\ast$ is a convolution operator with kernel $\kappa$ for deblurring problems.
We then go to the details of the discrete setting including the operators and spaces. In (6.1), $\mathcal{E}w=(\nabla w+\nabla w^{T})/2$ is a discrete symmetrized derivative. $\nabla$ is a discrete gradient operator with finite forward differences, i.e., $(\nabla u)=(\partial_{x}^{+}u,\partial_{y}^{+}u)^{T}$ [9]. For discretization, set image domain $\Omega\subset\mathbf{Z}^{2}$ as the discretized grid $\Omega=\{(i,j)\mid i,j\in\mathbb{N},0\leq i\leq d_{1}-1,0\leq j\leq d_{2}-1\}$, where $d_{1},d_{2}$ are the image dimensions [8]. The image space is defined as $U=\{u:\Omega\rightarrow\mathbf{R}\}$ with the standard $L^{2}$ scalar product. For $t\in[1,\infty)$, let us introduce the following norms: $\|u\|_{\infty}=\max_{(i,j)\in\Omega}|u_{i,j}|$,
$\begin{array}[]{l}\|u\|_{t}=(\sum_{(i,j)\in\Omega}|u_{i,j}|^{t})^{1/t},\ \ \|p\|_{1}=\sum_{l=1}^{2}\sum_{(i,j)\in\Omega}|p_{i,j}^{l}|,\ \ \|q\|_{1}=\sum_{l=1}^{4}\sum_{(i,j)\in\Omega}|q_{i,j}^{l}|.\end{array}$
The adjoint operator of $\nabla:U\rightarrow V$ is $\nabla^{*}=-$ div (e.g., $\operatorname{div}p=\partial_{x}^{-}p^{1}+\partial_{y}^{-}p^{2}$) with homogeneous Neumann and Dirichlet boundary conditions, respectively and we refer to [9] for details. Next we describe the discretization of the anisotropic symmetric derivative $\mathcal{E}$ in the space $W=\{u:\Omega\rightarrow S^{2\times 2}\}$, which leads to the anisotropic second-order TGV which makes it possible to separate dual variables. We identify elements $q\in W$ and employ the scalar product $\langle q,q^{\prime}\rangle_{W}=\langle q^{1},(q^{\prime})^{1}\rangle_{U}+\langle q^{2},(q^{\prime})^{2}\rangle_{U}+\langle q^{3},(q^{\prime})^{3}\rangle_{U}+\langle q^{4},(q^{\prime})^{4}\rangle_{U}$. The symmetrized derivative can be defined as follows
$\mathcal{E}w=\left[\begin{array}[]{cc}\partial_{x}^{+}w^{1}&\frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})\\ \frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})&\partial_{y}^{+}w^{2}\end{array}\right]=\left[\begin{array}[]{c}\partial_{x}^{+}w^{1}\\ \partial_{y}^{+}w^{2}\\ \frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})\\ \frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})\end{array}\right]$ (6.6)
where the second equation has to be understood in terms of the identification $W=U^{4}$. Consequently, the negative adjoint realizes a discrete negative divergence operator according to $\langle\mathcal{E}w,q\rangle_{W}=-\langle w,\operatorname{div}q\rangle_{V}$ for all $w\in V,q\in$ $W$. With $\Delta w=(\Delta w^{1},\Delta w^{2}),$ we can express the operator *$\mathcal{E}^{*}\mathcal{E}w=-\frac{1}{2}\Delta w-\frac{1}{2}\delta w$. Here $\operatorname{div}q$ and $\delta w$ are gives as
$\operatorname{div}q:=\left[\begin{array}[]{l}\partial_{x}^{-}q^{1}+\frac{1}{2}\partial_{y}^{-}q^{3}+\frac{1}{2}\partial_{y}^{-}q^{4}\\ \frac{1}{2}\partial_{x}^{-}q^{3}+\frac{1}{2}\partial_{x}^{-}q^{4}+\partial_{y}^{-}q^{2}\end{array}\right],\quad\delta w:=\left[\begin{array}[]{l}\partial_{x}^{-}\partial_{x}^{+}w^{1}+\partial_{y}^{-}\partial_{x}^{+}w^{2}\\ \partial_{y}^{-}\partial_{y}^{+}w^{2}+\partial_{x}^{-}\partial_{y}^{+}w^{1}\end{array}\right].$
Moreover, we confirm that $q^{3}=q^{4}$ and only $6$ dual variables will be randomly updated.
6.1 The preconditioner
Setting $N_{2}=I$, our goal is to find efficient preconditioners to solve $Tx^{k+1}=b^{k}$, with $x^{k+1}=(u^{k+1},w^{k+1})$ and $b^{k}=\bar{x}^{k}-\sigma\mathcal{K}^{*}\bar{y}^{k}=(b_{u}^{k},b_{w}^{k})$ and
$T=\left[\begin{array}[]{cc}\sigma\tau\mathcal{K}_{1}^{*}\mathcal{K}_{1}+I-\sigma\tau\Delta&\sigma\tau\mathrm{div}\\ -\sigma\tau\nabla&(1+\sigma\tau)I+\sigma\tau\mathcal{E}^{*}\mathcal{E}\end{array}\right].$ (6.7)
$T$ is challenging for preconditioning due to $\mathcal{E}^{*}\mathcal{E}$. Fortunately, it can be readily checked that
$T^{\prime}=\left[\begin{array}[]{cc}(1+\sigma\tau)I-\sigma\tau\Delta&\sigma\tau\operatorname{div}\\ -\sigma\tau\nabla&(1+\sigma\tau)I-\sigma\tau\Delta\end{array}\right]$ (6.8)
satisfies $T^{\prime}-T\geq 0$ [7]. It is clear that feasible preconditioners $M$ for $T^{\prime}$ are also feasible for $T$ (Since $M\geq T^{\prime}$, we have $M\geq T$). Consequently, we obtain
$x^{k+1}=x^{k}+M^{-1}\left(b^{k}+\left(T^{\prime}-T_{1}\right)x^{k}-T^{\prime}x^{k}\right),$ (6.9)
which is a preconditioned iteration for $T^{\prime}$. Now we adjust the right-hand side as follows.
$\left(b^{\prime}\right)^{k}$ $:=b^{k}+\left(T^{\prime}-T_{2}\right)x^{k}$ $=\left[\begin{array}[]{c}b_{u}^{k}+(\sigma\tau I-\sigma\tau\mathcal{K}_{1}^{*}\mathcal{K}_{1})u^{k}\\ b_{w}^{k}+\frac{\sigma\tau}{2}\left(\delta-\Delta\right)w^{k}\end{array}\right].$ (6.10)
Then the block symmetric Gauss–Seidel preconditioner for $T^{\prime}$ as in [7] is employed.
6.2 Parameters
Here let us set $U=\mathbb{R}^{d_{1}\times d_{2}}$ with $d_{1}=512$, $d_{2}=357$ and $\mathcal{K}_{1}=k\ast$ is a convolution kernel with $40\times 40$ motion blur. It can be checked that the Bregman distance for the ergodic sequence is bounded for a bounded domain containing a saddle point.
Set the sampling to be uniform, i.e., $p_{i}=\frac{1}{n},\,n=6$. The regularization parameters are $\alpha_{0}=10^{-4}$ and $\alpha_{1}=5\times 10^{-5}$. The random iteration sequence is determined by random seed 5. The parameters for compared algorithms are as follows:
$\bullet$ PDHG, SPDHG: Primal-Dual Hybrid Gradient algorithm (PDHG) [9] and stochastic PDHG [10]. We use constant step size $\tau=\sigma=1/\|\mathcal{K}\|$, $\|\mathcal{K}\|=\sqrt{14}$ for PDHG and $\tau=0.25$, $\sigma=0.12$ for SPDHG, which is much faster than the step size in [10].
$\bullet$ PDR, SPDR: Preconditioned Douglas–Rachford method (PDR) [8], Stochastic PDR as in (SPDR). Here, one-step symmetric Red-Black Gauss–Seidel is the preconditioner for the elliptic system equation. The step sizes are chosen as $\tau=0.1,\,\sigma=5$.
Now, let us turn to the numerical ergodic Bregman distance. It can be checked that $x^{k+1}_{\operatorname{test}}$, $s^{k+1}_{\operatorname{test}}$ and $y^{k+1}_{\operatorname{test},i},$ $i=\{2,\ldots,7\}$ produced by (SPDR) satisfy the constraints of $\mathcal{F}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{i},$ $i=\{2,\ldots,7\}$ correspondingly. Still, by the convexity of these functions, the ergodic sequence $(x^{k+1}_{\operatorname{test},K},y^{k+1}_{\operatorname{test},K})$ also satisfies the corresponding constraints. We thus conclude that the Bregman distance of ergodic sequence is finite and can be written as $\langle\mathcal{K}x_{\operatorname{test},K},y^{*}\rangle-\mathcal{G}_{1}(s^{*})-\langle\mathcal{K}x^{*},y_{\operatorname{test},K}\rangle+\mathcal{G}_{1}(s_{\operatorname{test},K})$. Here $(x^{*},y^{*})$ is the approximate saddle point which is calculated for more than $10^{5}$ epochs.
(a) Convergence rate of normalized ergodic Bregman distance concerning epochs.
(b) Convergence rate of normalized primal error concerning epochs.
(c) PSNR with respect to epochs.
Figure 1: TGV-KL deblurring problem. (a), $\emph{(}b)$ are normalized by $d_{1}\times d_{2}$ and in a double-logarithmic scale. Image $\emph{(}b)$ shows the speed comparison measured in terms of primal error, which is given by $(\mathcal{P}(x^{K})-\mathcal{P}(x^{*}))/\mathcal{P}(x^{*})$.
(a) Degraded image
(b) PDHG, 300 epochs
(c) PDR, 300 epochs
(d) Approximate primal part of saddle point
(e) SPDHG, 300 epochs
(f) SPDR, 300 epochs
(g) Details of a
(h) Details of PDHG
(i) Details of PDR
(j) Details of b
(k) Details of SPDHG
(l) Details of c
Figure 2: Examples of variational TGV-KL deblurring after 300 epochs with uniform sampling.
Figure 1 tells that the Bregman distance does converge with rate $\mathcal{O}(1/K)$ experimentally. It also can be seen that the proposed SPDR is highly efficient and competitive compared with SPDHG regarding either the Bregman distance or the normalized primal error. Furthermore, the SPDR is more efficient than PDR compared with epochs and SPDR can get a higher and more stable PSNR as shown in Figure 1. Figure 2 show the SPDR can give clearer reconstructions compared with PDR, SPDHG, and PDHG with the same epochs.
7 Conclusion
We proposed a stochastic and preconditioned DR framework to solve general saddle-point problems for separable dual variables cases. We showed the almost sure convergence of SPDR along with SPDRQ and given the $\mathcal{O}(1/k)$ convergence rate of the ergodic iteration sequences regarding the restricted primal-dual gap functions. For SPDR, we will consider local linear convergence involving metric subregularity as in [1] and design highly efficient preconditioners for more complicated operators in applications including the Radon transforms [16] for tomography.
Acknowledgements YD and HS acknowledge the support of the National Natural Science Foundation of China under grant No.  12271521, Beijing Natural Science Foundation No. Z210001, and National Key R&D Program of China (2022ZD0116800). YD also acknowledges the support of the Austrian Science Fund (FWF) project P 34981. The Institute of Mathematics and Scientific Computing, to which KB is affiliated, is a member of NAWI Graz (https://www.nawigraz.at/en/). The author KB is a member of BioTechMed Graz (https://biotechmedgraz.at/en/).