### v3

# A stochastic preconditioned Douglas-Rachford splitting method for saddle-point problems

Abstract
In this article, we propose and study a stochastic and relaxed preconditioned Douglas–Rachford splitting method to solve saddle-point problems that have separable dual variables. We prove the almost sure convergence of the iteration sequences in Hilbert spaces for a class of convex-concave and nonsmooth saddle-point problems. We also provide the sublinear convergence rate for the ergodic sequence concerning the expectation of the restricted primal-dual gap functions. Numerical experiments show the high efficiency of the proposed stochastic and relaxed preconditioned Douglas–Rachford splitting methods.
Key words.
Douglas–Rachford splitting, stochastic optimization, saddle-point problems, over-relaxation, almost sure convergence, preconditioning techniques.
MSCodes.
65K10, 65F08, 49K35, 90C25.
1 Introduction
The Douglas–Rachford (henceforth abbreviated as DR) splitting method is an iterative splitting method for finding a zero point of the sum of maximal set-valued monotone operators, which is a powerful optimization tool for convex optimization and has plenty of applications [4, 10, 17, 25, 19]. The method originated from splitting linear operators for solving heat conduction equations [18] and was extended to a more general nonlinear maximal monotone operator splitting framework [32, 41]. The convergence of the DR splitting method in Hilbert spaces can be found in [19] through a proximal point framework. We further refer to [11, 26, 27, 22] for general DR splitting methods for convex problems and [2, 33, 42, 31] for nonconvex problems. The DR splitting method is essentially an implicit and unconditionally stable splitting algorithm, which means the DR splitting method is convergent with arbitrary positive step sizes. Solving implicit equations including linear equations can be computationally demanding for DR methods. In [10], within the generalized proximal-point framework, an efficient preconditioned DR framework with a convergence guarantee aiming at a specific class of convex-concave saddle-point problems is developed. It can incorporate the classical preconditioning techniques for linear systems into the nonlinear DR iterations. We refer to [9] for applications of preconditioned DR and [7] for its accelerated variants.
Due to a wide range of applications in various fields including big data science and machine learning, there are a lot of studies on stochastic first-order algorithms [5, 15] and especially for saddle-point problems [1, 13, 15]. In this paper, we study a stochastic preconditioned DR splitting framework which includes the stochastic DR without preconditioners [15] as a special case for solving the following saddle-point problems with separable dual variables [1, 15]. We refer to [24] for the randomized DR methods aiming at solving linear systems.
$\min_{x\in\operatorname{dom}{\mathcal{F}}}\max_{y\in\operatorname{dom}{\mathcal{G}}}\mathcal{L}(x,y),\quad\mathcal{L}(x,y):=\mathcal{F}(x)+\sum_{i=1}^{n}(\langle\mathcal{K}_{i}x,y_{i}\rangle-\mathcal{G}_{i}(y_{i})).$ (1.1)
where $\mathcal{F}:\mathbb{X}\rightarrow\mathbb{R}_{\infty}:=\mathbb{R}\cup\{+\infty\}$ and each $\mathcal{G}_{i}:\mathbb{Y}_{i}\rightarrow\mathbb{R}_{\infty}$ are proper, convex and lower semi-continuous. Here, $\mathbb{X}$ and $\mathbb{Y}_{i}$ are real Hilbert spaces and each $\mathcal{K}_{i}:\mathbb{X}\rightarrow\mathbb{Y}_{i}$ is linear and continuous. The space $\mathbb{Y}$ is a product space, i.e., $\mathbb{Y}=\prod^{n}_{i=1}\mathbb{Y}_{i}$ with $y=(y_{1},y_{2},\ldots,y_{n})^{T}$. The saddle-point problem (1.1) can be reformulated as the following common primal-dual formulation [4]:
$\min_{x\in\operatorname{dom}{\mathcal{F}}}\max_{y\in\operatorname{dom}{\mathcal{G}}}\mathcal{L}(x,y),\quad\mathcal{L}(x,y):=\mathcal{F}(x)+\langle\mathcal{K}x,y\rangle-\mathcal{G}(y),$ (1.2)
where the dual function $\mathcal{G}$ is separable, i.e., $\mathcal{G}(y)=\sum^{n}_{i=1}\mathcal{G}_{i}(y_{i})$ with $\mathcal{G}:\mathbb{Y}\rightarrow\mathbb{R}_{\infty}$ and $\mathcal{K}=[\mathcal{K}_{1},\mathcal{K}_{2},\ldots,\mathcal{K}_{n}]^{T}:\mathbb{X}\rightarrow\mathbb{Y}$. When strong duality holds [4], the solutions of (1.1) are exactly the solutions of the following primal and dual problems:
$\min_{x\in\mathbb{X}}\mathcal{P}(x),\ \mathcal{P}(x):=\mathcal{F}(x)+\sum_{i=1}^{n}\mathcal{G}_{i}^{*}(\mathcal{K}_{i}x),\quad\max_{y\in\mathbb{Y}}\mathcal{D}(y),\ \mathcal{D}(y):=-\mathcal{F}^{*}(-\mathcal{K}^{*}y)-\sum_{i=1}^{n}\mathcal{G}_{i}(y_{i}),$ (1.3)
with $\mathcal{F}^{*}$ and $\mathcal{G}_{i}^{*}$ denoting the Fenchel conjugate functionals of $\mathcal{F}$ and $\mathcal{G}_{i}$, respectively.
Compared to related work, our contributions are as follows. First, different from the stochastic first-order primal-dual method [13] or stochastic classical DR splitting method without preconditioners [15], preconditioners are introduced. The challenges in the convergence analysis are addressed for the proposed stochastic DR splitting method. Preconditioners already introduced challenges for the deterministic preconditioned DR and extra effort is necessary for the corresponding convergence analysis [10, 7]. Second, we incorporate a general over-relaxed strategy into the stochastic and preconditioned DR splitting method with different relaxation parameters on different variables. It is well known that over-relaxation techniques can accelerate the deterministic DR splitting algorithm [19, 17, 33, 8]. For the over-relaxed, stochastic, and preconditioned DR (hereafter abbreviated as SRPDR), the almost sure convergence framework of the stochastic fixed point iteration based on firmly nonexpansive operators developed in [15] cannot be applied directly here. We circumvent these difficulties by analyzing the relation between the stochastic updates and the one-step deterministic updates based on the current stochastic iteration. We mainly employed the celebrated stochastic quasi-Fejér monotonicity [15, 16, 20] and Opial’s lemma [35]. Third, we show the SRPDR splitting method has a $\mathcal{O}(1/K)$ rate for the ergodic sequence respecting restricted primal-dual gap functions where $K$ is the iteration number. Since taking the expectation of the supremum of the primal-dual gap concerning transitional variables is a very subtle issue [1], a detailed analysis is devoted to this part. Furthermore, we also show the SRPDR splitting method has a $\mathcal{O}(1/K)$ rate for primal error when Lipschitz continuity is satisfied for dual functions. Finally, we show the high efficiency of the proposed SRPDR compared to state-of-art methods including the stochastic primal-dual method [13] with some numerical experiments on synthetic and real datasets.
The rest of the paper is organized as follows. In section 2, we give a brief introduction of the preconditioned deterministic DR and the proposed stochastic, over-relaxed, and preconditioned DR splitting method along with some necessary analysis tools. In section 3, we discuss the almost sure convergence of the proposed SRPDR splitting method to a fixed point using stochastic quasi-Fejér monotonicity [15] and Opial’s lemma [35]. In section 4, we show the sublinear convergence rate of ergodic sequences respecting the expectation of the restricted primal-dual gap and the primal error. In section 5, we extend these results to the stochastic relaxed and preconditioned DR method for quadratic-linear problems (henceforth abbreviated as SRPDRQ). In section 6, we present detailed numerical experiments for total generalized variation (TGV) regularized problems and binary classification with real datasets to show the competitive performance of the proposed stochastic and preconditioned DR methods. Finally, we give some conclusions in section 7.
2 Preconditioned DR framework
Now, let us turn to the deterministic preconditioned Douglas–Rachford splitting method (PDR), which aims at solving the saddle-point problem (1.2). It reads as follows [10, 9, 7]:
$\left\{\begin{array}[]{ll}x^{k+1}=N_{1}^{-1}[(N_{1}-I)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{\ast}y^{k+1}],\\ y^{k+1}=N_{2}^{-1}[(N_{2}-I)y^{k}+\bar{y}^{k}+\tau\mathcal{K}x^{k+1}],\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=(I+\tau\partial\mathcal{G})^{-1}[2y^{k+1}-\bar{y}^{k}],\\ \bar{x}^{k+1}=\bar{x}^{k}+x_{\operatorname{test}}^{k+1}-x^{k+1},\\ \bar{y}^{k+1}=\bar{y}^{k}+y_{\operatorname{test}}^{k+1}-y^{k+1},\end{array}\right.$ (PDR)
where $\sigma$ and $\tau$ are arbitrary positive primal and dual step sizes, $N_{1}\succeq I$, $N_{2}\succeq I$ with $N_{1}:\mathbb{X}\rightarrow\mathbb{X}$ and $N_{2}:\mathbb{Y}\rightarrow\mathbb{Y}$ being bounded, linear and self-adjoint operators, and $x_{\operatorname{test}}^{k+1}$ and $y_{\operatorname{test}}^{k+1}$ are transitional variables. Denoting by $z=(x,y)\in\mathbb{X}\times\mathbb{Y}$ and $\bar{z}=(\bar{x},\bar{y})\in\mathbb{X}\times\mathbb{Y}$, the iteration PDR with $N_{1}=I$, $N_{2}=I$ and $\sigma=\tau$ becomes the original DR splitting method [19, 32]. It was based on the following maximal monotone operator splitting
$0\in Az+Bz,\ \ \text{with}\quad A=\begin{pmatrix}0&\mathcal{K}^{*}\\ -\mathcal{K}&0\end{pmatrix},\quad B=\begin{pmatrix}\partial\mathcal{F}&0\\ 0&\partial\mathcal{G}\end{pmatrix},$
by the first-order optimality conditions of (1.2). Note that the first two equations of PDR are implicit. However, if we choose $N_{2}=I$ and an operator $M$ such that
$N_{1}=M-\sigma\tau\mathcal{K}^{*}\mathcal{K}$ (2.1)
where $M:X\rightarrow X$ is linear, continuous, self-adjoint and satisfies $M\succeq T:=I+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, we can update $x^{k+1}$ first with the following classical preconditioned iteration
$x^{k+1}=x^{k}+M^{-1}(b^{k}-Tx^{k}),\quad b^{k}=\bar{x}^{k}-\sigma\mathcal{K}^{*}\bar{y}^{k},$ (2.2)
and then obtain the linear update $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ according to the second equation of PDR followed by $x^{k+1}$ [10, 9, 7]. Classical preconditioning techniques can be employed with (2.2) which can be seen as an iterative method for solving $Tx^{k+1}=b^{k}$. This linear equation arises from the linear update step in the original Douglas–Rachford splitting method where $N_{1}=I$ and $N_{2}=I$. Preconditioned DR splitting can bring high efficiency with great flexibility and convergence guarantees using finitely many iterations of any feasible preconditioner [10, 9, 7]. A preconditioner is called feasible if $M\succeq T$ as in (2.2) includes a lot of classical efficient preconditioners [10]. Let us denote by $u:=(x,y,\bar{x},\bar{y})\in\mathbb{U}:=(\mathbb{X}\times\mathbb{Y})^{2}$ and introduce an associated positive semidefinite and possibly degenerate inner product as
$\langle u,u^{\prime}\rangle_{\mathbf{M}}:=\langle\mathbf{M}u,u^{\prime}\rangle=\frac{1}{\sigma}[\langle(N_{1}-I)x,x^{\prime}\rangle+\langle\bar{x},\bar{x}^{\prime}\rangle]+\frac{1}{\tau}[\langle(N_{2}-I)y,y^{\prime}\rangle+\langle\bar{y},\bar{y}^{\prime}\rangle],$ (2.3)
with $\mathbf{M}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\sigma}I,\tfrac{1}{\tau}I]$. Denoting by $\mathcal{T}:\mathbb{U}\rightarrow\mathbb{U}$ as the iteration operator that maps $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ to $u^{k+1}=(x^{k+1},y^{k+1},\bar{x}^{k+1},\bar{y}^{k+1})$ according to PDR, we can rewrite the iteration PDR as
$u^{k+1}:=\mathcal{T}u^{k}.$ (2.4)
Now, we present the following lemma and proposition, which are useful for analyzing the stochastic PDR method in the subsequent sections. The proofs for the case where $\sigma=\tau$ can be found in [10, 9], while the proofs for the case where $\sigma\neq\tau$ can be obtained easily.
Lemma 1 ([9, Lemma 2.1]).
For ${u^{k}={(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})}}$, we have
$\frac{1}{\sigma}\langle x^{k+1}-\bar{x}^{k+1},x_{\operatorname{test}}^{k+1}\rangle+\frac{1}{\tau}\langle y^{k+1}-\bar{y}^{k+1},y_{\operatorname{test}}^{k+1}\rangle=\langle u^{k}-u^{k+1},u^{k+1}\rangle_{\mathbf{M}}.$ (2.5)
Hence, employing Lemma 1 yields the following proposition.
Proposition 1 ([9, Lemma 2.2]).
If ${u={(x,y,\bar{x},\bar{y})}}$ with $(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$, and $\bar{y}=y-\tau\mathcal{K}x$, then we have
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})$ $\leq\langle u^{k}-u^{k+1},u^{k+1}-u\rangle_{\mathbf{M}}$
$=\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}}^{2}).$ (2.6)
The equality in (2.6) comes from the polarization identity with positive semi-definite weight.
Next, let us introduce the relaxed preconditioned Douglas-Rachford splitting method,
$u^{k+1}=(I-I_{\varrho_{k}})u^{k}+I_{\varrho_{k}}\mathcal{T}u^{k},$ (2.7)
where $I_{\varrho_{k}}:=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$. Throughout the paper, we assume the dimension of each identity operator $I$ depends on the corresponding variable. Suppose $\rho_{k}$, $\rho_{k,x}$, and $\rho_{k,y}$ are non-decreasing numbers with lower and upper limits $\rho_{l}$ and $\rho_{u}$, respectively, such that $20<\rho_{l}<\rho_{u}<$ 2. With $\mathcal{T}$ being the iteration operator of PDR given by (2.4), we then define $u^{k+1}_{t}=\mathcal{T}u^{k}:=(x^{k+1}_{t},y^{k+1}_{t},\bar{x}^{k+1}_{t},\bar{y}^{k+1}_{t})$. In the computation, we also introduce the following new transitional variables:
$\left\{\begin{array}[]{ll}x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x_{t}^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=(I+\tau\partial\mathcal{G})^{-1}[2y_{t}^{k+1}-\bar{y}^{k}].\end{array}\right.$ (2.8)
Writing (2.7) component-wisely, we get the relaxed and preconditioned Douglas-Rachford iteration as follows
$\left\{\begin{array}[]{ll}x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}x_{t}^{k+1},\\ y^{k+1}=(1-\rho_{k})y^{k}+\rho_{k}y_{t}^{k+1},\\ \bar{x}^{k+1}=\bar{x}^{k}+\rho_{k,x}[x_{\operatorname{test}}^{k+1}-x_{t}^{k+1}],\\ \bar{y}^{k+1}=\bar{y}^{k}+\rho_{k,y}[y_{\operatorname{test}}^{k+1}-y_{t}^{k+1}].\\ \end{array}\right.$ (RPDR)
Now we define the iteration operator RPDR as $\mathcal{T}_{R}$ such that $u^{k+1}:=\mathcal{T}_{R}u^{k}$. Note that PDR is a special case of RPDR when $I_{\varrho_{k}}=I$. Let us define
$\mathbf{M}_{\varrho_{k}}:=\mathbf{M}I^{-1}_{\varrho_{k}}=\operatorname{Diag}\left[\frac{1}{\sigma\rho_{k}}(N_{1}-I),\frac{1}{\tau\rho_{k}}(N_{2}-I),\frac{1}{\sigma\rho_{k,x}}I,\frac{1}{\tau\rho_{k,y}}I\right].$
We will present the following proposition, with the proof provided in the Appendix.
Proposition 2.
If ${u={(x,y,\bar{x},\bar{y})}}$ with $(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ and $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$, $\bar{y}=y-\tau\mathcal{K}x$, then we have
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})\leq\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}}^{2}).$ (2.9)
[…]
Lemma 2.
The iteration RPDR for problem 1.2 possesses the following properties:
  • (i)
    A point $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ is a fixed point of RPDR if and only if $(x^{*},y^{*})$ is a saddle point for (1.2) and $x^{*}+\sigma\mathcal{K}^{*}y^{*}=\bar{x}^{*},y^{*}-\tau\mathcal{K}x^{*}=\bar{y}^{*}$.
  • (ii)
    If w-${\lim_{j\rightarrow\infty}}(\bar{x}^{k_{j}},\bar{y}^{k_{j}})=(\bar{x}^{*},\bar{y}^{*})$ and ${\lim_{j\rightarrow\infty}}((N_{1}-I)(x^{k_{j}}-x^{k_{j}+1}),(N_{2}-I)(y^{k_{j}}-y^{k_{j}+1}),\bar{x}^{k_{j}}-\bar{x}^{k_{j}+1},\bar{y}^{k_{j}}-\bar{y}^{k_{j}+1})=(0,0,0,0)$, then $(x^{k_{j}+1},y^{k_{j}+1},\bar{x}^{k_{j}+1},\bar{y}^{k_{j}+1})$ converges weakly to a fixed point.
Remark 1.
If we choose $I_{\varrho_{k}}:=\textrm{Diag}[I,I,\rho_{k,x}I,\rho_{k,y}I]$ in relaxed preconditioned Douglas-Rachford method (2.7), we get partial relaxed preconditioned Douglas-Rachford method (p-RPDR). Further, if we choose $N_{1}=I$, $N_{2}=I$, and $\rho_{k,x}=\rho_{k,y}$, we derive the original relaxed Douglas-Rachford splitting that was studied in [19], of which the weak convergence can be seen in [19, Theorem 7] with an ergodic convergence rate.
[…]
To clarify, we refer to the general RPDR as the fully relaxed preconditioned Douglas-Rachford method (f-RPDR) when $I_{\varrho_{k}}:=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k}\neq 1$, $\rho_{k,x}\neq 1$, and $\rho_{k,y}\neq 1$.
In the special case where $\mathcal{F}$ is a quadratic-linear form as follows:
$\mathcal{F}(x)=\langle\frac{1}{2}Qx-f,x\rangle,$ (2.10)
where $Q$ is a self-adjoint, linear, and bounded positive semi-definite operator and $f\in X$, we can get a more compact preconditioned Douglas–Rachford splitting method for quadratic problems, which is called PDRQ [10, 9, 7]. Denoting by $u=(x,y,\bar{y})\in\mathbb{U}_{Q}:=\mathbb{X}\times\mathbb{Y}\times\mathbb{Y}$, $N_{1}\succeq 0$, $N_{2}\succeq I$, the PDRQ iteration reads as,
$\left\{\begin{array}[]{l}x^{k+1}=(N_{1}+\sigma Q)^{-1}[N_{1}x^{k}-\sigma(\mathcal{K}^{*}y^{k+1}-f)],\\ y^{k+1}=N_{2}^{-1}[(N_{2}-I)y^{k}+\bar{y}^{k}+\tau\mathcal{K}x^{k+1}],\\ y_{\operatorname{test}}^{k+1}=(I+\tau\partial G)^{-1}[2y^{k+1}-\bar{y}^{k}],\\ \bar{y}^{k+1}=\bar{y}^{k}+y_{\operatorname{test}}^{k+1}-y^{k+1}.\end{array}\right.$ (PDRQ)
Without the updates for $\bar{x}^{k+1}$, the PDRQ iterations are more compact and efficient compared to PDR [10]. Moreover, if we choose $N_{2}=I$ and $N_{1}$ such that
$N_{1}=M_{Q}-\sigma Q-\sigma\tau\mathcal{K}^{*}\mathcal{K},$ (2.11)
where $M_{Q}:X\rightarrow X$ is linear, continuous, self-adjoint and feasible meaning $M_{Q}\succeq T_{Q}:=\sigma Q+\sigma\tau\mathcal{K}^{*}\mathcal{K}$, then we can update $x^{k+1}$ first with the following classical preconditioned iteration
$x^{k+1}=x^{k}+M_{Q}^{-1}(b^{k}-Tx^{k}),\quad b^{k}=\sigma f-\sigma\mathcal{K}^{*}\bar{y}^{k}.$ (2.12)
The linear update of $y^{k+1}=\bar{y}^{k}+\tau\mathcal{K}x^{k+1}$ is then according to the second line in PDRQ [10, 9, 7]. Denoting
$\mathbf{M}_{Q}=\operatorname{Diag}[\tfrac{1}{\sigma}N_{1},\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\tau}I],$
we have the associated inner product again,
$\langle u,u^{\prime}\rangle_{\mathbf{M}_{Q}}=\frac{1}{\sigma}\langle N_{1}x,x^{\prime}\rangle+\frac{1}{\tau}[\langle(N_{2}-I)y,y^{\prime}\rangle+\langle\bar{y},\bar{y}^{\prime}\rangle].$ (2.13)
Let $\mathcal{T}_{Q}$ denote the iteration operator associated with PDRQ, we can rewrite the iteration PDRQ as
$u^{k+1}:=\mathcal{T}_{Q}u^{k}.$ (2.14)
Now, we introduce the relaxed PDRQ method,
$u^{k+1}=(I-I_{q,\varrho_{k}})u^{k}+I_{q,\varrho_{k}}\mathcal{T}_{Q}u^{k},$ (2.15)
where $I_{q,\varrho_{k}}:=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,y}I]$. Suppose $\rho_{k}$, $\rho_{k,y}$ are non-decreasing and have lower and upper limits $\rho_{l},\rho_{u}$, and $20<\rho_{l}<\rho_{u}<$ 2. We also set $u^{k+1}_{t}=\mathcal{T}_{{Q}}u^{k}:=(x^{k+1}_{t},y^{k+1}_{t},\bar{y}^{k+1}_{t})$ and
$y_{\operatorname{test}}^{k+1}=(I+\tau\partial\mathcal{G})^{-1}[2y_{t}^{k+1}-y^{k}],$
Writing (2.15) component-wisely, we get the relaxed and preconditioned Douglas-Rachford iteration for the quadratic problem as follows
$\left\{\begin{array}[]{ll}x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}x_{t}^{k+1},\\ y^{k+1}=(1-\rho_{k})y^{k}+\rho_{k}y_{t}^{k+1},\\ \bar{y}^{k+1}=\bar{y}^{k}+\rho_{k,y}[y_{\operatorname{test}}^{k+1}-y_{t}^{k+1}].\\ \end{array}\right.$ (RPDRQ)
We now define the iteration operator RPDRQ as $\mathcal{T}_{RQ}$ such that $u^{k+1}:=\mathcal{T}_{RQ}u^{k}$. Let $\mathbf{M}_{Q\varrho_{k}}:=\mathbf{M}_{Q}I^{-1}_{q,\varrho_{k}}=\operatorname{Diag}[\frac{1}{\sigma\rho_{k}}N_{1},\frac{1}{\tau\rho_{k}}(N_{2}-I),\tfrac{1}{\tau\rho_{k,y}}I]$. Similarly, we have the following proposition and lemma for the iteration sequence $\{u^{k}\}_{k}$ produced by RPDRQ.
Proposition 3.
If ${u={(x,y,\bar{y})}}$, with $\bar{y}=y-\tau\mathcal{K}x$, then we have the following estimate:
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})\leq\frac{1}{2}(\|u^{k}-u\|_{\mathbf{M}_{Q\varrho_{k}}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}_{Q\varrho_{k}}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}_{Q}(2I-I_{q,\varrho_{k}})I^{-2}_{q,\varrho_{k}}}^{2}).$ (2.16)
Lemma 3.
The iteration RPDRQ for 1.2 with $\mathcal{F}$ as in (2.10) possesses the following properties:
  • (i)
    A point $u^{*}=(x^{*},y^{*},\bar{y}^{*})$ is a fixed point of (2.14) if and only if $(x^{*},y^{*})$ is a saddle point for (1.2) and $Qx^{*}+\mathcal{K}^{*}y^{*}=f$, $y^{*}-\tau\mathcal{K}x^{*}=\bar{y}^{*}$.
  • (ii)
    If w-${\lim_{j\rightarrow\infty}}\bar{y}^{k_{j}}=\bar{y}^{*}$ and ${\lim_{j\rightarrow\infty}}(N_{1}(x^{k_{j}}-x^{k_{j}+1}),(N_{2}-I)(y^{k_{j}}-y^{k_{j}+1}),\bar{y}^{k_{j}}-\bar{y}^{k_{j}+1})=(0,0,0)$, then $(x^{k_{j}+1},y^{k_{j}+1},\bar{y}^{k_{j}+1})$ converges weakly to a fixed point.
Now, we introduce the restricted primal-dual gap function and the Bregman distance for the gap estimates. Suppose that the bounded set $\mathbb{B}:=\mathbb{B}_{1}\times\mathbb{B}_{2}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ contains a saddle point. We give the following restricted primal-dual gap $\mathfrak{G}_{\mathbb{B}_{1}\times\mathbb{B}_{2}}$ associated with $\mathbb{B}$:
$\mathfrak{G}_{\mathbb{B}_{1}\times\mathbb{B}_{2}}(x,y):=\sup_{\tilde{y}\in\mathbb{B}_{2}}\mathcal{L}(x,\tilde{y})-\inf_{\tilde{x}\in\mathbb{B}_{1}}\mathcal{L}(\tilde{x},y):=\sup_{\tilde{z}\in\mathbb{B}}\mathcal{H}(z,\tilde{z}),\quad\mathcal{H}(z,\tilde{z}):=\mathcal{L}(x,\tilde{y})-\mathcal{L}(\tilde{x},y),$ (2.17)
where $z=(x,y)$ as before and $\tilde{z}:=(\tilde{x},\tilde{y})$. For any proper, convex, and lower semi-continuous function $\zeta:\mathbb{X}\rightarrow\mathbb{R}_{\infty}$, the Bregman distance at $s$ with $q\in\partial\zeta(s)$ is defined as
$D_{\zeta}^{q}(t,s)=\zeta(t)-\zeta(s)-\langle q,t-s\rangle.$
If $\tilde{z}$ is a saddle point, we can interpret $\mathcal{H}(z,\tilde{z})$ as the following Bregman distance for $h(z):=\mathcal{F}(x)+\mathcal{G}(y)$ with respect to the point $q:=(-\mathcal{K}^{*}\tilde{y},\mathcal{K}\tilde{x})$ which is included in the subgradient $\partial h(\tilde{z})$ by optimality, i.e.,
$\mathcal{H}(z,\tilde{z}):=D_{h}^{q}(z,\tilde{z})=D_{\mathcal{F}}^{-\mathcal{K}^{*}\tilde{y}}(x,\tilde{x})+D_{\mathcal{G}}^{\mathcal{K}\tilde{x}}(y,\tilde{y}).$
The following theorem tells us that over-relaxation can bring certain acceleration for the deterministic PDR method and motivates us to focus on the stochastic and over-relaxed PDR. The proof is given in the Appendix.
Theorem 1.
If $u=(x,y,\bar{x},\bar{y}),$ with $\bar{x}=x+\sigma\mathcal{K}^{\ast}y,$ $\bar{y}=y-\tau\mathcal{K}x$, and $(x,y)$ being a saddle-point of (1.2). Let $u^{k}$ be the iteration sequence generated by RPDR. Denoting $K_{1}=K+1$, let $z^{k}_{\operatorname{test}}=(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ be the transitional variables of RPDR, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K+1}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in a restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, a bounded domain and containing a saddle-point, it holds that
$\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)\leq\frac{1}{2K_{1}\rho_{l}}\sup_{(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{1}}\|u^{0}-u\|_{\mathbf{M}}^{2}=\mathcal{O}(\frac{1}{K}).$ (2.18)
With these preparations for the deterministic relaxed PDR method, i.e., RPDR (or RPDRQ), let us now shift our focus to the stochastic and relaxed PDR method (or stochastic and relaxed PDRQ method).
2.1 The stochastic relaxed preconditioned DR (SRPDR) method
We begin by introducing the stochastic setting. Let $(\Omega,\mathfrak{F},\mathcal{P})$ denote the underlying probability space. Denote $\{i_{0},i_{1},i_{2},\ldots\}$ as index selection variables, i.e., independent random variables on the latter space with values in $\{1,\ldots,n\}$ which are identically distributed. We denote the probability of selecting an index $i\in\{1,\ldots,n\}$ by $p_{i}$ where we assume $10<p_{i}<$ 1 for each $i$ and recall that $\sum_{i=1}^{n}p_{i}=1$. Introduce $P=\operatorname{Diag}[p_{1},p_{2},\ldots,p_{n}]$ and observe that $P^{-1}=\operatorname{Diag}[\frac{1}{p_{1}},\frac{1}{p_{2}},\ldots,\frac{1}{p_{n}}]$. A Borel random vector $y$ in a separable Hilbert space $\mathbb{H}$ is a measurable map from $\Omega$ into $\mathbb{H}$ where $\Omega$ is equipped with $\mathfrak{F}$ and $\mathbb{H}$ is equipped with the Borel $\sigma$-algebra, i.e., the $\sigma$-algebra generated by the open sets of $\mathbb{H}$ (see [15] or [30, Chapter 2]). Further, $(\mathfrak{F}_{k})_{k\in\mathbb{N}}$ denotes the filtration generated by the indices $\{i_{0},i_{1},\ldots,i_{k}\}$, i.e., the $\sigma$-algebras generated by $i_{j}^{-1}\{i\}$ for $j=0,\ldots,k$ and $i=1,\ldots,n$ [40, Chapter 7]. Then, $\mathfrak{F}_{k}\subset\mathfrak{F}_{k+1}$ for each $k\in\mathbb{N}$. Henceforth, we denote $\mathbb{E}^{k}:=\mathbb{E}[\cdot\mid\mathfrak{F}_{k}]$ as the conditional expectation with respect to $\mathfrak{F}_{k}$, and by $\mathbb{E}$ the full expectation. Recall that if a Borel random vector $y$ is $\mathfrak{F}_{k}$-measurable, i.e., the preimages of Borel sets are in $\mathfrak{F}_{k}$, then $\mathbb{E}[y\mid\mathfrak{F}_{k}]=y$.
Let us first extend the relaxed preconditioned Douglas–Rachford splitting method, i.e., RPDR to a stochastic setting. For arbitrary positive step sizes $\sigma$, $\tau$ and $(x^{0},y^{0},\bar{x}^{0},\bar{y}^{0})\in(\mathbb{B}_{1}\times\mathbb{B}_{2})^{2}$, $y^{0}_{\operatorname{test}}\in\mathbb{B}_{2}$, the stochastic relaxed PDR (SRPDR) reads as follows
$\left\{\begin{array}[]{ll}\tilde{x}_{t}^{k+1}=\left(N_{1}+\sigma\tau\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(\left(N_{1}-I\right)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right),\\ x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}\tilde{x}_{t}^{k+1},\\ y^{k+1}=\left\{\begin{array}[]{ll}(1-\rho_{k})y_{i}^{k}+\rho_{k}\tilde{y}^{k+1}_{t,i},\quad\tilde{y}^{k+1}_{t,i}:=N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}\tilde{x}_{t}^{k+1}]&i=i_{k},\\ y^{k}_{i}&\forall i\neq i_{k},\\ \end{array}\right.\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2\tilde{x}_{t}^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2\tilde{y}^{k+1}_{t,i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{x}^{k+1}=\bar{x}^{k}+\rho_{k,x}(x_{\operatorname{test}}^{k+1}-\tilde{x}_{t}^{k+1}),\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+\rho_{k,y}(y_{\operatorname{test},i}^{k+1}-\tilde{y}_{t,i}^{k+1})&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k},\end{array}\right.\end{array}\right.$ (SRPDR)
where $\tilde{u}^{k+1}_{t}=\mathcal{T}u^{k}:=(\tilde{x}^{k+1}_{t},\tilde{y}^{k+1}_{t},\tilde{\bar{x}}^{k+1}_{t},\tilde{\bar{y}}^{k+1}_{t})$ and $\mathcal{T}$ is the iteration operator of PDR. $N_{1}\succeq I$ and $N_{2}=\operatorname{Diag}[N_{2,1},\ldots,N_{2,n}]$ with each $N_{2,i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ being self-adjoint, bounded and linear such that $N_{2,i}\succeq I$. Again, let $I_{\varrho_{k}}:=\text{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k}$, $\rho_{k,x}$, and $\rho_{k,y}$ are non-decreasing numbers with lower and upper limits $\rho_{l}$ and $\rho_{u}$, respectively, such that $20<\rho_{l}<\rho_{u}<$ 2. Although the updates of $x^{k+1}$ in the above SRPDR method seem complicated, it is convenient for analysis. Specifically, the updates of $\tilde{x}_{t}^{k+1}$ is based on preconditioned iteration given in (2.2) when $N_{2}=I$, and $N_{1}$ as in (2.1).
To clarify, we refer to the general SRPDR as the stochastic fully relaxed preconditioned Douglas-Rachford method (f-SRPDR) when $I_{\varrho_{k}}=\textrm{Diag}[\rho_{k}I,\rho_{k}I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k}\neq 1$, $\rho_{k,x}\neq 1$, and $\rho_{k,y}\neq 1$. While when $I_{\varrho_{k}}=\textrm{Diag}[I,I,\rho_{k,x}I,\rho_{k,y}I]$, where $\rho_{k,x}\neq 1$, and $\rho_{k,y}\neq 1$, we call it stochastic partial relaxed preconditioned Douglas-Rachford method (p-SRPDR). In addition, by letting $I_{\varrho_{k}}=I$, the stochastic PDR, i.e., SPDR reads as follows
$\left\{\begin{array}[]{ll}x^{k+1}=\left(N_{1}+\sigma\tau\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(\left(N_{1}-I\right)x^{k}+\bar{x}^{k}-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right),\\ y^{k+1}=\left\{\begin{array}[]{ll}N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}x^{k+1}]&i=i_{k},\\ y^{k}_{i}&\forall i\neq i_{k},\\ \end{array}\right.\\ x_{\operatorname{test}}^{k+1}=(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}],\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2y^{k+1}_{i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{x}^{k+1}=\bar{x}^{k}+(I+\sigma\partial\mathcal{F})^{-1}[2x^{k+1}-\bar{x}^{k}]-x^{k+1},\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+y_{\operatorname{test},i}^{k+1}-y_{i}^{k+1}&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k}.\end{array}\right.\end{array}\right.$ (SPDR)
Considering a simpler case when $N_{1}=I$ and $N_{2}=I$, we will derive the stochastic DR method introduced in [15].
With these preparations, now let us turn to our algorithms’ almost sure convergence and restricted gap estimates. Since we can treat SPDR as a special case of SRPDR, our analysis will focus solely on the convergence properties and gap estimates for SRPDR.
3 Almost sure convergence of stochastic RPDR
This section presents the almost sure convergence of iteration sequences generated by SRPDR to a solution of (1.2). To avoid any ambiguity, henceforth, the notation $u^{k}=(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$ along with $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ always represents the stochastic sequences produced by SRPDR. To write the iteration in a more compact form, we introduce the notation $\tilde{u}^{k+1}$ for one step RPDR update based on $u^{k}$, i.e.,
$\tilde{u}^{k+1}:=\mathcal{T}_{R}u^{k},\quad\text{with}\ \ \tilde{u}^{k+1}=(\tilde{x}^{k+1},\tilde{y}^{k+1},\tilde{\bar{x}}^{k+1},\tilde{\bar{y}}^{k+1}),$ (3.1)
where $\mathcal{T}_{R}$ is the iteration operator of RPDR, and let $(\tilde{x}^{k+1}_{\operatorname{test}},\tilde{y}^{k+1}_{\operatorname{test}})$ be the transitional variables of this one step RPDR update. Further, we introduce the notation $\tilde{u}_{t}^{k+1}$ for one step deterministic PDR update based on $u^{k}$, i.e.,
$\tilde{u}_{t}^{k+1}:=\mathcal{T}u^{k},\quad\text{with}\ \ \tilde{u}_{t}^{k+1}=(\tilde{x}_{t}^{k+1},\tilde{y}_{t}^{k+1},\tilde{\bar{x}}_{t}^{k+1},\tilde{\bar{y}}_{t}^{k+1}),$ (3.2)
where $\mathcal{T}$ is the same as in (2.4), and let $(\tilde{x}^{k+1}_{t,\operatorname{test}},\tilde{y}^{k+1}_{t,\operatorname{test}})$ be the transitional variables of this one step PDR update. Hence the stochastic update $u^{k+1}$ in SRPDR can then be written as
$x^{k+1}=\tilde{x}^{k+1},\quad x_{\operatorname{test}}^{k+1}=\tilde{x}_{\operatorname{test}}^{k+1}=\tilde{x}_{t,\operatorname{test}}^{k+1},\quad\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1},$ (3.3)
and
$y^{k+1}_{i}=\begin{cases}\tilde{y}^{k+1}_{i}&\text{if}\ i=i_{k},\\ y^{k}_{i}&\text{if}\ i\neq i_{k},\end{cases}\quad y^{k+1}_{\operatorname{test},i}=\begin{cases}\tilde{y}^{k+1}_{\operatorname{test},i}={\tilde{y}^{k+1}_{t,\operatorname{test},i}}&\text{if}\ i=i_{k},\\ y^{k}_{\operatorname{test},i}&\text{if}\ i\neq i_{k},\end{cases}\quad\bar{y}^{k+1}_{i}=\begin{cases}\tilde{\bar{y}}^{k+1}_{i}&\text{if}\ i=i_{k},\\ \bar{y}^{k}_{i}&\text{if}\ i\neq i_{k}.\end{cases}$ (3.4)
Note that $x^{k},x^{k}_{\operatorname{test}},\bar{x}^{k},y^{k},y^{k}_{\operatorname{test}},\bar{y}^{k}$ according to (3.3) and (3.4) are $\mathfrak{F}_{k}$-measurable Borel random vectors by continuity of $\mathcal{T}_{R}$ and construction.
The stochastic iteration sequence $\{u^{k}\}_{k}$ and the corresponding one-step deterministic update $\{\tilde{u}^{k+1}\}_{k+1}$ can be shown in the following diagram
[…]
$\downarrow\ \;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\;\quad\downarrow\;\;\;\qquad\qquad\downarrow{\mathcal{T}_{R}}$
$\tilde{u}^{1}\,\ \quad\tilde{u}^{2}\,\ \quad\tilde{u}^{3}\,\ \quad\tilde{u}^{4}\,\ \quad\cdots\,\ \quad\tilde{u}^{k+1}=\mathcal{T}_{R}u^{k}\cdots.$
Let us begin the convergence analysis with the following estimate.
Lemma 4.
For any $u=(x,y,\bar{x},\bar{y})$ with $(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$ and $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$ and $\bar{y}=y-\tau\mathcal{K}x$, it holds that
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\left(\|u^{k}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}\right),$ (3.6)
where $\mathbf{M}_{p}:=\operatorname{Diag}[\frac{1}{\sigma}(N_{1}-I),\frac{1}{\tau}(N_{2}-I)\mathbf{P}^{-1},\frac{1}{\sigma}I,\frac{1}{\tau}\mathbf{P}^{-1}]$, and $\mathbf{M}_{p,\varrho_{k}}:=I^{-1}_{\varrho_{k}}\mathbf{M}_{p}$. Here $\mathbf{P}=\operatorname{Diag}[p_{1}I,\ldots,p_{n}I]$, consequently, $\mathbf{P}^{-1}=\operatorname{Diag}[p_{1}^{-1}I,\ldots,p_{n}^{-1}I]$.
Proof.
First, noting that for $\mathfrak{F}_{k}$-measurable Borel random vectors $s^{k}=(s^{k}_{i})_{i=1,\ldots,n}$, $\tilde{s}^{k+1}=(\tilde{s}^{k+1}_{i})_{i=1,\ldots,n}$ on $\mathbb{Y}$, the update rule
$s_{i}^{k+1}=\left\{\begin{array}[]{ll}\tilde{s}_{i}^{k+1}&i=i_{k},\\ s_{i}^{k}&\forall i\neq i_{k},\end{array}\right.$
and any continuous $\varphi:Y\to\mathbb{R}$, it holds that
$\mathbb{E}^{k}\varphi(s_{i}^{k+1})=p_{i}\varphi(\tilde{s}_{i}^{k+1})+(1-p_{i})\varphi(s_{i}^{k}),$ (3.7)
which can be rewritten to
$\varphi(\tilde{s}_{i}^{k+1})=\frac{1}{p_{i}}\mathbb{E}^{k}\varphi(s_{i}^{k+1})-(\frac{1}{p_{i}}-1)\varphi(s_{i}^{k}).$ (3.8)
In particular, for any block diagonal operator $\mathbf{B}=\operatorname{Diag}[\mathbf{B}_{1},\ldots,\mathbf{B}_{n}]$ of symmetric and positive semidefinite operators $\mathbf{B}_{1},\ldots,\mathbf{B}_{n}$, this implies that
$\|\tilde{s}^{k+1}-\cdot\|_{\mathbf{B}}^{2}-\|s^{k}-\cdot\|_{\mathbf{B}}^{2}=\mathbb{E}^{k}\|s^{k+1}-\cdot\|_{\mathbf{P}^{-1}\mathbf{B}}^{2}-\|s^{k}-\cdot\|_{\mathbf{P}^{-1}\mathbf{B}}^{2}.$ (3.9)
Then with Proposition 2, $u=(x,y,\bar{x},\bar{y})$ and $\tilde{u}^{k+1}$ as in (3.2), (3.3) and (3.4), we arrive at
$\mathcal{H}\bigl{(}(x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y)\bigr{)}$ $\leq\frac{1}{2}\big{(}\|u^{k}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|\tilde{u}^{k+1}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|\tilde{u}^{k+1}-u^{k}\|_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}^{2}\big{)}$ (3.10)
Next, reformulating $\|\tilde{y}^{k+1}-\cdot\|_{(N_{2}-I)\rho^{-1}_{k}\tau^{-1}}^{2}$ and $\|\tilde{\bar{y}}^{k+1}-\cdot\|_{\rho^{-1}_{k,y}\tau^{-1}I}^{2}$ in the above inequality by using (3.9), then together with the fact that $x^{k+1}$ and $\bar{x}^{k+1}$ are $\mathfrak{F}_{k}$ measurable, we get
$\|\tilde{u}^{k+1}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}-\|u^{k}-u\|_{\mathbf{M}_{\varrho_{k}}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}-\|u^{k}-u\|_{\mathbf{M}_{p,\varrho_{k}}}^{2},$
$\|\tilde{u}^{k+1}-u^{k}\|_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}^{2}$ $=\mathbb{E}^{k}\|u^{k+1}-u^{k}\|_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}^{2},$
and consequently, (3.6). ∎
Moreover, we see, for instance,
$(N_{2}-I)\mathbf{P}^{-1}=(p_{1}^{-1}(N_{2,1}-I),\ldots,p_{n}^{-1}(N_{2,n}-I))=\mathbf{P}^{-1}(N_{2}-I).$
For this reason, we will tacitly use that $\mathbf{P}$ (and $\mathbf{P}^{-1}$) always commutes with the linear operators considered in this paper unless otherwise specified.
Let us give some notations first. Denote $\mathcal{Z}^{*}$ as the set for the saddle points of (1.1) and $\mathbb{F}\subset\mathbb{U}$ as the set for the fixed points for $\mathcal{T}_{R}$. Now, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$ to a fixed point of $\mathcal{T}_{R}$ as $k\to\infty$. Inspired by [15], we mainly employ the stochastic quasi-Fejér monotonicity [15], Opial’s lemma [3, Lemma 21] or [35], and Lemma 2 to show the almost sure convergence of sequence $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ first. We then “glue” $\{u^{k+1},k\in\mathbb{N}\}$ to $\{\tilde{u}^{k+1},k\in\mathbb{N}\}$ and finish the proof.
Theorem 2.
The iteration sequence $\{u^{k}\}_{k}=\{(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})\}_{k}$ produced by SRPDR weakly converges to an $\mathbb{F}$-valued random variable $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ almost surely, and $(x^{**},y^{**})$ is a $\mathcal{Z}^{*}$-valued random variable with $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**}$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover, the transitional sequence $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ converges weakly to a $\mathcal{Z}^{*}$-valued random variable $(x^{**},y^{**})$ almost surely.
Proof.
Inserting $u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})$ a fixed point of $\mathcal{T}_{R}$ in (3.6). By Lemma 2, $(x^{*},y^{*})$ is a saddle point of (1.1). Then it follows for every $(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}$, $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),(x^{*},y^{*}))\geq 0$. Employing Lemma 4 as well as $\mathbf{M}_{p,\varrho_{k+1}}\preceq\mathbf{M}_{p,\varrho_{k}}$ we have
$\mathbb{E}^{k}\|u^{k+1}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k+1}}}$ $\leq\mathbb{E}^{k}\|u^{k+1}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}$
$\leq\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}},$
denoting $\Delta^{k}=\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}$, and due to $\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}=\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}$, we get,
$\mathbb{E}(\Delta^{k+1}|\mathfrak{F}_{k})=\mathbb{E}^{k}\Delta^{k+1}$ $\leq\Delta^{k}-\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}.$ (3.11)
Using the Robbins–Siegmund theorem [38] (see also [15, 1]) on (3.11), we conclude that $\sum_{k=0}^{\infty}\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}}<\infty$. Due to $\rho_{k},\rho_{k,x},\rho_{k,y}$ are non-decreasing, thus $\frac{2-\rho_{k}}{\rho_{k}^{2}}$, $\frac{2-\rho_{k,x}}{\rho_{k,x}^{2}}$ and $\frac{2-\rho_{k,y}}{\rho_{k,y}^{2}}$ are decreasing and the lower bound is $\frac{2-\rho_{u}}{\rho_{u}^{2}}$. Then it follows that
$\sum_{k=0}^{\infty}\frac{2-\rho_{u}}{\rho_{u}^{2}}\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{\mathbf{M}}<\infty,\quad\|\mathcal{T}_{R}u^{k}-u^{k}\|^{2}_{\mathbf{M}}\rightarrow 0,$ (3.12)
almost surely in $\Omega$. Recalling the definition of $\mathbf{M}_{p}$, we immediately see that also $\|\mathcal{T}_{R}u^{k}-u^{k}\|_{\mathbf{M}_{p}}\to 0$ as $k\to\infty$ almost surely. Still with the Robbins–Siegmund theorem or Proposition 2.3 in [15], we conclude that $\Delta^{k}=\|u^{k}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}$ converges to a finite random variable almost surely. Moreover, $\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}$ also converges almost surely by the non-decreasing and convergence of $\rho_{k},\rho_{k,x},\rho_{k,y}$. Using further that $\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}=\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|$ for each $k$, we conclude that $\{\mathbf{M}_{p}^{1/2}u^{k}\}_{k}$ is also bounded almost surely and that the limit
$l(u^{*}):=\lim_{k\rightarrow+\infty}\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}$ (3.13)
exists almost surely. The existence and boudnedness of $\mathbf{M}_{p}^{1/2}$ can be guaranteed by the positive semi-definiteness of $N_{1}-I$ and $N_{2}-I$ and the positive definiteness of $\mathbf{P}$. In the following, we will establish the almost sure weak convergence of $\{u^{k}\}_{k}$.
Our first goal is to show that $\{\tilde{u}^{k}\}_{k}$ converges weakly to an $\mathbb{F}$-valued random variable, i.e., a fixed point of $\mathcal{T}_{R}$. We will begin with proving the almost sure weak convergence of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ with Opial’s lemma [3, Lemma 21]. This can be done by showing that almost surely, $\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|$ converges and every weak sequential limit point of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k+1}\}_{k}$ is a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable. Here $\mathbf{M}_{p}^{1/2}\mathbb{F}$ is defined as follows
$\mathbf{M}_{p}^{1/2}\mathbb{F}:=\{((N_{1}-I)^{\frac{1}{2}}x^{*}/\sigma^{\frac{1}{2}},(N_{2}-I)^{\frac{1}{2}}\mathbf{P}^{-\frac{1}{2}}y^{*}/\tau^{\frac{1}{2}},\bar{x}^{*}/\sigma^{\frac{1}{2}},\mathbf{P}^{-\frac{1}{2}}\bar{y}^{*}/\tau^{\frac{1}{2}})\ |\ u^{*}=(x^{*},y^{*},\bar{x}^{*},\bar{y}^{*})\in\mathbb{F}\},$
with $\mathbf{P}^{-\frac{1}{2}}=\operatorname{Diag}[p_{1}^{-\frac{1}{2}}I,\ldots,p_{n}^{-\frac{1}{2}}I]$. According to the triangle inequality, (3.12), we have
$\big{|}\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|-\|\mathbf{M}_{p}^{1/2}(u^{k}-u^{*})\|\big{|}\leq\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{k})\|=\|\mathcal{T}_{R}u^{k}-u^{k}\|_{\mathbf{M}_{p}}\rightarrow 0.$
Combining the fact of (3.13), we derive,
$\|\mathbf{M}_{p}^{1/2}(\tilde{u}^{k+1}-u^{*})\|\rightarrow l(u^{*})$ (3.14)
almost surely as $k\to\infty$. In particular, $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ is bounded almost surely. Next, we show that for each event such that (3.12), (3.13) and (3.14) hold, each weak accumulation point of $\{\tilde{u}^{k}\}_{k}$ is a fixed point. For that purpose, note that (3.14), we have for the sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ and each $\mathbf{M}_{p}^{1/2}u^{*}$ being a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable that $\|\mathbf{M}_{p}^{1/2}\tilde{u}^{k}-\mathbf{M}_{p}^{1/2}u^{*}\|$ converges. Now, let $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ be a weakly converging subsequence. Then, as $\|\mathbf{M}_{p}^{1/2}(\mathcal{T}_{R}u^{k}-u^{k})\|\to 0$ as $k\to\infty$, $\{\mathbf{M}_{p}^{1/2}u^{k_{j}-1}\}_{j}$ converges weakly to the same limit and looking at the definition of $\mathbf{M}_{p}$, we conclude that $\bigl{\{}\bigl{(}(N_{1}-I)x^{k_{j}-1},(N_{2}-I)y^{k_{j}-1},\bar{x}^{k_{j}-1},\bar{y}^{k_{j}-1}\bigr{)}\bigr{\}}_{j}$ converges weakly. We further have that $\bigl{(}(N_{1}-I)(\tilde{x}^{k_{j}}-x^{k_{j}-1}),(N_{2}-I)(\tilde{y}^{k_{j}}-y^{k_{j}-1}),\tilde{\bar{x}}^{k_{j}}-\bar{x}^{k_{j}-1},\tilde{\bar{y}}^{k_{j}}-\bar{y}^{k_{j}-1}\bigr{)}\to 0$ as $j\to\infty$ such that with Lemma 2 (ii), we obtain that $\{(\tilde{x}^{k_{j}},\tilde{y}^{k_{j}},\tilde{\bar{x}}^{k_{j}},\tilde{\bar{y}}^{k_{j}})\}$ converges weakly to a fixed point $(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}_{R}$. In particular, by Lemma 2 (i), $(x^{**},y^{**})$ is a saddle point for (1.1) and $x^{**}+\sigma\mathcal{K}^{*}y^{**}=\bar{x}^{**},y^{**}-\tau\mathcal{K}x^{**}=\bar{y}^{**}$. It is then immediate that the weak limit of $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k_{j}}\}_{j}$ is a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable. By Opial’s lemma, we have that the whole sequence $\{\mathbf{M}_{p}^{1/2}\tilde{u}^{k}\}_{k}$ converges weakly to a $(\mathbf{M}_{p}^{1/2}\mathbb{F})$-valued random variable, implying that $(\tilde{\bar{x}}^{k},\tilde{\bar{y}}^{k})$ converges weakly to $(\bar{x}^{**},\bar{y}^{**})$. Then by (3.12) again, we have $\bigl{(}(N_{1}-I)(\tilde{x}^{k}-x^{k-1}),(N_{2}-I)(\tilde{y}^{k}-y^{k-1}),\tilde{\bar{x}}^{k}-\bar{x}^{k-1},\tilde{\bar{y}}^{k}-\bar{y}^{k-1}\bigr{)}\to 0$ as $k\to\infty$. Applying Lemma 2 (ii) again, we conclude that $\{\tilde{u}^{k}\}_{k}$ converges weakly to a $\mathbb{F}$-valued random variable.
Finally, let us turn to the almost sure weak convergence of $\{u^{k}\}_{k}$. First, note that since for fixed $i=1,\ldots,n$, the probability $p_{i}$ of $i_{k}=i$ is positive and all $i_{k}$ are independent random variables, the probability of $i_{k}=i$ only occurring finitely many times has to be zero. For that reason, almost surely, for each $i=1,\ldots,n$, the sequences $\{k_{i,j}\}_{j}$ such that $k_{i,j}=\max\ \{k:k\leq j,\ i_{k-1}=i\}$, i.e., the sequences that yield the index of the last activation of $i$, are increasing and satisfy $\lim_{j\to\infty}k_{i,j}=\infty$. Fix an event for which $\{\tilde{u}^{k}\}_{k}$ converges weakly to a fixed point $u^{**}=(x^{**},y^{**},\bar{x}^{**},\bar{y}^{**})$ of $\mathcal{T}_{R}$ and for which $k_{i,j}$ tends to infinity as $j\to\infty$ for all $i=1,\ldots,n$. We will show that $\{u^{k}\}_{k}$ converges weakly, as $k\to\infty$, to $u^{*}$. For that purpose, observe that by (3.3), we have $x^{k}=\tilde{x}^{k}\to x^{**}$, $x^{k}_{\operatorname{test}}=\tilde{x}^{k}+\tilde{\bar{x}}^{k}-\bar{x}^{k-1}\to x^{**}$ and $\bar{x}^{k}=\tilde{\bar{x}}^{k}\to\bar{x}^{**}$ in the weak sense as $k\to\infty$. By (3.4) and the definition of $k_{i,j}$, we see that $y^{j}_{i}=\tilde{y}^{k_{i,j}}_{i}$, $y^{j}_{\operatorname{test},i}=\tilde{y}^{k_{i,j}}_{i}+\tilde{\bar{y}}^{k_{i,j}}_{i}-\bar{y}^{(k_{i,j})-1}_{i}$ and $\bar{y}^{j}_{i}=\tilde{\bar{y}}^{k_{i,j}}_{i}$ for each $j$. Hence, letting $j\to\infty$, we see that $y^{j}_{i}\to y^{**}_{i}$, $y^{j}_{\operatorname{test},i}\to y^{**}_{i}$ and $\bar{y}^{j}_{i}\to\bar{y}^{**}_{i}$ in the weak sense. Since these events occur almost surely, the proof is complete. ∎
Table 1: Existing research and our contributions
Preconditioner setting Almost sure convergence Restricted primal-dual gap
$N_{1}=I;N_{2}=I$ Already have in [15] Presented in this paper
$N_{1}\succ I;N_{2}\succ I$ Easy to prove with [1] Presented in this paper
$N_{1}\succeq I,N_{1}\neq I;N_{2}=I$ Presented in this paper Presented in this paper
As shown in Table 1, our framework covers some results in existing research. The commonly used preconditioners for stochastic (and relaxed) PDR are the cases of $N_{1}\succeq I,N_{1}\neq I$ and $N_{2}=I$, which brings out the degeneracy of $N_{1}-I$ and leads to the difficulties in this proof. For instance, the symmetric Gauss-Seidel preconditioner is widely used for an elliptic operator $T$ as in (2.1) and $M-T=N_{1}-I$ is only positive semidefinite [10, Proposition 2.13]. Next, let us turn to the convergence rate of the restricted primal-dual gap with respect to the transitional variables $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$. It is also a challenge due to the preconditioners and particularly because of the transitional variables, which are the main differences compared to the gap estimate reported in [1]. Since the relations derived from (3.11) are concerning $u^{k}$, i.e., $(x^{k},y^{k},\bar{x}^{k},\bar{y}^{k})$, without any relations of transitional variables $(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$, we introduce Lemma 5, Lemma 6 and Lemma 7 to address this issue.
4 Sublinear convergence of stochastic RPDR for restricted primal-dual gap
Taking full expectation in (3.11), induction with $k$ yields
$\mathbb{E}[\Delta^{k}]\leq\Delta^{0}=\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p,\varrho_{0}}},\ \ \sum_{i=0}^{k}\frac{2-\rho_{u}}{\rho_{u}^{2}}\mathbb{E}[\|u^{i+1}-u^{i}\|^{2}_{\mathbf{M}_{p}}]\leq\sum_{i=0}^{k}\mathbb{E}[\|u^{i+1}-u^{i}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}]\leq\Delta^{0}.$ (4.1)
Since SRPDR involves $u^{k+1}_{t}=\mathcal{T}u^{k}$ in the iteration, let us begin with the following lemma which gives the control of $u^{k+1}_{t}$ by the weighted $\mathbf{M}_{p}$ norm of $u^{k}$, where $u^{k}$ is the iteration sequence of SRPDR. Applying [10, Lemma 2.5] with $\mathbf{M}_{p}$ norm, we have the following estimate.
Lemma 5.
There exists a constant $C>0$ such that for all $u^{1},u^{2}\in\mathbb{U}$ and $\tilde{u}^{1}=\mathcal{T}u^{1}$, $\tilde{u}^{2}=\mathcal{T}u^{2}$ satisfies
$\|\tilde{u}^{1}-\tilde{u}^{2}\|^{2}\leq C\|u^{1}-u^{2}\|_{\mathbf{M}_{p}}^{2},$ (4.2)
where $\mathcal{T}$ is the iteration operator of PDR, the constant $C$ is only related to $N_{1}$, $N_{2}$, $\sigma$, $\tau$ and $\mathcal{K}$.
The following two lemmas can address the main challenges in $y^{k+1}_{\operatorname{test}}$ for the gap estimate.
Lemma 6.
Supposing $\delta\in(0,1)$ is a constant, given a non-decreasing sequence $S_{k}$ and a bounded sequence $h_{k}$ satisfying $0<S_{k}\leq h_{k}+(1-\delta)S_{k-1}$ for $k\geq 1$, we have
$S_{k}\leq\sum_{i=0}^{k-1}(1-\delta)^{i}h_{k-i}+(1-\delta)^{k}S_{0}.$ (4.3)
Proof.
An induction of $S_{k}$ gives (4.3). Furthermore, if $h_{k}$ is bounded by $M$, we have $S_{k}\leq[1-(1-\delta)^{k}]M/{\delta}+(1-\delta)^{k}S_{0}\leq{M}/{\delta}+(1-\delta)^{k}S_{0}$. ∎
Lemma 7.
The estimates of $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ and $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|$ are as follows. First, with notations $s^{k}$ and $\varphi$ defined in (3.7), we deduce
  • (i)
    $\mathbb{E}\varphi(s_{i}^{k+1})=\mathbb{E}[\sum_{j=0}^{k}p_{i}(1-p_{i})^{j}\varphi(\tilde{s}_{i}^{k+1-j})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$
  • (ii)
    $\mathbb{E}\|s^{k+1}-s^{k}\|^{2}_{P^{-1}-I}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k}\|^{2}_{I-P}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k+1}\|^{2}.$
Then, it follows
  • (iii)
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}\leq\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C$ is as in Lemma 5.
  • (iv)
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}\leq\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}$, where $D>0$ is a positive constant.
  • (v)
    $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}\leq M_{K}$ with $M_{K}\leq M_{\infty}$ and $M_{\infty}$ being a positive constant.
Proof.
For (i), since $\mathbb{E}\mathbb{E}^{k}[\cdot]=\mathbb{E}[\cdot]$, a direct induction gives,
$\mathbb{E}\varphi(s_{i}^{k+1})$ $=\mathbb{E}\mathbb{E}^{k}\varphi(s_{i}^{k+1})=\mathbb{E}[p_{i}\varphi(\tilde{s}_{i}^{k+1})+(1-p_{i})\varphi(s_{i}^{k})]$
$=\mathbb{E}p_{i}\varphi(\tilde{s}_{i}^{k+1})+\mathbb{E}\mathbb{E}^{k-1}(1-p_{i})\varphi(s_{i}^{k})$
$=\mathbb{E}[p_{i}\varphi(\tilde{s}_{i}^{k+1})+p_{i}(1-p_{i})\varphi(\tilde{s}_{i}^{k})+(1-p_{i})^{2}\varphi(s_{i}^{k-1})]=\cdots$
$=\mathbb{E}[\sum_{j=0}^{k}p_{i}(1-p_{i})^{j}\varphi(\tilde{s}_{i}^{k+1-j})]+(1-p_{i})^{k+1}\varphi(s_{i}^{0}),$ (4.4)
which leads to (i).
For (ii), the trick is to find the relations among those three items. That is, for every $k$, $\|\tilde{s}^{k+1}-s^{k}\|^{2}=\|s^{k+1}-s^{k}\|^{2}+\|\tilde{s}^{k+1}-s^{k+1}\|^{2}$ since $\|s_{i}^{k+1}-s_{i}^{k}\|^{2}$ will be zero except $i=i_{k}$ and $\|\tilde{s}_{i}^{k+1}-s_{i}^{k+1}\|^{2}$ will be zero except $i\neq i_{k}$. Using $\mathbb{E}\|s^{k+1}-s^{k}\|^{2}=\mathbb{E}\|\tilde{s}^{k+1}-s^{k}\|_{P}^{2}$, the statement (ii) follows. Now, let us prove $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ is uniformly bounded for all $k$, based on (3.4) and the iteration formula of PDR, we have
$\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ $=\mathbb{E}\|\tilde{y}^{k+1}_{t,\operatorname{test}}-y^{*}\|^{2}=\mathbb{E}\|\tilde{\bar{y}}_{t}^{k+1}-\bar{y}^{k}+\tilde{y}_{t}^{k+1}-y^{*}\|^{2}$
$\leq 3\mathbb{E}\|\tilde{\bar{y}}_{t}^{k+1}-\bar{y}^{*}\|^{2}+3\mathbb{E}\|\tilde{y}_{t}^{k+1}-y^{*}\|^{2}+3\mathbb{E}\|\bar{y}^{k}-\bar{y}^{*}\|^{2}$
$\leq 3\mathbb{E}\|\mathcal{T}u^{k}-u^{*}\|^{2}+3\tau\rho_{k,y}\mathbb{E}\|\bar{y}^{k}-\bar{y}^{*}\|_{\tau^{-1}\rho_{k,y}^{-1}\mathbf{P}^{-1}}^{2}$
$\leq 3C\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p}}^{2}+3\tau\rho_{u}\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}$
$\leq 3C\rho_{u}\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}+3\tau\rho_{u}\mathbb{E}\|u^{k}-u^{*}\|_{\mathbf{M}_{p,\varrho_{k}}}^{2}$
$=(3C+3\tau)\rho_{u}\mathbb{E}\Delta_{k}\leq(3C+3\tau)\rho_{u}\Delta^{0}\leq\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}$
where the third inequality is deduced from Lemma 5 while the last one is from (4.1). Applying (i) with $\varphi(\cdot)=\|\cdot-y^{*}\|^{2}$ on $y^{k+1}_{\operatorname{test}}$ and noting $10<\sum_{j=0}^{k}p_{i}(1-p_{i})^{j}<$ 1, 0<($10<(1-p_{i})^{k+1}<$ 1, then $0\prec\sum_{j=0}^{k}\mathbf{P}(I-\mathbf{P})^{j}\prec I$, $0\prec(I-\mathbf{P})^{k+1}\prec I$, we arrive at
$\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ $=\mathbb{E}\sum_{j=0}^{k}\|\tilde{y}^{k+1-j}_{\operatorname{test}}-y^{*}\|_{\mathbf{P}(I-\mathbf{P})^{j}}^{2}+\|y^{0}_{\operatorname{test}}-y^{*}\|_{(I-\mathbf{P})^{k+1}}^{2},$
$\leq\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2}.$ (4.5)
Now, let us turn to (iv). Remembering the relation between $\tilde{y}^{k+1}_{\operatorname{test}}$ and $\tilde{y}^{k+1}_{t,\operatorname{test}}$ as in (3.4), the updates of PDR and the non-expansiveness of the resolvent operator $(I+\tau\partial\mathcal{G})^{-1}$, it can be readily checked that
$\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{t,\operatorname{test}}-\tilde{y}^{k}_{t,\operatorname{test}}\|^{2}\leq\mathbb{E}\sum_{k=1}^{K}\|(2\tilde{y}_{t}^{k+1}-\bar{y}^{k})-(2\tilde{y}_{t}^{k}-\bar{y}^{k-1})\|^{2}$
$=\mathbb{E}\sum_{k=1}^{K}\|A_{1}(N_{1}-I)(x^{k}-x^{k-1})+A_{1}(\bar{x}^{k}-\bar{x}^{k-1})+A_{2}(N_{2}-I)(y^{k}-y^{k-1})$
$\quad+A_{3}(\bar{y}^{k}-\bar{y}^{k-1})\|^{2}\leq D\mathbb{E}\sum_{k=1}^{K}\|u^{k}-u^{k-1}\|_{\mathbf{M}_{p}}^{2},$ (4.6)
where $A_{1}=\tau A_{2}\mathcal{K}N_{1}^{-1}$, $A_{2}=2(N_{2}+\tau\sigma\mathcal{K}N_{1}^{-1}\mathcal{K}^{*})^{-1}$, $A_{3}=A_{2}-I$, and
$D=4\max({\sigma(\|A_{1}\|^{2}\|N_{1}-I\|),\ {\sigma}\|A_{1}\|^{2},\ \tau\|A_{2}\|^{2}\|N_{2}-I\|,\ {\tau}\|A_{3}\|^{2}}).$ (4.7)
Combining (4.6) and (4.1), we obtain (iv).
For (v), since $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|_{P}^{2}$, we henceforth focus on $\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$. By the inequality $\|a+b\|^{2}\leq(1+\epsilon)\|a\|^{2}+(1+{1}/{\epsilon})\|b\|^{2}$, for any $\epsilon>0$ and $K\geq 1$, we arrive at
$\sum_{k=1}^{K}\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}\leq\sum_{k=1}^{K}[\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+\mathbb{E}(1+\frac{1}{\epsilon})\|\tilde{y}^{k}_{\operatorname{test}}-y^{k}_{\operatorname{test}}\|^{2}]$
$=\sum_{k=1}^{K}\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+\sum_{k=1}^{K}\mathbb{E}(1+\frac{1}{\epsilon})\|\tilde{y}^{k}_{\operatorname{test}}-y^{k-1}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}$ (4.8)
$=\sum_{k=1}^{K}\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+(\sum_{k=1}^{K-1}\mathbb{E}(1+\frac{1}{\epsilon})\|\tilde{y}^{k+1}_{\operatorname{test}}-y^{k}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}+\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}),$
where the first equality is obtained by (ii). For the convenience of proofs, letting $\epsilon={1}/{\underline{p}}$ with $\underline{p}:=\min_{i}\{p_{i},\ i=1,\ldots,n\}$ henceforth, then $(1+{1}/{\epsilon})(I-\mathbf{P})\leq(1+{1}/{\epsilon})(1-\underline{p})I=(1-\underline{p}^{2})I$.
For employing Lemma 6, choosing $\delta=\underline{p}^{2}$ and $S_{0}=0$, we denote for $K\geq 1$,
$S_{K}:=\sum_{k=1}^{K}\mathbb{E}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2},\ \ h_{K}:=\sum_{k=1}^{K}\mathbb{E}(1+\epsilon)\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}+\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}.$
The equation (4.8) thus turns to
$S_{K}-(1-\delta)S_{K-1}\leq h_{K}.$ (4.9)
For $\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}$, we have
$\mathbb{E}(1+{1}/{\epsilon})\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{I-\mathbf{P}}^{2}\leq\mathbb{E}\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|_{(1-\delta)I}^{2}=\|\tilde{y}^{1}_{\operatorname{test}}-y^{0}_{\operatorname{test}}\|^{2}_{(1-\underline{p}^{2})I}.$ (4.10)
Then by (iv), $h_{K}$ is bounded, i.e.,
$h_{K}\leq(1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}.$ (4.11)
Thus by (4.9), (4.11), Lemma 6 and the fact $S_{0}=0$, we have
$S_{K}$ $\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)$ (4.12)
$\leq\frac{1}{\underline{p}^{2}}\left((1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right).$
Hence
$\mathbb{E}\sum_{k=0}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}+\mathbb{E}\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}$
$\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+\underline{p}^{-1})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
Finally, with (ii), we derive
$\mathbb{E}$ $\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}=\mathbb{E}\sum_{k=0}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|_{\mathbf{P}}^{2}\leq\mathbb{E}\sum_{k=0}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}$
$\leq\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}\left((1+1/\underline{p})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{K},$ (4.13)
$\leq{1}/{\underline{p}^{2}}\left((1+1/\underline{p})\frac{D\rho^{2}_{u}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{(1-\underline{p}^{2})I}^{2}\right)+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}:=M_{\infty}.$
Let us extend a lemma in [1] originated from [34] to positive semidefinite weights.
Lemma 8.
With notations $s^{k}$ defined in Lemma 4, picking up a point $s^{0}_{\dagger}\in\mathbb{Y}$ at will, let us define an iteration sequence $v^{k+1}=s^{k}-\tilde{s}^{k+1}-\mathbf{P}^{-1}(s^{k}-s^{k+1})$ with ${s}_{\dagger}^{k+1}={s}_{\dagger}^{k}-\mathbf{P}v^{k+1}$. Introduce a positive semidefinite, self-adjoint, and linear weight $\mathfrak{D}=\emph{Diag}[\mathfrak{d}_{1},\ldots,\mathfrak{d}_{n}]$ with $\mathfrak{d}_{i}:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$, and a non-increasing positive weight $\Theta^{k}=\emph{Diag}[\vartheta_{1}^{k}I,\ldots,\vartheta_{n}^{k}I]$ with $\vartheta_{i}^{k}I:\mathbb{Y}_{i}\rightarrow\mathbb{Y}_{i}$ and $0<\vartheta_{l}\leq\vartheta_{i}^{k+1}\leq\vartheta_{i}^{k}\leq\vartheta_{u}$ for $i=1,\ldots,n$ and $k=0,\ldots,K-1$. Here $\vartheta_{l}$ and $\vartheta_{u}$ are positive constants. We also assume that the weights $\Theta^{k}$, $\mathfrak{D}$ and the probability matrix $\mathbf{P}$ can communicate with each other, i.e., $\mathbf{P}\Theta^{k}=\Theta^{k}\mathbf{P}$, $\mathfrak{D}\Theta^{k}=\Theta^{k}\mathfrak{D}$ together with $\mathfrak{D}\mathbf{P}=\mathbf{P}\mathfrak{D}$ as before. Then, it follows that for any $s\in\mathbb{Y}$
$\sum_{k=0}^{K}\langle{s}_{\dagger}^{k}-s,v^{k+1}\rangle_{\mathfrak{D}\Theta^{k}}\leq\frac{1}{2}\|{s}_{\dagger}^{0}-s\|_{\mathfrak{D}\Theta^{0}\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{s}_{\dagger}^{K+1}-s\|_{\mathfrak{D}\Theta^{K}\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\Theta^{k}\mathbf{P}}^{2},$ (4.14)
where the above inequality in (4.14) becomes equality while $\Theta^{k}\equiv\Theta^{0}$. We further have
$\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|v^{k+1}\|_{\mathfrak{D}\Theta^{k}\mathbf{P}}^{2}]\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{2}\|s^{k}-s^{k+1}\|_{\mathfrak{D}\Theta^{k}\mathbf{P}^{-1}}^{2}].$ (4.15)
Moreover, $v^{k}$ and ${s}_{\dagger}^{k}$ are $\mathfrak{F}_{k}$ measurable and $\mathbb{E}^{k}[v^{k+1}]=0$. It follows that
$\mathbb{E}^{k}\langle{s}_{\dagger}^{k},v^{k+1}\rangle=0.$ (4.16)
Proof.
By the definition of $s^{k+1}_{\dagger}$, we have the estimate,
$\frac{1}{2}\|{s}_{\dagger}^{k+1}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}$ $=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}-\langle\mathbf{P}v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}+\frac{1}{2}\|\mathbf{P}v^{k+1}\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}$
$=\frac{1}{2}\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}-\langle v^{k+1},{s}_{\dagger}^{k}-s\rangle_{\mathfrak{D}\Theta^{k}}+\frac{1}{2}\|v^{k+1}\|_{\mathbf{P}\mathfrak{D}\Theta^{k}}^{2}.$ (4.17)
Summing the equality (4), we arrive at
$\sum_{k=0}^{K}\langle{s}_{\dagger}^{k}-s,v^{k+1}\rangle_{\mathfrak{D}\Theta^{k}}=\frac{1}{2}\sum_{k=0}^{K}\left[\|{s}_{\dagger}^{k}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}-\|{s}_{\dagger}^{k+1}-s\|_{\mathbf{P}^{-1}\mathfrak{D}\Theta^{k}}^{2}\right]+\frac{1}{2}\sum_{k=0}^{K}\|v^{k+1}\|_{\mathbf{P}\mathfrak{D}\Theta^{k}}^{2}.$
Noting that $\Theta^{k}\succeq\Theta^{k+1}$ together with $\mathbf{P}\mathfrak{D}\Theta^{k}\succeq\mathbf{P}\mathfrak{D}\Theta^{k+1}$ for $k=0,\ldots,K-1$, we obtain (4.14) from the above equation. For the second result, similar to $\mathbb{E}^{k}\|v^{k+1}\|^{2}=\mathbb{E}^{k}\|\mathbb{E}^{k}\xi-\xi\|^{2}\leq\mathbb{E}^{k}\|\xi\|^{2}$, with $\xi^{k}=\mathbf{P}^{-1}(s^{k}-s^{k+1})$ and $\mathbb{E}^{k}\xi=s^{k}-\tilde{s}^{k+1}$, we obtain
$\mathbb{E}[\sum_{k=0}^{K}\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}\Theta^{k}}^{2}]=\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|v^{k+1}\|_{\mathfrak{D}\mathbf{P}\Theta^{k}}^{2}]]$
$\leq\sum_{k=0}^{K}\mathbb{E}[\mathbb{E}^{k}[\|\mathbf{P}^{-1}(s^{k+1}-s^{k})\|_{\mathfrak{D}\mathbf{P}\Theta^{k}}^{2}]]=\sum_{k=0}^{K}\mathbb{E}[\|s^{k+1}-s^{k}\|_{\mathfrak{D}\mathbf{P}^{-1}\Theta^{k}}^{2}].$ (4.18)
By direct calculation, $\mathbb{E}^{k}[v^{k+1}]=0$. Since ${s}_{\dagger}^{k}$ is $\mathfrak{F}_{k}$ measurable, we conclude $\mathbb{E}^{k}\langle{s}_{\dagger}^{k},v^{k+1}\rangle=\langle{s}_{\dagger}^{k},\mathbb{E}^{k}v^{k+1}\rangle=0$. ∎
With the preceding lemmas, we now prove the sublinear convergence of SRPDR.
Theorem 3.
If $u=(x,y,\bar{x},\bar{y}),$ with $\bar{x}=x+\sigma\mathcal{K}^{\ast}y,$ $\bar{y}=y-\tau\mathcal{K}x$. Let $u^{k}$ be the iteration sequence generated by SRPDR, and let $z^{k}_{\operatorname{test}}=(x^{k}_{\operatorname{test}},y^{k}_{\operatorname{test}})$ be the corresponding transitional variables. Denote $K_{1}:=K+1$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K+1}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}=\frac{1}{K_{1}}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}\subset\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, a bounded domain that contains a saddle-point, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K_{1}}=\mathcal{O}(\frac{1}{K}),$ (4.19)
where the constant is given by
$C_{\mathbb{B}}=\sup_{z\in\mathbb{B}}\frac{1}{2\rho_{l}}\Big{[}\|x^{0}-x\|_{\frac{N_{1}-I}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{x}^{0}-x-\sigma\mathcal{K}^{*}y\|^{2}_{\frac{I}{\sigma}}+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}$
$+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}\Big{]}+\sup_{z\in\mathbb{B}}\frac{1}{2}(\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2})$
$+\frac{\rho_{u}^{2}}{2(2-\rho_{u})\rho^{2}_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}}M_{K}$
$+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}.$
Here $C$ is produced in Lemma 5, $M_{K}$ is from (4.13), $(x^{*},y^{*})$ is a saddle point, and $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$, $y^{0}_{\dagger,\operatorname{test}}$ are picked from $\mathbb{B}_{2}$ randomly. The same rate also holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
Proof.
Using the following two identities,
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)=\mathcal{F}(x_{\operatorname{test}}^{k+1})+\langle\mathcal{K}x_{\operatorname{test}}^{k+1},y\rangle-\mathcal{G}(y)-\mathcal{F}(x)-\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}\rangle+\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}),$
$\mathcal{H}((x_{\operatorname{test}}^{k+1},{y}_{\operatorname{test}}^{k+1}),z)=\mathcal{F}(x_{\operatorname{test}}^{k+1})+\langle\mathcal{K}x_{\operatorname{test}}^{k+1},y\rangle-\mathcal{G}(y)-\mathcal{F}(x)-\langle\mathcal{K}x,{y}_{\operatorname{test}}^{k+1}\rangle+\mathcal{G}({y}_{\operatorname{test}}^{k+1}),$
we obtain,
$\mathcal{H}((x_{\operatorname{test}}^{k+1},{y}_{\operatorname{test}}^{k+1}),z)=\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)+\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})+\mathcal{G}(y_{\operatorname{test}}^{k+1}).$ (4.20)
We now separate $\mathcal{H}((x_{\operatorname{test}}^{k+1},{y}_{\operatorname{test}}^{k+1}),z)$ into three parts including $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$, $\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle$ and $\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})$. By the convexity of $\mathcal{H}$ and the property of supremum, we have
$\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}(z^{K}_{\operatorname{test}},z)]\leq\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]$
$+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]+\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))].$ (4.21)
We will discuss these three parts one by one. First, denote by
$\mathbf{x}^{k}:=(x^{k},\bar{x}^{k}),\quad\mathbf{y}^{k}:=(y^{k},\bar{y}^{k}),\quad\tilde{\mathbf{y}}^{k}:=(\tilde{y}^{k},\tilde{\bar{y}}^{k}),\quad\mathbf{x}:=(x,\bar{x}),\quad\mathbf{y}:=(y,\bar{y}),$ (4.22)
and the following diagonal weights $\mathbf{w}_{x}:=\operatorname{Diag}[\tfrac{1}{\sigma}(N_{1}-I),\tfrac{1}{\sigma}I]$, $\mathbf{w}_{y}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I),\tfrac{1}{\tau}I]$, $\mathbf{R}_{x}^{k}:=\operatorname{Diag}[\tfrac{1}{\rho_{k}}I,\tfrac{1}{\rho_{k,x}}I]$, $\mathbf{R}_{y}^{k}:=\operatorname{Diag}[\tfrac{1}{\rho_{k}}I,\tfrac{1}{\rho_{k,y}}I]$, $\mathbf{r}_{x}^{k}:=\operatorname{Diag}[\frac{2-\rho_{k}}{\rho_{k}^{2}}I,\frac{2-\rho_{k,x}}{\rho_{k,x}^{2}}I]$, and $\mathbf{r}_{y}^{k}:=\operatorname{Diag}[\frac{2-\rho_{k}}{\rho_{k}^{2}}I,\frac{2-\rho_{k,y}}{\rho_{k,y}^{2}}I]$. Therefore, $\|u^{k}-u\|^{2}_{\mathbf{M_{p}}}=\|\mathbf{x}^{k}-\mathbf{x}\|^{2}_{\mathbf{w}_{x}}+\|\mathbf{y}^{k}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}}$, in which $\mathbf{w}_{y}\mathbf{P}:=\operatorname{Diag}[\tfrac{1}{\tau}(N_{2}-I)\mathbf{P},\tfrac{1}{\tau}\mathbf{P}]$. For $\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$, we focus on the iteration $u^{k}\rightarrow\tilde{u}^{k+1}$. Combining Proposition 1 and (3.3), we obtain
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)\leq\Delta_{\mathbf{x}}^{k}+\Delta_{\mathbf{y}}^{k},$ (4.23)
where $\Delta_{\mathbf{x}}$ and $\Delta_{\mathbf{y}}$ are defined as
$\begin{array}[]{c}\Delta_{\mathbf{x}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}\mathbf{r}_{x}^{k}}^{2}\Big{)},\\ \Delta_{\mathbf{y}}^{k}:=\frac{1}{2}\Big{(}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}\Big{)}.\\ \end{array}$ (4.24)
Let us first reformulate $\Delta_{\mathbf{y}}^{k}$ as follows
$\Delta_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}+\|\mathbf{y}^{k}-\ y\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}$
$+(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}$
$=-\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}+\frac{1}{2}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\frac{1}{2}\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}+\epsilon_{\mathbf{y}}^{k},$ (4.25)
where $\epsilon_{\mathbf{y}}^{k}$ is defined as follows with $\mathbf{v}^{k+1}:=\mathbf{y}^{k}-\tilde{\mathbf{y}}^{k+1}-\mathbf{P}^{-1}(\mathbf{y}^{k}-\mathbf{y}^{k+1})$ for integer $k\geq 0$,
$\epsilon_{\mathbf{y}}^{k}$ $=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-(\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}$
$=\frac{1}{2}\Big{[}\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})-2\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}\Big{]}.$ (4.26)
Now, with (4.23), (4.24) and (4), we arrive at
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq\frac{1}{2}\Big{(}\|\mathbf{x}^{k}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}\|_{\mathbf{w}_{x}\mathbf{R}_{x}^{k}}^{2}-\|\mathbf{x}^{k+1}-\mathbf{x}^{k}\|_{\mathbf{w}_{x}\mathbf{r}_{x}^{k}}^{2}$
$+\|\mathbf{y}^{k}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}-\mathbf{y}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}^{2}\Big{)}+\epsilon_{\mathbf{y}}^{k}$
$=\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\|u^{k+1}-u\|^{2}_{\mathbf{M}_{p,\varrho_{k}}}-\|u^{k+1}-u^{k}\|^{2}_{(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}\mathbf{M}_{p}}\Big{)}$
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}+\frac{1}{2}\Big{(}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}}$
$+\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}\Big{)},$ (4.27)
where the equality is by adding and subtracting $\frac{1}{2}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}$ and the definition of $\epsilon_{\mathbf{y}}^{k}$ in (4.26). In (4.27), dropping $-\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{p}(2I-I_{\varrho_{k}})I^{-2}_{\varrho_{k}}}$ and summing over $k=0,\ldots,K$ yields,
$\sum^{K}_{k=0}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)$ $\leq-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p,\varrho_{K}}}+\frac{1}{2}\|u^{0}-u\|^{2}_{\mathbf{M}_{p,\varrho_{0}}}+\mathcal{E}^{K}+\sum_{k=0}^{K}[-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}],$ (4.28)
where we use $\mathbf{M}_{p,\varrho_{k+1}}\leq\mathbf{M}_{p,\varrho_{k}}$ for all $k=0,\dots,K$. $\mathcal{E}^{K}$ is defined as follows
$\frac{1}{2}\sum_{k=0}^{K}\Big{[}$ $(\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}-\|\tilde{\mathbf{y}}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{r}_{y}^{k}})$
$+(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\tilde{\mathbf{y}}^{k+1}\|_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}+\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}.$
Noting there are no items involving $(x,y)$ in $\mathcal{E}^{K}$, we can take expectation directly without concerning taking the supremum over $(x,y)$. With (3.9) and the fact $\mathbb{E}[\mathbb{E}^{k}[\cdot]]=\mathbb{E}[\cdot]$, we get
$\mathbb{E}\mathcal{E}^{K}/K_{1}$ $=\frac{1}{2K_{1}}\mathbb{E}\sum_{k=0}^{K}\Big{[}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}-\mathbb{E}^{k}\|\mathbf{y}^{k+1}-\mathbf{y}^{k}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{r}_{y}^{k}}$
$+\mathbb{E}^{k}(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})-(\|\mathbf{y}^{k}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}-\|\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2})\Big{]}=0.$ (4.29)
For the sum of inner products in (4.28), introducing $\mathbf{y}^{k+1}_{\dagger}=\mathbf{y}^{k}_{\dagger}-\mathbf{P}\mathbf{v}^{k+1}$ and selecting a $\mathbf{y}^{0}_{\dagger}=(y^{0}_{\dagger},\bar{y}^{0}_{\dagger})\in\mathbb{B}_{2}\times\mathbb{B}_{2}$ randomly, we have
$-\langle\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}=\langle\mathbf{y}_{\dagger}^{k}-\mathbf{y},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}.$
In (4.28), dropping $-\frac{1}{2}\|u^{K+1}-u\|^{2}_{\mathbf{M}_{p}\mathbf{R}^{k}}$, dividing both sides by $K_{1}$, taking supremum with respect to $z$, and employing (4.14) on the above inner product, we arrive at
$\mathbb{E}[$ $\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\sup_{z\in\mathbb{B}}\frac{1}{2K_{1}}[\|u^{0}-u\|^{2}_{\mathbf{M}_{p,\varrho_{0}}}+\|\mathbf{y}^{0}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{0}}]$ (4.30)
$-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}+\mathbb{E}\sum_{k=0}^{K}\frac{1}{K_{1}}[\frac{1}{2}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}\mathbf{R}_{y}^{k}}-\langle\mathbf{y}_{\dagger}^{k},\mathbf{v}^{k+1}\rangle_{\mathbf{w}_{y}\mathbf{R}_{y}^{k}}].$
By (4.15), non-decreasing of $\rho_{k}$, $\rho_{k,y}$ and (4.1) we have,
$\mathbb{E}\sum_{k=0}^{K}\|\mathbf{v}^{k+1}\|^{2}_{\mathbf{w}_{y}\mathbf{P}\mathbf{R}_{y}^{k}}\leq\mathbb{E}[\sum_{k=0}^{K}\|\mathbf{y}^{k}-\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}^{2}]\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{\rho_{l}}\|\mathbf{y}^{k}-\mathbf{y}^{k+1}\|_{\mathbf{w}_{y}\mathbf{P}^{-1}}^{2}]$ (4.31)
$\leq\mathbb{E}[\sum_{k=0}^{K}\frac{1}{\rho_{l}}\|u^{k}-u^{k+1}\|_{\mathbf{M}_{p}}^{2}]\leq\frac{\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\Delta^{0}=\frac{\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}^{2}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}.$
Thus in (4.30), dropping $-\mathbb{E}\|\mathbf{y}^{K+1}_{\dagger}-\mathbf{y}\|^{2}_{\mathbf{w}_{y}\mathbf{P}^{-1}\mathbf{R}_{y}^{k}}$, together with (4.16) and (4.31), we can obtain
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}_{\operatorname{test}}^{k+1}),z)]\leq\frac{1}{2K_{1}\rho_{l}}\big{[}\sup_{z\in\mathbb{B}}(\|u^{0}-u\|^{2}_{\mathbf{M}_{p}}+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}$ (4.32)
$+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}})+\frac{\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}\big{]}.$
Second, let us give the upper bounds for $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$. Denote by $v_{\operatorname{test}}^{k+1}=y_{\operatorname{test}}^{k}-\tilde{y}_{\operatorname{test}}^{k+1}-\mathbf{P}^{-1}(y_{\operatorname{test}}^{k}-y_{\operatorname{test}}^{k+1})$, $y^{k+1}_{\dagger,\operatorname{test}}=y^{k}_{\dagger,\operatorname{test}}-\mathbf{P}v_{\operatorname{test}}^{k+1}$ and select a $y^{0}_{\dagger,\operatorname{test}}\in\mathbb{B}_{2}$ randomly. By the definition of $y^{k}_{\dagger,\operatorname{test}}$, $v^{k}_{\operatorname{test}}$ and (4.14), we obtain
$\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle$
$=$ $\sum_{k=0}^{K}-\langle\mathcal{K}x,y_{\operatorname{test}}^{k}-\tilde{y}_{\operatorname{test}}^{k+1}-\mathbf{P}^{-1}(y_{\operatorname{test}}^{k}-{y}_{\operatorname{test}}^{k+1})\rangle+\sum_{k=0}^{K}\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k})\rangle$
$=$ $\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}}-\mathcal{K}x,v^{k+1}_{\operatorname{test}}\rangle-\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}},v^{k+1}_{\operatorname{test}}\rangle+\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-{y}_{\operatorname{test}}^{0})\rangle$
$\leq$ $\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}-\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{K+1}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\sum_{k=0}^{K}\frac{1}{2}\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}-\sum_{k=0}^{K}\langle y^{k}_{\dagger,\operatorname{test}},v^{k+1}_{\operatorname{test}}\rangle$
$+\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-{y}_{\operatorname{test}}^{0})\rangle.$ (4.33)
For the last term in (4.33), we have
$\langle\mathcal{K}x,(\mathbf{P}^{-1}-I)(y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0})\rangle\leq\frac{1}{2}\|\mathcal{K}x\|^{2}+\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}.$ (4.34)
In (4.33), dropping $-\frac{1}{2}\|{y}_{\dagger,\operatorname{test}}^{K+1}-\mathcal{K}x\|^{2}$ and by (4.34), (4.16), we get
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]$ $\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K_{1}}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K_{1}}\|\mathcal{K}x\|^{2}]$
$+\frac{1}{2K_{1}}\mathbb{E}\sum_{k=0}^{K}\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}+\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K_{1}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}.$ (4.35)
For the last item in (4.35), inserting $y^{*}$ and employing Lemma 7, we get,
$\mathbb{E}\frac{(1-\underline{p})^{2}}{2\underline{p}^{2}K_{1}}\|y_{\operatorname{test}}^{K+1}-y_{\operatorname{test}}^{0}\|^{2}$ $\leq\mathbb{E}\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}\|y_{\operatorname{test}}^{K+1}-y^{*}\|^{2}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}\|y^{*}-y_{\operatorname{test}}^{0}\|^{2}$
$\leq\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2}).$ (4.36)
Noting $\mathbb{E}\sum_{k=0}^{K}[\|v_{\operatorname{test}}^{k+1}\|_{\mathbf{P}}^{2}]\leq\mathbb{E}\sum_{k=0}^{K}\|y_{\operatorname{test}}^{k}-y_{\operatorname{test}}^{k+1}\|_{\mathbf{P}^{-1}}^{2}$ by (4.15) together with (4.13), finally, we arrive at
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}\langle\mathcal{K}x,\tilde{y}_{\operatorname{test}}^{k+1}-{y}_{\operatorname{test}}^{k+1}\rangle]\leq\sup_{z\in\mathbb{B}}[\frac{1}{2K_{1}}\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\frac{1}{2K_{1}}\|\mathcal{K}x\|^{2}]$
$+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}K_{1}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}K_{1}}M_{K}.$ (4.37)
Third, we show $\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]$ is also bounded. In fact,
$\mathbb{E}[\sup_{z\in\mathbb{B}}\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1}))]=\mathbb{E}\frac{1}{K_{1}}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})].$ (4.38)
Employing (3.8) on $\mathcal{G}(\tilde{y}^{k+1}_{\operatorname{test}})$, we get $\mathcal{G}(\tilde{y}^{k+1}_{\operatorname{test}})=\mathbb{E}^{k}\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}})-\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})$. Here $\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}}):=\sum^{n}_{i=1}1/p_{i}\mathcal{G}_{i}(y^{k+1}_{\operatorname{test},i})$ and $\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k+1}_{\operatorname{test}}):=\sum^{n}_{i=1}(1/p_{i}-1)\mathcal{G}_{i}(y^{k+1}_{\operatorname{test},i})$. Thus, using the fact $\mathbb{E}\mathbb{E}[\cdot]=\mathbb{E}[\cdot]$ again, we arrive at
$\mathbb{E}\frac{1}{K_{1}}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]=\frac{1}{K_{1}}\sum_{k=0}^{K}\mathbb{E}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathbb{E}^{k}\mathcal{G}_{\mathbf{P}^{-1}}(y^{k+1}_{\operatorname{test}})+\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]$
$=\frac{1}{K_{1}}\sum_{k=0}^{K}[-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k+1}_{\operatorname{test}})+\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{k}_{\operatorname{test}})]=\frac{1}{K_{1}}(\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\mathbb{E}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{{K}+1}_{\operatorname{test}})).$ (4.39)
Since $\mathcal{K}_{i}x^{*}\in\partial\mathcal{G}_{i}(y^{*}_{i})$, one has
$\mathcal{G}_{i}(y_{\operatorname{test},i}^{K+1})\geq\mathcal{G}_{i}(y_{i}^{*})+\langle\mathcal{K}_{i}x^{*},y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\rangle\geq\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\|y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\|.$
We thus obtain
$\mathbb{E}[\mathcal{G}_{\mathbf{P}^{-1}-I}(y_{\operatorname{test}}^{K+1})$ $]=\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathbb{E}[\mathcal{G}_{i}(y_{\operatorname{test},i}^{K+1})]\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\mathbb{E}[\|y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\|])$
$\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\sqrt{\mathbb{E}[\|y_{\operatorname{test},i}^{K+1}-y_{i}^{*}\|^{2}]})$
$\geq\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)(\mathcal{G}_{i}(y_{i}^{*})-\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}),$ (4.40)
where the second inequality is by Jensen’s inequality and the third inequality is by Lemma 7. We thus have
$\mathbb{E}\frac{1}{K_{1}}\sum_{k=0}^{K}[\mathcal{G}(y_{\operatorname{test}}^{k+1})-\mathcal{G}(\tilde{y}_{\operatorname{test}}^{k+1})]$ $\leq\frac{1}{K_{1}}\Big{(}\mathcal{G}_{\mathbf{P}^{-1}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})$
$+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y^{*}_{i}\|^{2}}\Big{)}.$ (4.41)
Finally, with (4.21) and combining (4.32), (4.37) and (4.41), we get (4.19) with the constant $C_{\mathbb{B}}$. For the Bregman distance, replacing $z=(x,y)$ by $(x^{*},y^{*})$, one can obtain the same sublinear convergence rate for the ergodic sequences. ∎
In the next theorem, we investigate the convergence rate of expected primal error when Lipschitz smoothness is satisfied for $\mathcal{G}$.
Theorem 4.
For the ergodic sequence $z_{\operatorname{test},K}=(x_{\operatorname{test},K},y_{\operatorname{test},K})$ defined in Theorem 3, $K_{1}=K+1$ and any $z=(x,y)\in\operatorname{dom}\mathcal{F}\times\operatorname{dom}\mathcal{G}$, if $\mathcal{G}^{*}$ is $\beta$-Lipschitz continuous with respect to the norm $\|\cdot\|_{\frac{I}{\tau}}$, and $y^{0}$, $(\bar{y}^{0}+\tau\mathcal{K}x^{*})$, $y^{0}_{\dagger}$, $(\bar{y}^{0}_{\dagger}+\tau\mathcal{K}x^{*})\in\operatorname{dom}\mathcal{G}$, and if $y^{*}_{\dagger}\in\operatorname{dom}\mathcal{G}$, such that $\sigma\mathcal{K}^{*}y^{*}_{\dagger}=\bar{x}^{0}-x^{*}$, then the expected primal error also exhibits a convergence rate of $\mathcal{O}(1/K)$, i.e,
$\mathbb{E}[\mathcal{F}(x_{\operatorname{test},K})+\mathcal{G}^{*}(\mathcal{K}x_{\operatorname{test},K})-\mathcal{F}(x^{*})-\mathcal{G}^{*}(\mathcal{K}x^{*})]\leq\frac{C_{p,2}}{K_{1}}=\mathcal{O}(1/K),$ (4.42)
where
$C_{p,2}=C_{p}+\frac{1}{2\rho_{l}}(\|x^{0}-x^{*}\|_{\frac{N_{1}-I}{\sigma}}^{2}+4\sigma\tau\beta^{2}\|\mathcal{K}^{*}\|^{2}+8\|N_{2}-I\|\beta^{2}/\underline{p}+8\beta^{2}/\underline{p})$
$\quad\qquad+\frac{1}{2}(\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x^{*}\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x^{*}\|^{2}),\ \text{with}$
$C_{p}:=\frac{\rho_{u}^{2}}{2(2-\rho_{u})\rho^{2}_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+\frac{(1-\underline{p})^{2}}{\underline{p}^{2}}(\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{p}}+2\|y^{0}_{\operatorname{test}}-y^{*}\|^{2})+\frac{1}{2\underline{p}}M_{K}$
$+\mathcal{G}_{\mathbf{P}^{{-1}}-I}(y^{0}_{\operatorname{test}})-\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\mathcal{G}_{i}(y_{i}^{*})+\sum_{i=1}^{n}(\frac{1}{p_{i}}-1)\|\mathcal{K}_{i}x^{*}\|\sqrt{\frac{(3C+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|_{\mathbf{M}_{p}}^{2}+\|y^{0}_{\operatorname{test},i}-y_{i}^{*}\|^{2}}.$
Proof.
Let us first define a new distance function ¯$\bar{\mathcal{H}}(z_{\operatorname{test},K},z)$,
¯$\bar{\mathcal{H}}(z_{\operatorname{test},K},z):=\mathcal{H}(z_{\operatorname{test},K},z)-\frac{1}{2\rho_{l}K_{1}}(\|x^{0}-x\|_{\frac{N_{1}-I}{\sigma}}^{2}+\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}+\|\bar{x}^{0}-x-\sigma\mathcal{K}^{*}y\|^{2}_{\frac{I}{\sigma}}$
$+\|\bar{y}^{0}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}+\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}+\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}})-\frac{1}{2K_{1}}(\|{y}_{\dagger,\operatorname{test}}^{0}-\mathcal{K}x\|_{\mathbf{P}^{-1}}^{2}+\|\mathcal{K}x\|^{2}),$
where $y^{0}_{\dagger}$, $\bar{y}^{0}_{\dagger}$ and ${y}_{\dagger,\operatorname{test}}^{0}$ are given as in Theorem 3. Following the same proof of Theorem 3, we can drive, $\mathbb{E}\sup_{z\in\mathbb{B}}\bar{\mathcal{H}}(z_{\operatorname{test},K},z)\leq\frac{C_{p}}{K_{1}}$. Then substituting $z=(x^{*},y)$ for any $y\in\mathbb{B}_{2}$ in ¯$\bar{\mathcal{H}}(z_{\operatorname{test},K},z)$, we derive $\mathbb{E}\sup_{y\in\mathbb{B}_{2}}\bar{\mathcal{H}}(z_{\operatorname{test},K},(x^{*},y))\leq\frac{C_{p}}{K_{1}}$. By the definition of $y^{*}_{\dagger}$, we arrive at $\|\bar{x}^{0}-x^{*}-\sigma\mathcal{K}^{*}y\|^{2}_{\frac{I}{\sigma}}\leq\sigma\|\mathcal{K}^{*}\|^{2}\|y^{*}_{\dagger}-y\|^{2}$. With [4, Corollary 17.19] and noting that $\mathcal{G}^{*}_{i}$ is $\beta$-Lipschitz continuous on the norm $\|\cdot\|_{\frac{I}{\tau}}$, we have $\|y^{*}_{\dagger}-y\|^{2}\leq 4\tau\beta^{2}$, $\|y^{0}-y\|_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}^{2}\leq 4\|N_{2}-I\|\beta^{2}/\underline{p}$, $\|\bar{y}^{0}-y+\tau\mathcal{K}x^{*}\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}\leq 4\beta^{2}/\underline{p}$, $\|y^{0}_{\dagger}-y\|^{2}_{\frac{(N_{2}-I)\mathbf{P}^{-1}}{\tau}}\leq 4\|N_{2}-I\|\beta^{2}/\underline{p}$ and $\|\bar{y}^{0}_{\dagger}-y+\tau\mathcal{K}x^{*}\|^{2}_{\frac{\mathbf{P}^{-1}}{\tau}}\leq 4\beta^{2}/\underline{p}$. We thus obtain
$\mathbb{E}\sup_{y\in\mathbb{B}_{2}}\mathcal{H}(z_{\operatorname{test},K},(x^{*},y))$
$=\mathbb{E}\sup_{y\in\mathbb{B}_{2}}[\mathcal{F}(x_{\operatorname{test},K})+\langle\mathcal{K}x_{\operatorname{test},K},y\rangle-\mathcal{G}(y)-\mathcal{F}(x^{*})-\langle\mathcal{K}x^{*},y_{\operatorname{test},K}\rangle+\mathcal{G}(y_{\operatorname{test},K})]\leq\frac{C_{p,2}}{K_{1}}.$ (4.43)
By Lipschitzness of $\mathcal{G}^{*}$, we pick $y\in\partial\mathcal{G}^{*}(\mathcal{K}x_{\operatorname{test},K})\neq\mathbf{0}$ s.t., $\langle\mathcal{K}x_{\operatorname{test},K},y\rangle-\mathcal{G}(y)=\mathcal{G}^{*}(\mathcal{K}x_{\operatorname{test},K})$, and by Fenchel-Young inequality, $\mathcal{G}(y_{\operatorname{test},K})-\langle\mathcal{K}x^{*},y_{\operatorname{test},K}\rangle\geq-\mathcal{G}^{*}(\mathcal{K}x^{*})$. Employing these inequalities in (4.43), we derive (4.42). ∎
5 Stochastic PDR for quadratic problems (SRPDRQ)
In this section, we will focus on the following relaxed and stochastic PDRQ method directly for compactness
$\left\{\begin{array}[]{ll}\tilde{x}^{k+1}=\left(N_{1}+\sigma Q+\mathcal{K}^{*}N_{2}^{-1}\mathcal{K}\right)^{-1}\left(N_{1}x^{k}+\sigma f-\sigma\mathcal{K}^{*}N_{2}^{-1}\left(\left(N_{2}-I\right)y^{k}+\bar{y}^{k}\right)\right)\\ x^{k+1}=(1-\rho_{k})x^{k}+\rho_{k}\tilde{x}_{t}^{k+1},\\ y^{k+1}=\left\{\begin{array}[]{ll}(1-\rho_{k})y_{i}^{k}+\rho_{k}\tilde{y}^{k+1}_{t,i},\quad\tilde{y}^{k+1}_{t,i}:=N_{2,i}^{-1}[(N_{2,i}-I)y_{i}^{k}+\bar{y}_{i}^{k}+\tau\mathcal{K}_{i}\tilde{x}^{x+1}]&i=i_{k}\\ y^{k}_{i}&\forall i\neq i_{k}\\ \end{array}\right.\\ y_{\operatorname{test}}^{k+1}=\left\{\begin{array}[]{ll}(I+\tau\partial\mathcal{G}_{i})^{-1}[2\tilde{y}^{k+1}_{t,i}-\bar{y}^{k}_{i}]&i=i_{k},\\ y_{\operatorname{test},i}^{k}&\forall i\neq i_{k},\\ \end{array}\right.\\ \bar{y}^{k+1}=\left\{\begin{array}[]{ll}\bar{y}_{i}^{k}+\rho_{k,y}(y_{\operatorname{test},i}^{k+1}-\tilde{y}_{t,i}^{k+1})&i=i_{k},\\ \bar{y}^{k}_{i}&\forall i\neq i_{k},\end{array}\right.\\ \end{array}\right.$ (SRPDRQ)
where $N_{1}\succeq 0$, $N_{2}\succeq I$ and $N_{2}$ is separable, i.e., $N_{2}y=(N_{2,1}y_{1},\ldots,N_{2,n}y_{n})^{T}$. Here $\rho_{k}$, $\rho_{k,y}$ are non-decreasing and have lower and upper limit $\rho_{l},\rho_{u}$, and $20<\rho_{l}<\rho_{u}<$ 2 which is similar as the SRPDR. It can be observed that the updates for $\bar{x}^{k+1}$ are not required in the SRPDRQ iterations, unlike in SRPDR. As a result, SRPDRQ can be more efficient in terms of iteration time.
To clarify, we refer to the general SRPDRQ as the stochastic fully relaxed preconditioned Douglas-Rachford method for quadratic problems (f-SRPDRQ) when $\rho_{k}\neq 1$ and $\rho_{k,y}\neq 1$. While when $\rho_{k}=1$, and $\rho_{k,y}=1$, we call it the stochastic preconditioned Douglas-Rachford method for quadratic problems (SPDRQ).
If we set $N_{2}=I$ and $N_{1}$ as in (2.11), we derive the preconditioned iteration (2.12). Define $x_{\operatorname{test}}^{k+1}=x^{k+1}$ and $\mathcal{T}_{Q}$ as in (2.14), we do one-step deterministic update based on the $u^{k}$ and get $\tilde{u}^{k+1}$, i.e.,
$\tilde{u}^{k+1}:=\mathcal{T}_{Q}u^{k},\quad\text{with}\ \ \tilde{u}^{k+1}=(\tilde{x}^{k+1},\tilde{y}^{k+1},\tilde{\bar{y}}^{k+1}),$ (5.1)
where $x^{k+1}=\tilde{x}^{k+1}$ and $\bar{x}^{k+1}=\tilde{\bar{x}}^{k+1}$. Noting $I_{q,\varrho_{k}}:=\textrm{Diag}[\rho_{k},\rho_{k},\rho_{k,y}]$ as before, with Proposition 3 and Lemma 3, we can derive the following lemma and theorem similarly.
Lemma 9.
For any $(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}$, a bounded domain, $\bar{x}=x+\sigma\mathcal{K}^{\ast}y$ and $\bar{y}=y-\tau\mathcal{K}x$, denoting $\mathbf{M}_{Qp}:=\operatorname{Diag}[\tfrac{1}{\sigma}N_{1},\tfrac{1}{\tau}(N_{2}-I)\mathbf{P}^{-1},\tfrac{1}{\tau}\mathbf{P}^{-1}]$, it holds that
$\mathcal{H}((x_{\operatorname{test}}^{k+1},\tilde{y}^{k+1}_{\operatorname{test}}),(x,y))$ $\leq\frac{1}{2}\Big{(}\|u^{k}-u\|^{2}_{\mathbf{M}_{Qp}I_{q,\varrho_{k}}^{-1}}-\mathbb{E}^{k}\|u^{k+1}-u\|^{2}_{\mathbf{M}_{Qp}I_{q,\varrho_{k}}^{-1}}$ (5.2)
$-\mathbb{E}^{k}\|u^{k+1}-u^{k}\|^{2}_{\mathbf{M}_{Qp}(2I-I_{q,\varrho_{k}})I_{q,\varrho_{k}}^{-2}}\Big{)}.$
Denote $\mathcal{Z}^{*}$ as the set for the saddle points of (1.1) and $\mathbb{F}\subset\mathbb{U}$ as the set of the fixed points for $\mathcal{T}_{Q}$.
Theorem 5.
The iteration sequence $u^{k}=(x^{k},y^{k},\bar{y}^{k})$ produced by SRPDRQ converges weakly to an $\mathbb{F}$-valued random variable $u^{**}=(x^{**},y^{**},\bar{y}^{**})$ almost surely, and $(x^{**},y^{**})$ is a $\mathcal{Z}^{*}$-valued random variable with $Qx^{**}+\mathcal{K}^{*}y^{**}=f$ and ${y}^{**}-\tau\mathcal{K}{x}^{**}=\bar{y}^{**}$. Moreover the sequence $(x^{k+1}_{\operatorname{test}}$, $y^{k+1}_{\operatorname{test}})$ also converges weakly to the $\mathcal{Z}^{*}$-valued random variable $(x^{**},y^{**})$ almost surely.
For the gap estimate, we give the following two lemmas which are similar to the SRPDR case.
Lemma 10.
There exists a constant $C_{Q}>0$ such that for all $u^{1},u^{2}\in\mathbb{U}_{Q}$ and $\tilde{u}^{1}=\mathcal{T}_{Q}u^{1}$, $\tilde{u}^{2}=\mathcal{T}_{Q}u^{2}$ and it leads to
$\|\tilde{u}^{1}-\tilde{u}^{2}\|^{2}\leq C_{Q}\|u^{1}-u^{2}\|_{\mathbf{M}_{Qp}}^{2}$ (5.3)
where $C_{Q}$ is only related to $N_{1}$, $N_{2}$, $f$, $\sigma$, $\tau$ and $\mathcal{K}$.
Lemma 11.
Similar to Lemma 7, we give the following estimates of $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}$ and $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|$.
  • (i)
    $\mathbb{E}\|y^{k+1}_{\operatorname{test}}-y^{*}\|^{2}<\frac{(3C_{Q}+3\tau)\rho_{u}}{\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{Q_{p}}}+\|y^{0}_{\operatorname{test}}-y^{*}\|^{2},$ where $C_{Q}$ is given in Lemma 10.
  • (ii)
    $\mathbb{E}\sum_{k=1}^{K}\|\tilde{y}^{k+1}_{\operatorname{test}}-\tilde{y}^{k}_{\operatorname{test}}\|^{2}<\frac{D_{Q}\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{Q_{p}}}$, where
    $D_{Q}=3\max(\sigma\|N_{1}\|\|A_{1}\|^{2},\ \tau\|N_{2}-I\|\|A_{2}\|^{2},\ \tau\|A_{3}\|^{2})$
    and $A_{2}=2(N_{2}+\tau\sigma\mathcal{K}(N_{1}+\sigma Q)^{-1}\mathcal{K}^{*})^{-1}$, $A_{1}=A_{2}\tau\mathcal{K}(N_{1}+\sigma Q)^{-1}$, $A_{3}=A_{2}-I$.
  • (iii)
    $\mathbb{E}\sum_{k=0}^{K}\|y^{k+1}_{\operatorname{test}}-{y}^{k}_{\operatorname{test}}\|^{2}<M_{Q,K}$, where
    $M_{Q,K}=\frac{1-(1-\underline{p}^{2})^{K-1}}{\underline{p}^{2}}((1+1/\underline{p})\frac{D_{Q}\rho_{u}^{2}}{(2-\rho_{u})\rho_{l}}\|u^{0}-u^{*}\|^{2}_{\mathbf{M}_{Q_{p}}}+\|\tilde{y}_{\operatorname{test}}^{1}-y^{0}_{\operatorname{test}}\|_{1-\underline{p}^{2}}^{2})+\|\tilde{y}^{1}_{\operatorname{test}}-{y}^{0}_{\operatorname{test}}\|^{2}.$
Concerning the restricted gap estimate for SRPDRQ, we have the following similar theorem.
Theorem 6.
If $u=(x,y,\bar{y}),$ with $\bar{y}=y-\tau\mathcal{K}x$, let $u^{k}$ be the iteration sequence generated by SRPDRQ, and let $z^{k}_{\operatorname{test}}=(x^{k},y^{k}_{\operatorname{test}})$ be the corresponding transitional variables. Denote $K_{1}:=K+1$, then the ergodic sequence $z_{\operatorname{test},K}=\frac{1}{K+1}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}=\frac{1}{K_{1}}\sum_{k=0}^{K}z^{k}_{\operatorname{test}}$ converges with rate $\mathcal{O}({1}/{K})$ in an expected restricted primal-dual gap sense, i.e., for any $z=(x,y)\in\mathbb{B}_{1}\times\mathbb{B}_{2}=\mathbb{B}$, a bounded domain, it holds that
$\mathbb{E}\mathfrak{G}_{z\in\mathbb{B}}(z_{\operatorname{test},K})=\mathbb{E}[\sup_{z\in\mathbb{B}}\mathcal{H}(z_{\operatorname{test},K},z)]\leq\frac{C_{\mathbb{B}}}{K_{1}}=\mathcal{O}({1}/{K}),$ (5.4)
where the constant $C_{\mathbb{B}}$ is a positive constant depending on $u^{0}$, $\mathcal{K}$, $\mathbf{P}$, $\rho_{l}$, $\rho_{u}$, $\mathbf{M}_{Q_{p}}$, $u^{*}$, $z$, and $C_{Q}$, $M_{Q,K}$. Here $C_{Q}$ is produced from Lemma 10, and $M_{Q,K}$ is from Lemma 11. Therefore the same rate holds for the Bregman distance $\mathcal{H}(z_{\operatorname{test},K},(x^{*},y^{*}))$.
Remark 2.
Theorem 2, 3, 5, and 6 can also cover the situation of random sets of indices to be selected and updated. Sample the dual variable $y=(y_{1},\dots,y_{n})$ properly [36, 37, 13] into multiple blocks $\mathbb{S}_{1}$, $\mathbb{S}_{2},\dots,\mathbb{S}_{b}$, such that $\mathbb{S}_{1}\cup\mathbb{S}_{2}\cup\dots\cup\mathbb{S}_{b}=\{y_{1},y_{2},\dots,y_{n}\}$ and $\mathbb{S}_{i}\cap\mathbb{S}_{j}=\emptyset$ for any $i,j\in\{1,\dots,b\}$, $i\neq j$. We denote the probability of selecting a block $i\in\{1,\ldots,b\}$ by $p_{i}$ where we assume $10<p_{i}<$ 1 for each $i$ and $\sum_{i=1}^{b}p_{i}=1$. Following the same routine, we can derive the same results on almost sure convergence and sublinear convergence of restricted primal-dual gap.
6 Numerical Experiments
In this section, we will illustrate two experiments, one is a deblurring problem, and the other is a Binary classification problem with real data.
6.1 TGV-Deblur Problem
In the first experiments, we focus on the following image deblurring problem (6.1) based on the Kullback-Leibler (KL) divergence and the anisotropic second-order Total Generalized Variation (TGV) [6] regularization
$\min_{u\in U}\mathcal{F}(u)+\mathcal{G}_{1}^{*}(\mathcal{K}_{1}u)+\operatorname{TGV}_{\alpha}(u)=\min_{u\in U,w\in V}\mathcal{F}(u)+\mathcal{G}_{1}^{*}(\mathcal{K}_{1}u)+\alpha_{1}\|\nabla u-w\|_{1}+\alpha_{0}\|\mathcal{E}w\|_{1}.$ (6.1)
A similar denoising problem in multi-modal electron tomography with KL divergence and TGV regularization is studied in [29]. Here $U$ and $V$ are finite-dimensional Hilbert spaces and $w=(w^{1},w^{2})$. The two-dimensional problem (6.1) is standard and all definitions along with discretizations will be made clear afterwards. Let us first reformulate the problem (6.1) as the saddle-point problem. $\mathcal{G}_{1}^{*}$ is the appropriate data term measure with blurred data given by the following KL divergence [29, Equation (9)], which is a variant of [28, Equation (13)].
$\mathcal{G}^{*}_{1}(s)=\sum_{i=1}^{N}\begin{cases}s_{i}-b_{i}\log s_{i}&\text{if }s_{i}\geq 0,\\ \infty&\text{ else.}\end{cases}$ (6.2)
where $b_{i}$ is the blurred image and we use the convention $00\log 0=$ 0 and $-a\log(0)=\infty$ for $a>0$. It can be readily checked that its Fenchel conjugate is
$\mathcal{G}_{1}(z)=\sum_{i=1}^{N}\begin{cases}\frac{-b_{j}z_{j}}{1-z_{j}}-\frac{b_{j}}{1-z_{j}}+b_{j}\log(\frac{b_{j}}{1-z_{j}}),&\text{ if }z_{j}\leq 1\text{ and }(b_{j}=0\text{ or }z_{j}<1),\\ \infty&\text{ else, }\end{cases}$ (6.3)
and its resolvent is $(I+\tau\partial\mathcal{G}_{1})^{-1}(z)=\frac{1}{2}(z+1-\sqrt{(z-1)^{2}+4\tau b})$. Here $\mathcal{F}(u)=\mathcal{I}_{\{0\leq u\leq 1\}}(u)$ is a box constraint. It forces $u$ to be non-negative by the requirement of KL divergence and to have an upper bound $1$. The equivalent saddle-point problem of (6.1) is as follows
$\min_{u\in U,w\in V}\max_{s\in U,p\in V,q\in W}\langle\mathcal{K}_{1}u,s\rangle_{U}+\langle\nabla u-w,p\rangle_{V}+\langle\mathcal{E}w,q\rangle_{W}+\mathcal{F}(u)$ (6.4)
$-\mathcal{G}_{1}(s)-\sum_{i=1}^{2}\mathcal{I}_{\{\|p^{i}\|_{\infty}\leq\alpha_{1}\}}(p^{i})-\sum_{j=1}^{4}\mathcal{I}_{\{\|q^{j}\|_{\infty}\leq\alpha_{0}\}}(q^{j}).$
Let us reformulate (6.4) into (1.1). We have the data $\mathcal{F}(x)=\mathcal{F}(u)=\mathcal{I}_{\{0\leq u\leq 1\}}(u)$ with $x=(u,w)$ and $\mathcal{G}^{*}_{i}(y_{i})=\alpha_{1}\|y_{i}\|_{1},$ $i=2,3$, $\mathcal{G}^{*}_{i}(y_{i})=\alpha_{0}\|y_{i}\|_{1},$ $i=4,5,6,7$. $\mathcal{G}_{i}$ is the Fenchel conjugate of $\mathcal{G}_{i}^{*}$. The dual variables $y=(y_{1},\ldots,y_{7})=(s,p,q)$ with $p=(p^{1},p^{2})\in V:=U^{2}$ and $q=(q^{1},q^{2},q^{3},q^{4})\in W:=U^{4}$ is thus obtained. The linear operator $\mathcal{K}:X\rightarrow Y$ as in (1.1) is
$\mathcal{K}=\left(\begin{array}[]{cc}\mathcal{K}_{1}&0\\ \nabla&-I\\ 0&\mathcal{E}\end{array}\right),\quad X=U\times V,\quad Y=U\times V\times W,$ (6.5)
where $\mathcal{K}_{1}=\kappa\ast$ is a convolution operator with kernel $\kappa$ for deblurring problems.
We then go to the details of the discrete setting including the operators and spaces. In (6.1), $\mathcal{E}w=(\nabla w+\nabla w^{T})/2$ is a discrete symmetrized derivative. $\nabla$ is a discrete gradient operator with finite forward differences, i.e., $(\nabla u)=(\partial_{x}^{+}u,\partial_{y}^{+}u)^{T}$ [12]. For discretization, set image domain $\Omega\subset\mathbf{Z}^{2}$ as the discretized grid $\Omega=\{(i,j)\mid i,j\in\mathbb{N},0\leq i\leq d_{1}-1,0\leq j\leq d_{2}-1\}$, where $d_{1},d_{2}$ are the image dimensions [10]. The image space is defined as $U=\{u:\Omega\rightarrow\mathbf{R}\}$ with the standard $L^{2}$ scalar product. For $t\in[1,\infty)$, let us introduce the following norms: $\|u\|_{\infty}=\max_{(i,j)\in\Omega}|u_{i,j}|$,
$\begin{array}[]{l}\|u\|_{t}=(\sum_{(i,j)\in\Omega}|u_{i,j}|^{t})^{1/t},\ \ \|p\|_{1}=\sum_{l=1}^{2}\sum_{(i,j)\in\Omega}|p_{i,j}^{l}|,\ \ \|q\|_{1}=\sum_{l=1}^{4}\sum_{(i,j)\in\Omega}|q_{i,j}^{l}|.\end{array}$
The adjoint operator of $\nabla:U\rightarrow V$ is $\nabla^{*}=-$ div (e.g., $\operatorname{div}p=\partial_{x}^{-}p^{1}+\partial_{y}^{-}p^{2}$) with homogeneous Neumann and Dirichlet boundary conditions, respectively and we refer to [12] for details. Next we describe the discretization of the anisotropic symmetric derivative $\mathcal{E}$ in the space $W=\{u:\Omega\rightarrow S^{2\times 2}\}$, which leads to the anisotropic second-order TGV which makes it possible to separate dual variables. We identify elements $q\in W$ and employ the scalar product $\langle q,q^{\prime}\rangle_{W}=\langle q^{1},(q^{\prime})^{1}\rangle_{U}+\langle q^{2},(q^{\prime})^{2}\rangle_{U}+\langle q^{3},(q^{\prime})^{3}\rangle_{U}+\langle q^{4},(q^{\prime})^{4}\rangle_{U}$. The symmetrized derivative can be defined as follows
$\mathcal{E}w=\left[\begin{array}[]{cc}\partial_{x}^{+}w^{1}&\frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})\\ \frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})&\partial_{y}^{+}w^{2}\end{array}\right]=\left[\begin{array}[]{c}\partial_{x}^{+}w^{1}\\ \partial_{y}^{+}w^{2}\\ \frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})\\ \frac{1}{2}(\partial_{y}^{+}w^{1}+\partial_{x}^{+}w^{2})\end{array}\right]$ (6.6)
where the second equation has to be understood in terms of the identification $W=U^{4}$. Consequently, the negative adjoint realizes a discrete negative divergence operator according to $\langle\mathcal{E}w,q\rangle_{W}=-\langle w,\operatorname{div}q\rangle_{V}$ for all $w\in V,q\in$ $W$. With $\Delta w=(\Delta w^{1},\Delta w^{2}),$ we can express the operator $\mathcal{E}^{*}\mathcal{E}w=-\frac{1}{2}\Delta w-\frac{1}{2}\delta w$. Here $\operatorname{div}q$ and $\delta w$ are gives as
$\operatorname{div}q:=\left[\begin{array}[]{l}\partial_{x}^{-}q^{1}+\frac{1}{2}\partial_{y}^{-}q^{3}+\frac{1}{2}\partial_{y}^{-}q^{4}\\ \frac{1}{2}\partial_{x}^{-}q^{3}+\frac{1}{2}\partial_{x}^{-}q^{4}+\partial_{y}^{-}q^{2}\end{array}\right],\quad\delta w:=\left[\begin{array}[]{l}\partial_{x}^{-}\partial_{x}^{+}w^{1}+\partial_{y}^{-}\partial_{x}^{+}w^{2}\\ \partial_{y}^{-}\partial_{y}^{+}w^{2}+\partial_{x}^{-}\partial_{y}^{+}w^{1}\end{array}\right].$
Moreover, we confirm that $q^{3}=q^{4}$ and only $6$ dual variables will be randomly updated.
6.1.1 The preconditioner
Setting $N_{2}=I$, our goal is to find efficient preconditioners to solve $Tx^{k+1}=b^{k}$, with $x^{k+1}=(u^{k+1},w^{k+1})^{T}$ and $b^{k}=\bar{x}^{k}-\sigma\mathcal{K}^{*}\bar{y}^{k}=(b_{u}^{k},b_{w}^{k})^{T}$ and
$T=\left[\begin{array}[]{cc}\sigma\tau\mathcal{K}_{1}^{*}\mathcal{K}_{1}+I-\sigma\tau\Delta&\sigma\tau\mathrm{div}\\ -\sigma\tau\nabla&(1+\sigma\tau)I+\sigma\tau\mathcal{E}^{*}\mathcal{E}\end{array}\right].$ (6.7)
$T$ is challenging for preconditioning due to $\mathcal{E}^{*}\mathcal{E}$. Fortunately, it can be readily checked that
$T^{\prime}=\left[\begin{array}[]{cc}(1+\sigma\tau)I-\sigma\tau\Delta&\sigma\tau\operatorname{div}\\ -\sigma\tau\nabla&(1+\sigma\tau)I-\sigma\tau\Delta\end{array}\right]$ (6.8)
satisfies $T^{\prime}-T\succeq 0$ [9]. It is clear that feasible preconditioners $M$ for $T^{\prime}$ are also feasible for $T$ (Since $M\succeq T^{\prime}$, we have $M\succeq T$). Consequently, we obtain
$x^{k+1}=x^{k}+M^{-1}\left(b^{k}+\left(T^{\prime}-T\right)x^{k}-T^{\prime}x^{k}\right),$ (6.9)
which is a preconditioned iteration for $T^{\prime}$. Now we adjust the right-hand side as follows.
$\left(b^{\prime}\right)^{k}$ $:=b^{k}+\left(T^{\prime}-T\right)x^{k}$ $=\left[\begin{array}[]{c}b_{u}^{k}+(\sigma\tau I-\sigma\tau\mathcal{K}_{1}^{*}\mathcal{K}_{1})u^{k}\\ b_{w}^{k}+\frac{\sigma\tau}{2}\left(\delta-\Delta\right)w^{k}\end{array}\right].$ (6.10)
Now solving $Tx^{k+1}=b^{k}$ by preconditioned iteration: $x^{k+1}=x^{k}+M^{-1}(b^{k}-Tx^{k})$ is transformed into a new preconditioned iteration: $x^{k+1}=x^{k}+M^{-1}((b^{\prime})^{k}-T^{\prime}x^{k})$, which is much easier to compute and implement. We refer to [9, Section 5.2] for the block symmetric Gauss–Seidel preconditioners.
6.1.2 Parameters
Here let us set $U=\mathbb{R}^{d_{1}\times d_{2}}$ with $d_{1}=512$, $d_{2}=357$ and $\mathcal{K}_{1}=k\ast$ is a convolution kernel with $40\times 40$ motion blur. It can be checked that the Bregman distance for the ergodic sequence is bounded for a bounded domain containing a saddle point.
Set the sampling to be uniform, i.e., $p_{i}=\frac{1}{n},\,n=6$. The regularization parameters are $\alpha_{0}=10^{-4}$ and $\alpha_{1}=5\times 10^{-5}$. The random iteration sequence is determined by random seed 5. The initial value of all algorithms is set to be zero vector, except the initial primal update term of $u$ in (6.4) is set to be the blurred image. We mainly compared with the SPDHG as in [13] and we refer to [21] for the case stochastic updates both on the primal and dual variables. The step sizes for compared algorithms are as follows:
$\bullet$ PDHG, SPDHG: Primal-Dual Hybrid Gradient algorithm (PDHG) [12] and stochastic PDHG [13]. We use constant step size $\tau=\sigma=1/\|\mathcal{K}\|$, $\|\mathcal{K}\|=\sqrt{14}$ for PDHG and $\tau=0.25$, $\sigma=0.12$ for SPDHG, which is much faster than the step size in [13].
$\bullet$ PDR, SPDR: One-step block symmetric Red-Black Gauss–Seidel is the preconditioner for the elliptic system equation. The step sizes are chosen as $\tau=0.1,\,\sigma=5$.
$\bullet$ p-RPDR (f-RPDR), p-SRPDR (f-SRPDR): partial (fully) relaxed PDR, partial (fully) stochastic relaxed PDR. Again, the preconditioner and the step sizes are the same as PDR. The relaxation parameters are all set to 1.9.
Now, let us turn to the numerical ergodic Bregman distance. It can be checked that $x^{k+1}_{\operatorname{test}}$, $s^{k+1}_{\operatorname{test}}$ and $y^{k+1}_{\operatorname{test},i},$ $i=\{2,\ldots,7\}$ produced by SPDR satisfy the constraints of $\mathcal{F}$, $\mathcal{G}_{1}$ and $\mathcal{G}_{i},$ $i=\{2,\ldots,7\}$ correspondingly. Still, by the convexity of these functions, the ergodic sequence $(x^{k+1}_{\operatorname{test},K},y^{k+1}_{\operatorname{test},K})$ also satisfies the corresponding constraints. We thus conclude that the Bregman distance of ergodic sequence is finite and can be written as $\langle\mathcal{K}x_{\operatorname{test},K},y^{*}\rangle-\mathcal{G}_{1}(s^{*})-\langle\mathcal{K}x^{*},y_{\operatorname{test},K}\rangle+\mathcal{G}_{1}(s_{\operatorname{test},K})$. Here $(x^{*},y^{*})$ is the approximate saddle point which is calculated for more than $10^{5}$ epochs.
(a) Convergence rate of normalized ergodic Bregman distance concerning epochs.
(b) Convergence rate of normalized primal error concerning epochs.
(c) PSNR with respect to epochs.
Figure 1: TGV-KL deblurring problem. (a), $\emph{(}b)$ are normalized by $d_{1}\times d_{2}$ and in a double-logarithmic scale. Image $\emph{(}b)$ shows the speed comparison measured in terms of primal error, which is given by $(\mathcal{P}(x^{K})-\mathcal{P}(x^{*}))/\mathcal{P}(x^{*})$.
(a) Degraded image
(b) PDHG, 300 epochs
(c) PDR, 300 epochs
(d) p-RPDR, 300 epochs
(e) f-RPDR, 300 epochs
(f) Approximate primal part of saddle point
(g) SPDHG, 300 epochs
(h) SPDR, 300 epochs
(i) p-SRPDR,300 epochs
(j) f-SRPDR, 300 epochs
(k) Details of a
(l) Details of PDHG
(m) Details of PDR
(n) Details of P-PDR
(o) Details of f-RPDR
(p) Details of f
(q) Details of SPDHG
(r) Details of SPDR
(s) Details of p-SRPDR
(t) Details of f-SRPDR
Figure 2: Examples of variational TGV-KL deblurring after 300 epochs with uniform sampling.
Figure 1 (a) shows that the Bregman distance converges at a rate of $\mathcal{O}(1/K)$ in practice. Figure 1 (b) shows that the primal errors are not continuous at the beginning epochs. This happens because $\mathcal{G}_{i}$ in (6.2) can be infinite. Additionally, the proposed SPDR and SRPDR methods demonstrate high efficiency and strong performance compared to SPDHG, both in terms of Bregman distance and normalized primal error. Notably, f-SRPDR achieves a higher and more stable PSNR, as shown in Figure 1. Figure 2 shows that the proposed SPDR, p-SRPDR, and f-SRPDR can give clearer reconstructions compared with PDR, f-RPDR, f-RPDR, SPDHG, and PDHG with the same epochs.
6.2 Binary Classification with real data
In this subsection, we present the results of solving a binary classification problem using the real datasets Gisette [23] and Madelon [23], both part of the LIBSVM data collection [14]. These two datasets are selected to illustrate different relations between sample size and feature dimensionality. Each data point is represented as a pair $(a_{i},b_{i})$, where $a_{i}\in\mathbb{R}^{d}$ is the feature vector and $b_{i}$ is the binary class label, taking values in $\{-1,1\}$. Our objective is to solve the following regularized empirical risk minimization problem:
$\min_{x\in\mathbb{R}^{d}}\frac{1}{n}\sum^{n}_{i=1}\phi_{i}(a_{i}^{T}x)+\frac{\lambda}{2}\|x\|_{2}^{2},\ \ \text{where}\ \phi_{i}(z)=\begin{cases}0&\text{if}b_{i}z\geq 1\\ \frac{1}{2}-b_{i}z&\text{if }b_{i}z\leq 0\\ \frac{1}{2}(1-b_{i}z)^{2}&\text{otherwise},\end{cases}$ (6.11)
is the smoothed hinge loss, originally introduced by [39] and addressed with data term by [43, Section 7.2]. Define $A=[a_{1},\dots,a_{n}]^{T}$. The corresponding $\mathcal{K}$, $\mathcal{G}_{i}^{*}(\mathcal{K}_{i}x)$ and $\mathcal{F}(x)$ in the form of (1.3) read as
$\mathcal{K}=\frac{1}{n}A,\quad\mathcal{G}^{*}_{i}(\mathcal{K}_{i}x)=\frac{1}{n}\phi_{i}(n\mathcal{K}_{i}x)=\frac{1}{n}\phi_{i}(a_{i}^{T}x),\quad\mathcal{F}(x)=\frac{\lambda}{2}\|x\|^{2}_{2}.$
The conjugate function of hinge lose $\phi_{i}$ is given as $\phi^{*}_{i}(z)=b_{i}z+\frac{1}{2}z^{2}+\mathcal{I}_{\{b_{i}z\in[-1,0]\}}(z)$, and $\mathcal{G}_{i}(z)=(1/n)\phi^{*}_{i}(z)$. Then the corresponding saddle point problem of (6.11) is written as,
$\min_{x\in\mathbb{R}^{d}}\max_{y\in\operatorname{dom}{\mathcal{G}}}\frac{\lambda}{2}\|x\|^{2}_{2}+\sum^{n}_{i=1}(\langle\mathcal{K}_{i}x,y_{i}\rangle-\frac{1}{n}\phi_{i}^{*}(y_{i})),\ \text{with}\quad\mathcal{G}(y):=\sum^{n}_{i=1}\mathcal{G}_{i}(y_{i}),\ \text{and}\ y=(y_{1},\ldots,y_{n})^{T}.$ (6.12)
Further, the resolvent of $\mathcal{F}$ is derived as $(I+\tau\partial\mathcal{F})^{-1}(z)={z}/(1+\lambda\tau)$. Set $l_{i}=\min(-b_{i},0)$ and $r_{i}=\max(-b_{i},0)$ thus the resolvent of $\mathcal{G}_{i}$ is derived as $(I+\sigma\partial\mathcal{G}_{i})^{-1}(y)=\text{Project}_{[l_{i},r_{i}]}((ny-\sigma b_{i})/(n+\sigma))$.
6.2.1 Characteristics of datasets and parameters for the algorithms
The Gisette dataset represents the case where $n<d$, with $n=1000$ and $d=5000$, while the Madelon dataset represents the case where $n>d$, with $n=2000$ and $d=500$. We will evaluate our algorithms on these two datasets using two regularization parameters: $\lambda=10^{-6}$ and $\lambda=10^{-4}$. The initial values for all algorithms are set to be zero matrices. The step sizes for each algorithm are provided below.
$\bullet$ PDHG, SPDHG: We use a step size of $\sigma={2}/\|\mathcal{K}\|$ and $\tau={1}/({2\|\mathcal{K}\|})$ for PDHG and SPDHG on the dataset Gisette to achieve better performance. For the dataset Madelon, we employ $\sigma={2}/\|\mathcal{K}\|$ and $\tau={1}/({2\|\mathcal{K}\|})$ for PDHG. In contrast, for SPDHG, we set $\sigma=\tau={1}/(15\|\mathcal{K}\|)$, due to the significant fluctuations observed with larger step sizes on this dataset.
$\bullet$ PDR, SPDR: The step sizes are chosen as $\sigma=\tau=0.3$ for both algorithms and both datasets Madelon and Gisette. The precondition iteration is set to be the two-step Richardson [10, Example 2.11, Proposition 2.15]. Thus the corresponding preconditioner $M$ in preconditioned iteration (2.2) is given by $(1+\sigma\tau\|\mathcal{K}\|^{2})I$.
$\bullet$ p-RPDR (f-RPDR), p-SRPDR (f-SRPDR): The step sizes and the preconditioners are set to be the same as those used in PDR and SPDR. The relaxation parameters are all set to 1.9.
With these settings, we tested the primal error of eight algorithms in these two datasets. In Figure 3 and Figure 4, the vertical axis is the relative primal error calculated by $\mathcal{P}(x^{k})-\mathcal{P}(x^{*})$ in logarithmic scale. The minimum loss $\mathcal{P}(x^{*})$ for both datasets is calculated by f-RPDR after 750,000 epochs for Gisette and after 500,000 epochs for Madelon.
Figure 3: Convergence rate of relative primal error concerning epochs for dataset Gisette and Madelon. The regularization parameter is set to be $1\times 10^{-6}$.
Figure 4: Convergence rate of relative primal error concerning epochs for dataset Gisette and Madelon. The regularization parameter is set to be $1\times 10^{-4}$.
From Figure 3 and Figure 4, We observe that our proposed SRPDR and SPDR methods are highly efficient and stable, with fewer fluctuations compared to SPDHG. We also find that the performance of these algorithms depends a lot on step sizes. For some algorithms, especially deterministic ones, the primal errors decrease quickly at first but then slow down considerably. As a result, the loss for our proposed stochastic algorithms after 1000 epochs can reach a level nearly the same as f-RPDR after 750,000 epochs on dataset Gisette. This is why the relative primal errors of our stochastic algorithms can drop below $10^{-16}$ on dataset Giesette. It is worth noting that SRPDRQ can also be used here. SRPDRQ is more compact because it does not require the updates of $\bar{x}^{k+1}$ compared to SRPDR. However, since SRPDR is commonly used and the difference in primal error between SRPDR and SRPDRQ is not significant, we only present SPDR, f-SRPDR and p-SRPDR in this experiment. A comparison between SRPDR and SRPDRQ is included in the Appendix 3.
7 Conclusion
We proposed a stochastic and relaxed PDR framework to solve general saddle-point problems for separable dual variables cases. We showed the almost sure convergence of SRPDR along with SRPDRQ. We gave the $\mathcal{O}(1/K)$ convergence rate of the ergodic iteration sequences regarding the restricted primal-dual gap functions and the primal errors. For SRPDR, we will consider local linear convergence involving metric subregularity as in [1] and design highly efficient preconditioners for more complicated operators in applications including the Radon transforms [29] for tomography.
8 Appendix
Proof of Proposition 2.
By the equation (2.7) and the iteration of RPDR, we see
$\mathcal{T}u^{k}=I_{\varrho_{k}}^{-1}(\mathcal{T}_{R}-I)u^{k}+u^{k}=I_{\varrho_{k}}^{-1}(u^{k+1}-u^{k})+u^{k}.$ (8.1)
Substituting $u_{t}^{k+1}=\mathcal{T}u^{k}$ to the equation (2.6) in Proposition 1, we arrive at
$\mathcal{L}(x_{\operatorname{test}}^{k+1},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k+1})$ $\leq\langle u^{k}-(I_{\varrho_{k}}^{-1}(u^{k+1}-u^{k})+u^{k}),(I_{\varrho_{k}}^{-1}(u^{k+1}-u^{k})+u^{k})-u\rangle_{\mathbf{M}}$
$=\langle I_{\varrho_{k}}^{-1}(u^{k}-u^{k+1}),I_{\varrho_{k}}^{-1}(u^{k+1}-u^{k})+u^{k}-u\rangle_{\mathbf{M}}$
$=\langle I_{\varrho_{k}}^{-1}(u^{k}-u^{k+1}),(I_{\varrho_{k}}^{-1}-I)(u^{k+1}-u^{k})+u^{k+1}-u\rangle_{\mathbf{M}}$
$=\langle I_{\varrho_{k}}^{-1}(u^{k}-u^{k+1}),u^{k+1}-u\rangle_{\mathbf{M}}-\langle I_{\varrho_{k}}^{-1}(u^{k+1}-u^{k}),(I_{\varrho_{k}}^{-1}-I)(u^{k+1}-u^{k})\rangle_{\mathbf{M}}$
$=\langle u^{k}-u^{k+1},u^{k+1}-u\rangle_{\mathbf{M}I_{\varrho_{k}}^{-1}}-\|u^{k+1}-u^{k}\|_{\mathbf{M}I_{\varrho_{k}}^{-1}(I_{\varrho_{k}}^{-1}-I)}^{2}.$ (8.2)
By the polarization identity and the positive semidefiniteness of $\mathbf{M}I_{\varrho_{k}}^{-1}$, we obtain
$\langle u^{k}-u^{k+1},u^{k+1}-u\rangle_{\mathbf{M}I_{\varrho_{k}}^{-1}}=\frac{1}{2}\big{(}\|u^{k}-u\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}\big{)}.$ (8.3)
Substituting the above equality to (8.2) and noting $\mathbf{M}I_{\varrho_{k}}^{-1}+2\mathbf{M}I_{\varrho_{k}}^{-1}(I_{\varrho_{k}}^{-1}-I)=\mathbf{M}(2I-I_{\varrho_{k}})I^{-2}_{\rho_{k}}$, we get (2.9). ∎
We need the following lemma first to prove Lemma 2.
Lemma 12.
Suppose $\rho_{k}\in(0,2)$ be a nondecreasing sequence, and $\rho_{k}\rightarrow\bar{\rho}$ with $\bar{\rho}\in(0,2)$. If $\mathbf{c}^{k}\rightharpoonup\mathbf{c}$, $\bar{\rho}\neq 1$, and the sequence $\{\mathbf{a}^{k}\}$ satisfy
$\mathbf{a}^{k+1}-(1-\rho_{k})\mathbf{a}^{k}=\rho_{k}\mathbf{c}^{k},$ (8.4)
then we have $\mathbf{a}^{k}\rightharpoonup\mathbf{c}$. For the special case $\rho_{k}\equiv 1$ with $\bar{\rho}=1$, we also have $\mathbf{a}^{k}\rightharpoonup\mathbf{c}$.
Proof.
It can be seen that
$\rho_{k}\mathbf{c}^{k}=(\rho_{k}-\bar{\rho})\mathbf{c}^{k}+\bar{\rho}\mathbf{c}^{k}.$
Since $\mathbf{c}^{k}$ is bounded and $\rho_{k}\rightarrow\bar{\rho}$, we have $\rho_{k}\mathbf{c}^{k}\rightharpoonup\bar{\rho}\mathbf{c}$ with the assumption $\mathbf{c}^{k}\rightharpoonup\mathbf{c}$. While $\rho_{k}\equiv 1$ with $\bar{\rho}=1$, we conclude that $\mathbf{a}^{k+1}=\mathbf{c}^{k}\rightharpoonup\mathbf{c}$.
If $\bar{\rho}\neq 1$, then there exists $k_{0}$, such that either $\rho_{k}\in(1,2)$ while $\bar{\rho}\in(1,2)$ or $\rho_{k}\in(0,\bar{\rho})$ while $\bar{\rho}\in(0,1)$ for all $k\geq k_{0}$. Henceforth, denote $\alpha_{k+1}:=(1-\rho_{k})$, $\mathbf{b}^{k+1}:=\mathbf{a}^{k+1}-\alpha_{k+1}\mathbf{a}^{k}$ for $k\geq k_{0}$ and define $\mathbf{b}^{k_{0}}:=\mathbf{a}^{k_{0}}$ along with $\alpha_{k_{0}}=1$ for convenience. By induction with direct calculation from $k_{0}$, we obtain
$\mathbf{a}^{k_{0}+k}=\bigl{(}\sum_{k^{\prime}=0}^{k}\mathbf{b}^{k_{0}+k^{\prime}}\prod_{k^{\prime\prime}=0}^{k^{\prime}}\alpha_{k_{0}+k^{\prime\prime}}^{-1}\bigr{)}/\prod_{k^{\prime}=0}^{k}\alpha_{k_{0}+k^{\prime}}^{-1},\ \forall k\geq 0.$
It can be seen that $\mathbf{b}^{k+1}\rightharpoonup\mathbf{b}:=\bar{\rho}\mathbf{c}$ due to $\mathbf{b}^{k+1}=\rho_{k}\mathbf{c}^{k}$ for $k>k_{0}$ with (8.4) and $\alpha_{k}\rightarrow\alpha:=(1-\bar{\rho})$. For the case $\bar{\rho}<1$, i.e., $\alpha_{k}<1$ for $k>k_{0}$, for any $f$, applying Stolz-Cesàro theorem, we have
$\lim_{k\rightarrow\infty}\langle\mathbf{a}^{k_{0}+k},f\rangle=\lim_{k\to\infty}\frac{\langle\mathbf{b}^{k_{0}+k},f\rangle\prod_{k^{\prime\prime}=0}^{k}\alpha_{k_{0}+k^{\prime\prime}}^{-1}}{\prod_{k^{\prime\prime}=0}^{k}\alpha_{k_{0}+k^{\prime\prime}}^{-1}-\prod_{k^{\prime\prime}=0}^{k-1}\alpha_{k_{0}+k^{\prime\prime}}^{-1}}=\lim_{k\to\infty}\frac{\langle\mathbf{b}^{k_{0}+k},f\rangle}{1-\alpha_{k_{0}+k}}=\frac{\langle\mathbf{b},f\rangle}{1-\alpha}=\langle\mathbf{c},f\rangle.$
We thus conclude that $\mathbf{a}^{k}\rightharpoonup\mathbf{c}$. For the case $\bar{\rho}\in(1,2)$, since $\alpha_{k}<0$ for $k\geq k_{0}$, the proof is similar and is omitted here. ∎
Proof of Lemma 2.
Regarding the first point, notice the fixed points of (2.7) (i.e., RPDR) satisfy
$u^{*}=(I-I_{\varrho_{k}})u^{*}+I_{\varrho_{k}}\mathcal{T}u^{*},$
which is equivalent to $\mathcal{T}u^{*}=u^{*}$ by the invertibility of $I_{\varrho_{k}}$. Thus the first point can be obtained with [10, Lemma 2.7] (or [7, Lemma 3]).
For the second point of this lemma, writing (2.7) component-wisely and denoting $\bar{I}_{k_{i}}:=\text{Diag}[\rho_{k_{i},x},\rho_{k_{i},y}]$, we have
$\begin{pmatrix}x^{k_{i}+1}\\ y^{k_{i}+1}\end{pmatrix}$ $=(1-\rho_{k_{i}})\begin{pmatrix}x^{k_{i}}\\ y^{k_{i}}\end{pmatrix}+\rho_{k_{i}}\begin{pmatrix}N_{1}&\sigma K^{*}\\ -\tau K&N_{2}\end{pmatrix}^{-1}\begin{pmatrix}(N_{1}-I)x^{k_{i}}+\bar{x}^{k_{i}}\\ (N_{2}-I)y^{k_{i}}+\bar{y}^{k_{i}}\end{pmatrix},$ (8.5)
$\begin{pmatrix}\bar{x}^{k_{i}+1}\\ \bar{y}^{k_{i}+1}\end{pmatrix}$ $=\begin{pmatrix}\bar{x}^{k_{i}}\\ \bar{y}^{k_{i}}\end{pmatrix}+\bar{I}_{k_{i}}\begin{pmatrix}(I+\sigma\partial\mathcal{F})^{-1}(2x_{t}^{k_{i}+1}-\bar{x}^{k_{i}})-x_{t}^{k_{i}+1}\\ (I+\tau\partial\mathcal{G})^{-1}(2y_{t}^{k_{i}+1}-\bar{y}^{k_{i}})-y_{t}^{k_{i}+1}\end{pmatrix}.$ (8.6)
Denote $A_{N}:=\begin{pmatrix}N_{1}&\sigma K^{*}\\ -\tau K&N_{2}\end{pmatrix}$ and $A:=\begin{pmatrix}I&\sigma K^{*}\\ -\tau K&I\end{pmatrix}$. Multiplying $A_{N}$ to both sides of (8.5), by rearranging and cancelling some terms, we arrive at
$\begin{pmatrix}N_{1}(x^{k_{i}+1}-x^{k_{i}})+\sigma K^{*}(y^{k_{i}+1}-y^{k_{i}})\\ -\tau K(x^{k_{i}+1}-x^{k_{i}})+N_{2}(y^{k_{i}+1}-y^{k_{i}})\end{pmatrix}=\rho_{k_{i}}\begin{pmatrix}\bar{x}^{k_{i}}-x^{k_{i}}-\sigma K^{*}y^{k_{i}}\\ \bar{y}^{k_{i}}-y^{k_{i}}+\tau Kx^{k_{i}}\end{pmatrix}.$ (8.7)
Adding $(x^{k_{i}}-x^{k_{i}+1},y^{k_{i}}-y^{k_{i}+1})^{T}$ to both sides of above equation, we have
$\begin{pmatrix}I&\sigma K^{*}\\ -\tau K&I\end{pmatrix}\begin{pmatrix}x^{k_{i}+1}\\ y^{k_{i}+1}\end{pmatrix}+h^{k_{i}}=\rho_{k_{i}}\begin{pmatrix}\bar{x}^{k_{i}}\\ \bar{y}^{k_{i}}\end{pmatrix}+(1-\rho_{k_{i}})\begin{pmatrix}I&\sigma K^{*}\\ -\tau K&I\end{pmatrix}\begin{pmatrix}x^{k_{i}}\\ y^{k_{i}}\end{pmatrix},$ (8.8)
where $h^{k_{i}}=((N_{1}-I)(x^{k_{i}+1}-x^{k_{i}}),(N_{2}-I)(y^{k_{i}+1}-y^{k_{i}}))^{T}\rightarrow 0$ by the assumption. We conclude that $A$ has a bounded inverse since $I+\sigma\tau K^{*}K$ is positive definite. We thus can rewrite (8.8) as follows
$\begin{pmatrix}x^{k_{i}+1}\\ y^{k_{i}+1}\end{pmatrix}-(1-\rho_{k_{i}})\begin{pmatrix}x^{k_{i}}\\ y^{k_{i}}\end{pmatrix}=\rho_{k}A^{-1}\begin{pmatrix}\bar{x}^{k_{i}}\\ \bar{y}^{k_{i}}\end{pmatrix}-A^{-1}h^{k_{i}}.$ (8.9)
By the assumption of this lemma, $\{\bar{y}^{k}\}_{k}$ and $\{\bar{x}^{k}\}_{k}$ are bounded. There thus exist weakly convergent subsequences and denote them as $\{\bar{x}^{k}_{i}\}_{k}$ and $\{\bar{y}^{k}_{i}\}_{k}$ with corresponding weak limit $\bar{x}$ and $\bar{y}$. Since $A^{-1}$ is linear and bounded, then $\rho_{k}A^{-1}(\bar{x}^{k}_{i},\bar{y}^{k}_{i})^{T}-A^{-1}h^{k}$ is weakly convergent with weak limit $\rho^{*}A^{-1}(\bar{x},\bar{y})^{T}$. Furthermore, applying Lemma 12 to each component of (8.9), we arrive at
$\begin{pmatrix}x^{k}_{i}\\ y^{k}_{i}\end{pmatrix}\rightharpoonup\begin{pmatrix}x^{*}\\ y^{*}\end{pmatrix}:=A^{-1}\begin{pmatrix}\bar{x}^{*}\\ \bar{y}^{*}\end{pmatrix}.$ (8.10)
We now turn to prove $(x^{*},y^{*})$ is a saddle point for (1.2). Actually we just need to prove that $(x_{\operatorname{test}}^{k_{i}+1},y_{\operatorname{test}}^{k_{i}+1})$ weakly converge to $(x^{*},y^{*})$. With (8.6), we can obtain
$(\bar{x}^{k_{i}+1}-\bar{x}^{k_{i}})/\rho_{k_{i},x}+x_{t}^{k_{i}+1}=x_{\operatorname{test}}^{k_{i}+1},$ (8.11)
$(\bar{y}^{k_{i}+1}-\bar{y}^{k_{i}})/\rho_{k_{i},y}+y_{t}^{k_{i}+1}=y_{\operatorname{test}}^{k_{i}+1},$ (8.12)
and $(x_{t}^{k_{i}+1},y_{t}^{k_{i}+1})$ have the same weak limits as $(x_{\operatorname{test}}^{k_{i}+1},y_{\operatorname{test}}^{k_{i}+1})$. Together with (8.5), we see $(x^{k_{i}},y^{k_{i}})$ have the same weak limits with $(x_{t}^{k_{i}+1},y_{t}^{k_{i}+1})$. We thus conclude that
$x_{\operatorname{test}}^{k_{i}}\rightharpoonup x^{*},\quad y_{\operatorname{test}}^{k_{i}}\rightharpoonup y^{*}.$ (8.13)
Still with (8.5) and the definitions of $x_{\operatorname{test}}^{k_{i}+1}$ and $y_{\operatorname{test}}^{k_{i}+1}$, we get
$2x_{t}^{k_{i}+1}-\bar{x}^{k_{i}}-x_{\operatorname{test}}^{k_{i}+1}$ $\in\sigma\partial\mathcal{F}(x_{\operatorname{test}}^{k_{i}+1})$ (8.14)
$2y_{t}^{k_{i}+1}-\bar{y}^{k_{i}}-y_{\operatorname{test}}^{k_{i}+1}$ $\in\tau\partial\mathcal{G}(y_{\operatorname{test}}^{k_{i}+1}).$
Combining (8.11) and (8.14), we have
$\sigma\partial\mathcal{F}(x_{\operatorname{test}}^{k_{i}+1})+\sigma K^{*}y_{\operatorname{test}}^{k_{i}+1}\ni 2x_{t}^{k_{i}+1}-\bar{x}^{k_{i}}-x_{\operatorname{test}}^{k_{i}+1}+\sigma K^{*}y_{\operatorname{test}}^{k_{i}+1}$
$=x_{t}^{k_{i}+1}-\bar{x}^{k_{i}}+\sigma K^{*}y_{t}^{k_{i}+1}-(\bar{x}^{k_{i}+1}-\bar{x}^{k_{i}})/\rho_{k_{i},x}+\sigma K^{*}(\bar{y}^{k_{i}+1}-\bar{y}^{k_{i}})/\rho_{k_{i},y}.$
Still with (8.5), we obtain
$x^{k_{i}+1}-x^{k_{i}}=\rho_{k}(x_{t}^{k_{i}+1}-x^{k_{i}}),\quad y^{k_{i}+1}-y^{k_{i}}=\rho_{k}(y_{t}^{k_{i}+1}-y^{k_{i}}).$ (8.15)
Remembering the updates of $x_{t}^{k_{i}+1}$ and $y_{t}^{k_{i}+1}$ with $u^{t+1}=\mathcal{T}u^{k}$, we have
$x_{t}^{k_{i}+1}-\bar{x}^{k_{i}}+\sigma K^{*}y_{t}^{k_{i}+1}=(N_{1}-I)(x^{k_{i}}-x_{t}^{k_{i}+1})=(N_{1}-I)(x^{k_{i}}-x^{k_{i}+1})/\rho_{k_{i}}.$ (8.16)
Together with (8.11) and (8.15), we see
$\sigma\partial\mathcal{F}(x_{\operatorname{test}}^{k_{i}+1})+\sigma K^{*}y_{\operatorname{test}}^{k_{i}+1}\rightarrow 0.$
Similarly, we can get $\tau\partial\mathcal{G}(y_{\operatorname{test}}^{k_{i}+1})-\tau Kx_{\operatorname{test}}^{k_{i}+1}\rightarrow 0$. Since $(x,y)\mapsto\bigl{(}\mathcal{K}^{*}y+\partial\mathcal{F}(x),-\mathcal{K}x+\partial\mathcal{G}(y)\bigr{)}$ is maximally monotone, it is particular weak-strong closed. Together with (8.13), we finally arrive at
$0\in\mathcal{K}^{*}y^{*}+\partial\mathcal{F}(x^{*}),\quad 0\in-\mathcal{K}x^{*}+\partial\mathcal{G}(y^{*}).$
Proof of Theorem 1.
With the convexity of $\mathcal{F}$ and $\mathcal{G}$, we conclude that $\mathcal{L}(x,y)$ is convex in $x$ and concave in $y$. Together with the definitions of $x_{\operatorname{test}}^{K}$ and $y_{\operatorname{test}}^{K}$, noting that $\mathbf{M}I_{\varrho_{k}}^{-1}\succeq\mathbf{M}I_{\varrho_{k+1}}^{-1}$ for $k=0,\ldots,K$ and $K_{1}:=K+1$, it can be seen that
$\mathcal{H}(z_{\operatorname{test},K},z)=\mathcal{L}(x_{\operatorname{test}}^{K},y)-\mathcal{L}(x,y_{\operatorname{test}}^{K})\leq\frac{1}{K_{1}}\sum_{k=0}^{K}(\mathcal{L}(x_{\operatorname{test}}^{k},y)-\mathcal{L}(x,y_{\operatorname{test}}^{k}))$
$\leq\frac{1}{2K_{1}}\sum_{k=0}^{K}\left({\|u^{k}-u\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}-\|u^{k+1}-u\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}-\|u^{k+1}-u^{k}\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}}\right)$
$\leq\frac{1}{2K_{1}}\left({\|u^{0}-u\|_{\mathbf{M}I_{\varrho_{0}}^{-1}}^{2}-\|u^{K+1}-u\|_{\mathbf{M}I_{\varrho_{K}}^{-1}}^{2}-\sum_{k^{\prime}=0}^{K}\|u^{k+1}-u^{k}\|_{\mathbf{M}I_{\varrho_{k}}^{-1}}^{2}}\right)\leq\frac{1}{2K_{1}\rho_{l}}\|u^{0}-u\|_{\mathbf{M}}.$
Taking supremum regarding $z\in\mathbb{B}$ in the above inequality leads to this theorem. ∎
Remark 3.
Our experiments shows that SRPDR and SRPDRQ have no significant differences in primal errors with respect to the number of epochs. However SRPDRQ has lower complexity compared to SRPDR; further details can be found in Section 5.
To illustrate, we consider a single experiment on the binary classification problem (6.11) using the dataset Madelon. The regularization parameter is set to be $1\times 10^{-6}$. We compare the relative primal error between SPDRQ (f-SRPDRQ) and SPDR (f-SRPDR), furthermore, we compute the absolute difference in primal error between SPDRQ (f-SRPDRQ) and SPDR (f-SRPDR). The relaxation parameters for stochastic relaxed algorithms are all set to 1.9. The results are shown in Figure 5. We can see that the absolute difference between the given methods is of level $10^{-10}$ to $10^{-5}$, which does not have a significant difference.
(a) Relative primal error of SPDR and SPDRQ
(b) Relative primal error of f-SRPDR and f-SRPDRQ
(c) Absolut difference on primal error between SPDRQ (f-SRPDRQ) and SPDR (f-SRPDR)
Figure 5: Comparison between SRPDR and SPRDRQ on primal errors with respect to the epochs.